### v1

# Training on test proteins improves fitness, structure, and function prediction

Abstract
Data scarcity and distribution shifts often hinder the ability of machine learning models to generalize when applied to proteins and other biological data. Self-supervised pre-training on large datasets is a common method to enhance generalization. However, striving to perform well on all possible proteins can limit model’s capacity to excel on any specific one, even though practitioners are often most interested in accurate predictions for the individual protein they study. To address this limitation, we propose an orthogonal approach to achieve generalization. Building on the prevalence of self-supervised pre-training, we introduce a method for self-supervised fine-tuning at test time, allowing models to adapt to the test protein of interest on the fly and without requiring any additional data. We study our test-time training (TTT) method through the lens of perplexity minimization and show that it consistently enhances generalization across different models, their scales, and datasets. Notably, our method leads to new state-of-the-art results on the standard benchmark for protein fitness prediction, improves protein structure prediction for challenging targets, and enhances function prediction accuracy.
1 Introduction
[…]
[] Refer to caption
Figure 1: Example of test-time training (TTT) applied to protein folding. ESMFold poorly predicts the structure of the CASP14 target T1074 (shown in white) because the underlying language model ESM2 poorly fits the sequence, as indicated by the high perplexity (Fig. 2E in Lin et al. (2023) and the left panel here). Self-supervised test-time training of ESM2 on the single sequence of T1074 directly minimizes the perplexity, leading to improved structure prediction (better TM-score alignment and higher pLDDT predicted confidence). The same test-time training approach is also broadly applicable to other tasks, such as protein fitness and function prediction.
A comprehensive understanding of protein structure, function, and fitness is essential for advancing research in the life sciences (Subramaniam & Kleywegt, 2022; Tyers & Mann, 2003; Papkou et al., 2023). While machine learning models have demonstrated remarkable potential in protein research, they are typically optimized for achieving the best average performance across large datasets (Jumper et al., 2021; Watson et al., 2023; Yang et al., 2024; Kouba et al., 2023). However, biologists often focus their research on individual proteins or protein complexes involved for example in metabolic disorders (Ashcroft et al., 2023; Gunn & Neher, 2023), oncogenic signalling (Hoxhaj & Manning, 2020; Keckesova et al., 2017), neurodegeneration (Gulen et al., 2023; oh Seo et al., 2023), and other biological phenomena (Gu et al., 2022). In these scenarios, detailed insights into a single protein can lead to significant scientific advances.
Nonetheless, general machine learning models for proteins often struggle to generalize to individual case studies due to data scarcity (Bushuiev et al., 2023; Chen & Gong, 2022) and distribution shifts (Tagasovska et al., 2024; Feng et al., 2024). Bridging the gap between broad, dataset-wide optimizations and the precision required for studying single proteins in practical applications remains a critical challenge in integrating machine learning into biological research (Sapoval et al., 2022).
By contrast, in other application domains of machine learning, such as computer vision and natural language processing, customization and adaptation approaches have emerged as powerful tools to improve model performance in specific contexts (Ruiz et al., 2023; Hardt & Sun, 2023). Drawing inspiration from the test-time training (TTT) approach developed in computer vision to mitigate distribution shifts (Sun et al., 2020; Gandelsman et al., 2022), in this work we propose the TTT approach for proteins. Our method enables adapting protein models to one protein at a time, on the fly, and without the need for additional data. Given a model that has been pre-trained using masked language modeling, our method minimizes the perplexity of the model on a given test protein through self-supervised fine-tuning, which, in turn, results in improved downstream performance without updating the downstream task head.
The prevalence of masked modeling in protein machine learning makes our method broadly applicable to various downstream tasks. Empirically, we demonstrate its effectiveness across three key challenges in protein machine learning. First, TTT achieves state-of-the-art results on the ProteinGym dataset (Notin et al., 2024), a well-established benchmark for protein fitness prediction. Second, TTT enhances protein structure predictions with ESMFold (Lin et al., 2023) and ESM3 (Hayes et al., 2024) on challenging targets. Third, the application of TTT to protein function predictors results in improved classification of terpene synthase (TPS) substrates and protein subcellular localization.
In summary, the key contributions of this work are three-fold:
  1. 1.
    Motivated by the generalization challenges and distribution shifts prevalent in protein machine learning, we introduce a new test-time training (TTT) method [^1] https://github.com/anton-bushuiev/ProteinTTT that enables models to adapt to individual proteins on the fly and without requiring additional data.
  2. 2.
    We establish a link between our TTT approach and perplexity minimization, providing an insight into why this approach enhances model effectiveness.
  3. 3.
    We empirically validate TTT, achieving state-of-the-art results in protein fitness prediction, improving the protein structure prediction capabilities of well-established folding models, and enhancing protein function predictions in the tasks of terpene synthase substrate classification and protein localization prediction.
2 Background and related work
In this section, we present the context and related work that highlight the feasibility and broad applicability of test-time training (TTT) in the domain of machine learning on proteins. The widespread adoption of Y-shaped architectures relying on masked modeling enables the development of a general method for adapting protein models at test time via masking-based self-supervised fine-tuning.
The Y-shaped paradigm of learning.
In machine learning applied to biology, architectures often follow a Y-shaped paradigm (Gandelsman et al., 2022), consisting of a backbone feature extractor $f$, a self-supervised head $g$, and an alternative fine-tuning head $h$. During training, $g\circ f$ is first pre-trained, after which the pre-trained backbone $f$ is reused to fine-tune $h\circ f$ toward a downstream task. Here, $\circ$ denotes a composition of two machine learning modules (e.g., $g$ is applied on top of $f$ in $g\circ f$). At test time, the final model $h\circ f$ is fixed. Generalization is achieved by leveraging the rich knowledge encoded in the backbone $f$ and the task-specific priors acquired in the fine-tuning head $h$. This paradigm enables overcoming data scarcity during fine-tuning and underlies breakthrough approaches in protein structure prediction (Lin et al., 2023), protein design (Watson et al., 2023), protein function prediction (Yu et al., 2023), and other protein-related tasks (Hayes et al., 2024).
The backbone $f$ is typically a large neural network pre-trained in a self-supervised way on a large dataset using a smaller pre-training projection head $g$ (Hayes et al., 2024). The fine-tuning head $h$, however, depends on the application. In some cases, $h$ is a large neural network, repurposing the pre-trained model entirely (Watson et al., 2023; Lin et al., 2023); in others, $h$ is a minimal projection with few parameters (Cheng et al., 2023), or even without any parameters at all (i.e., a zero-shot setup, (Meier et al., 2021; Dutton et al., 2024)). In some cases, the fine-tuning head $h$ may also be a machine learning algorithm other than a neural network (Samusevich et al., 2024).
Masked modeling.
While the objective of fine-tuning $h\circ f$ is determined by the downstream application, the choice of pre-training objective for $g\circ f$ is less straightforward. Nevertheless, most methods employ various forms of masked modeling, i.e., optimizing the model weights to accurately reconstruct missing parts of proteins, regardless of the downstream application. Masked modeling pre-traning underpins models for protein structure (Lin et al., 2023) and function (Samusevich et al., 2024) prediction, as well as for protein design (Hayes et al., 2024). For example, in AlphaFold2, most of the loss function weight is put onto masked modeling of multiple sequence alignments (MSAs) (Jumper et al., 2021), and the model has been effectively fine-tuned for various tasks beyond structure prediction (Jing et al., 2024; Cheng et al., 2023; Motmaen et al., 2023).
Masked modeling is a dominant pre-training objective not only across different tasks but also across various protein representations. Sequence models applied to proteins are typically pre-trained to predict randomly masked amino acids in a random or autoregressive manner (Lin et al., 2023; Rao et al., 2021; Elnaggar et al., 2023; Madani et al., 2023; Ferruz et al., 2022; Rives et al., 2021; Rao et al., 2020). Models utilizing graph neural networks or 3D convolutions on protein structures are also commonly pre-trained to fill in missing structural fragments (Dieckhaus et al., 2024; Diaz et al., 2023; Bushuiev et al., 2023; Hsu et al., 2022; Shroff et al., 2020). The most recent approaches combine both sequential and structural information under masked modeling (Hayes et al., 2024; Su et al., 2023; Heinzinger et al., 2023).
Model adaptation.
In many scenarios, machine learning models for proteins benefit from being adapted to a specific protein of interest. This adaptation is commonly achieved in two ways: either via additional input features or via protein-specific fine-tuning. Multiple sequence alignments (MSAs) containing sequences similar to the target protein provide a common way of supplying a model with protein-specific features (Abramson et al., 2024; Jumper et al., 2021; Rao et al., 2021). Another approach for injecting protein-specific knowledge into the model is standard supervised fine-tuning (i.e., via the $h\circ f$ track) on protein-specific data (Notin et al., 2024; Kirjner et al., 2023; Rao et al., 2019). An alternative is self-supervised fine-tuning (i.e., via the $g\circ f$ track) on proteins from the MSA (Notin et al., 2022b; Frazer et al., 2021; Alley et al., 2019) or on proteins sharing another property with the target protein, such as common family (Sevgen et al., 2023) or class (Samusevich et al., 2024). However, constructing MSAs is time-consuming (Fang et al., 2023), and similar proteins may not be available for many targets (Durairaj et al., 2023; Lin et al., 2023).
Here, we propose an extreme case of self-supervised fine-tuning: learning from a single target protein, without the need for any additional data. To the best of our knowledge, this approach has not been employed in the field of machine learning applied to biology; however, similar methods have been developed in computer vision (Chi et al., 2024; Wang et al., 2023; Xiao et al., 2022; Karani et al., 2021) and natural language processing (Hardt & Sun, 2023; Ben-David et al., 2022; Banerjee et al., 2021). The paradigm of test-time training (TTT), developed to mitigate distribution shifts in computer vision applications (Gandelsman et al., 2022; Sun et al., 2020), is a main inspiration for our work. Here, we demonstrate that TTT is highly relevant for machine learning on proteins even without the presence of explicit distribution shift. We investigate the link of TTT to perplexity minimization and show that TTT improves performance on several downstream tasks.
3 Test-time training (TTT) on proteins
As discussed in the previous section, many machine learning models for proteins employ Y-shaped architectures, consisting of a backbone $f$ with a self-supervised head $g$ and a supervised head $h$. This design facilitates the use of self-supervised fine-tuning across various tasks and models. Notably, most of these models leverage masked modeling as a pre-training objective, which enables the introduction of a broadly applicable test-time training (TTT) method based on masking. Our method adapts models to specific test proteins through masked modeling (Figure 2). In this section, we first formally define the proposed TTT approach (Section 3.1), followed by its application to a range of well-established models (Section 3.2). Finally, we provide an insight into the effectiveness of our method by linking it to perplexity minimization (Section 3.3).
3.1 Self-supervised fine-tuning on test proteins
Figure 2: Overview of our test-time training (TTT) for proteins. Test-time training for proteins builds on the prevalence of Y-shaped architectures relying on masked modeling (i.e., self-supervised masking-based pre-training of $g\circ f$ followed by supervised fine-tuning of $h\circ f$, sharing the backbone $f$). Given a single test protein $x$, TTT adapts the backbone $f$ to the protein using self-supervised fine-tuning. This adaptation leads to better generalization for the downstream task, such as protein fitness, structure, or function prediction.
At test time, we assume a Y-shaped model with a backbone $f$ that has been pre-trained via the self-supervised track $g\circ f$, followed by task-specific fine-tuning through the supervised track $h\circ f$. The goal of test-time training (TTT) is to adapt the backbone $f$ to a single test example $x$ before performing test-time inference on a downstream task via the supervised track.
To achieve this, we first fine-tune the backbone $f$ using the self-supervised track $g\circ f$ on the single example $x$. This step customizes the backbone $f$ to the test sample $x$, and, as demonstrated in Section 4, enhances the generalization of $h\circ f$ without modifying the weights of the task-specific head $h$. Figure 2 illustrates our method. Although the concept of TTT is relatively simple, it involves several important design choices, such as selecting the optimizer and efficiently fine-tuning large backbones, which we describe in the following paragraphs.
Training objective.
We fine-tune $g\circ f$ on a test sample $x$ via minimizing the masked language modeling objective (Devlin, 2018; Rives et al., 2021):
$\mathcal{L}(x)=\mathbb{E}_{M}\Biggl{[}\sum_{i\in M}{-\log{p(x_{i}|x_{\setminus M})}}\Biggl{]},$ (1)
where $x$ denotes a sequence of protein tokens (typically amino acid types), and $\mathbb{E}_{M}$ represents the expectation over randomly sampled masking positions $M$. The loss function $\mathcal{L}(x)$ maximizes the log-probabilities $\log{p(x_{i}|x_{\setminus M})}$ of the true tokens $x_{i}$ at the masked positions $i\in M$ in the partially masked sequence $x_{\setminus M}$. Please note that here we focus on bi-directional masked modeling models, which employ random masking, but the method can be straightforwardly extended to models employing autoregressive masking.
In practice, $\mathbb{E}_{M}$ can follow different distributions, such as sampling a fixed proportion (e.g., 15%) of random amino acid tokens (Lin et al., 2023), or dynamically varying the number of sampled tokens based on another distribution (e.g., a beta distribution) (Hayes et al., 2024). During test-time training, we replicate the masking distribution used during the pre-training. If relevant, we also replicate other pre-training tricks, such as replacing 10% of masked tokens with random tokens and another 10% with the original tokens (Devlin, 2018; Lin et al., 2023; Su et al., 2023) or cropping sequences to random 1024-token fragments (Lin et al., 2023; Su et al., 2023).
Optimization.
We minimize the loss defined in Equation 1 using stochastic gradient descent (SGD) with zero momentum and zero weight decay (Ruder, 2016). While a more straightforward option might be to use the optimizer state from the final pre-training step, this approach is often impractical because the optimizer parameters are usually not provided with the pre-trained model weights (Hayes et al., 2024; Lin et al., 2023). Moreover, many models are pre-trained using the Adam optimizer (Kingma & Ba, 2015) or its variants (Loshchilov & Hutter, 2019). However, it has been shown that Adam results in less predictable behavior of test-time training (TTT) compared to the SGD optimizer, possibly due to its more exploratory behavior (Gandelsman et al., 2022).
Because each TTT experiment assumes only one test example available, we are not able to halt the training using early stopping on any validation sample. Therefore, for each choice of task-specific $f$ and $h$, we tune the optimal number of TTT steps using the entire validation set beforehand or rely on available performance estimates (e.g., pLDDT in the case of protein structure prediction; Section 4.2) to select the optimal number of optimization steps.
Fine-tuning large models.
We aim for test-time training to be applicable on the fly, i.e., without the need for any pre-computation and on a single GPU with a minimum computational overhead. This poses two challenges when using large models. First, when using pre-trained transformers on a single GPU, even for the forward pass, the batch size is typically limited to only several samples due to the quadratic complexity of the inference (Vaswani, 2017). Second, for the backward pass, even a batch size of one is not always feasible for large models. To address the first challenge, we perform forward and backward passes through a small number of training examples and accumulate gradients to simulate updates with any batch size. We address the second challenge by employing low-rank adaptation (LoRA, Hu et al. (2021)), which in practice enables fine-tuning of any model for which a forward pass on a single sample is feasible, due to a low number of trainable parameters.
Figure 3: Test-time training (TTT) improves protein structure prediction by reducing protein sequence perplexity. ESMFold fails to predict the structure of chain B from PDB entry 7EBL in the CAMEO validation set, as shown at TTT step 0, where the perplexity is high and the TM-score is low. By applying TTT on the single target sequence, the model iteratively improves the structure prediction quality, as demonstrated by the increasing TM-score, associated with reduced perplexity. At step 7, the predicted structure achieves the highest TM-score, as well as the highest predicted confidence metric pLDDT, enabling the selection of this step as the final prediction by ESMFold + TTT.
3.2 Inference on downstream tasks
Once the backbone $f$ is adapted to a test protein via self-supervised fine-tuning, it can be used in conjunction with a pre-trained downstream head $h$, as $h\circ f$. The key idea of TTT is not to update the head $h$ during test time, but rather to leverage improved input representations from $f$.
Since Y-shaped architectures are prevalent in protein machine learning, TTT can be straightforwardly applied to numerous tasks in protein research. In this work, we address three primary challenges: protein fitness, structure, and function prediction, applying TTT to corresponding well-established models. For fitness prediction, we apply TTT to ESM2 (Lin et al., 2023) and SaProt (Su et al., 2023); for folding, we apply it to ESMFold (Lin et al., 2023) and ESM3 (Hayes et al., 2024); and for function prediction, we apply TTT to ESM-1v-based (Meier et al., 2021) TerpeneMiner (Samusevich et al., 2024) and ESM-1b-based (Rives et al., 2021) Light attention (Stärk et al., 2021).
In all the models we consider, $f$ is a transformer encoder that takes a protein sequence as input (except for SaProt, which also uses structural tokens), while $g$ is a masked language modeling head (a layer mapping token embeddings to amino acid types). The downstream task heads $h$, however, vary significantly across tasks. For fitness prediction, $h$ outputs a single value for a mutated sequence, measuring how well the protein supports an organism’s functioning. Both ESM2 and SaProt use zero-shot inference from $g\circ f$ via log odds, where $h$ is a simple adaptation of $g$ without any additional parameters. For structure prediction, $h$ is a protein structure decoder: in ESMFold, it is an AlphaFold2-like structure prediction module (Jumper et al., 2021), while in ESM3, it is a VQ-VAE decoder (Razavi et al., 2019). The function predictors are classification models: in TerpeneMiner (Samusevich et al., 2024), $h$ is a random forest that outputs substrate probabilities, and in Light attention (Stärk et al., 2021), $h$ is a light attention module predicting localization class probabilities. Detailed descriptions of the models and their TTT adaptation are provided in Appendix A.
3.3 Link between test-time training and perplexity minimization
[…]
[] Refer to caption
Figure 4: The quality of protein structure prediction, as measured by TM-score, correlates with the perplexity of the underlying language model on the challenging targets from the CAMEO validation set. Higher TM-scores are associated with lower perplexity, indicating that better predictions are linked to lower uncertainty in the language model’s understanding of the protein sequence.
While the approach of test-time training has been extensively investigated in computer vision and other domains, the reasons behind its effectiveness remain unclear (Liu et al., 2021; Zhao et al., 2023). Here, we offer a potential explanation by linking test-time training to perplexity minimization within the context of protein sequence modeling.
Perplexity has traditionally been used in natural language processing to evaluate how well models comprehend test sentences (Brown, 2020; Chelba et al., 2013). Protein language modeling has adopted this metric to assess how effectively models understand amino acid sequences (Hayes et al., 2024; Lin et al., 2023). For bidirectional, random masking language models, which are the focus of this study, we consider the following definition of perplexity  [^2] Please note that this is an approximation of perplexity, which is computationally intractable for bidirectional models, and is often referred to as pseudo-perplexity (Lin et al., 2023; Salazar et al., 2019).:
$\text{Perplexity}(x)=\exp\Biggl{(}\frac{1}{|x|}\sum_{i=1}^{|x|}{-\log{p(x_{i}|x_{\setminus i})}}\Biggl{)},$ (2)
where $|x|$ is the length of the input protein sequence $x$ and $p(x_{i}|x_{\setminus i})$ represents the probability that the model correctly predicts the token $x_{i}$ at position $i$ when it is masked on the input $x_{\setminus i}$. Perplexity ranges from 1 to infinity (the lower the better), providing an intuitive measure of how well a model understands, on average, positions within a given sequence. A perplexity value of 1 indicates that the model perfectly understands the sequence, accurately predicting all the true tokens.
Several studies have shown that lower perplexity on held-out protein sequences (calculated through the self-supervised track $g\circ f$) correlates with better performance on downstream tasks (via the supervised track $h\circ f$), such as predicting protein contacts (Rao et al., 2020), structure (Lin et al., 2023), or fitness (Kantroo et al., 2024). To provide an example, we analyze the correlation between perplexity and structure prediction performance (Figure 4). A strong correlation suggests that reducing a model’s perplexity on a single test sample $x$ can lead to improved performance on the downstream task (Figure 3; Figure 11).
Since we consider only a single test example $x$, the minimization of the masked language modeling loss $\mathcal{L}(x)$ (Equation 1) on this example is directly linked to minimizing the perplexity Perplexity$(x)$ (Equation 2). For instance, in the case of a single masked position (i.e., $|M|=1$), the loss is equal to the logarithm of perplexity. More generally, it can be shown formally that by minimizing the masked language modeling objective, one learns to approximate the conditional marginals of the language (of proteins), including the leave-one-out probabilities evaluated in perplexity (Hennigen & Kim, 2023). As a result, applying test-time training (TTT) through $g\circ f$ enhances the representation of the test protein in the backbone $f$, leading to improved downstream performance via the fine-tuning track $h\circ f$.
4 Experiments
Building on the broad applicability of our test-time training (TTT) approach, we apply it to three example downstream tasks in protein machine learning: fitness, structure, and function prediction. The experimental setup and results for each task are presented in the following subsections.
4.1 Protein Fitness Prediction
Protein fitness refers to the ability of a protein to efficiently perform its biological function, which is determined by its structure, stability, and interactions with other molecules. Predicting protein fitness allows researchers to understand how mutations affect protein function, aiding in protein engineering (Notin et al., 2024). In this paper, we demonstrate that applying test-time training (TTT) to representative models, such as ESM2 (Lin et al., 2023) and SaProt (Su et al., 2023), enhances their protein fitness prediction capabilities. ESM2 is a protein language model trained on protein sequences, while SaProt is an extension of ESM2 that incorporates 3D information via additional structural tokens encoding structures predicted by AlphaFold2 (Jumper et al., 2021).
Evaluation Setup.
We evaluate the models using ProteinGym, the standard benchmark for fitness prediction (Notin et al., 2024), focusing specifically on its zero-shot setup. Since the zero-shot setup only provides a test set without any data split, we aim to validate TTT on independent data. To achieve this, we create a new fitness prediction dataset mined from MaveDB, a public repository containing datasets from Multiplexed Assays of Variant Effect (MAVEs) (Esposito et al., 2019). The quality of the new dataset is validated by confirming that both ESM2 and SaProt generalize well to the new data, achieving comparable performance (Appendix A).
Given a protein and its variants, fitness prediction models output one real value per variant to estimate fitness. ProteinGym uses Spearman correlation between predicted and experimentally measured fitness values as the main evaluation metric for assessing the capabilities of models to score mutations. The correlation is first calculated for each protein and then aggregated per types of measured fitness: activity, binding, expression, organismal fitness, and stability. The final Spearman correlation metric is obtained by averaging across these five categories. We adopt this metric in our benchmarking.
In our evaluation, we also include other top-performing baselines on the ProteinGym benchmark: TraceptEVE (Notin et al., 2022b) and GEMME (Laine et al., 2019). TraceptEVE combines language model Tranception (Notin et al., 2022a) with the protein-specific variational autoencoder, EVE, capturing the evolutionary information via MSAs (Frazer et al., 2021). GEMME is a statistical method deriving fitness predictions from evolutionary trees.
Results.
Test-time training (TTT) consistently enhances the protein fitness prediction performance of both ESM2 and SaProt models across varying model scales (35M and 650M parameters) and both datasets, test ProteinGym (Section 4.1 left) and validation MaveDB (Table 6 in Section B.2). Notably, SaProt (650M) + TTT sets a new state-of-the-art on the ProteinGym benchmark, achieving a 40% higher improvement compared to the previous leaderboard update (SaProt (650M) against TranceptEVE L). When examining performance across different phenotype categories, TTT yields improvements specifically in the categories where the baseline performance is weakest: “Organismal Fitness”, “Binding”, and “Activity” (Section 4.1 right). This improvement indicates the ability of TTT to enhance predictions on challenging targets. Additionally, we observe an inverse correlation between the degree of TTT enhancement and the depth of the MSA (i.e., the number of available homologous sequences) available for each test protein, suggesting that TTT primarily improves predictions for proteins with fewer similar sequences available in the training data (Table 5 in Section B.1).
[…]
[] Avg. Spearman $\uparrow$ Spearman by phenotype $\uparrow$ Activity Binding Expression Organismal Stability Fitness ESM2 (35M) (Lin et al., 2023) 0.3211 0.3137 0.2907 0.3435 0.2184 0.4392 ESM2 (35M) + TTT (Ours) 0.3407 ± 0.00014 0.3407 0.2942 0.3550 0.2403 0.4733 SaProt (35M) (Su et al., 2023) 0.4062 0.3721 0.3568 0.4390 0.2879 0.5749 SaProt (35M) + TTT (Ours) 0.4106 ± 0.00004 0.3783 0.3569 0.4430 0.2955 0.5795 ESM2 (650M) (Lin et al., 2023) 0.4139 0.4254 0.3366 0.4151 0.3691 0.5233 ESM2 (650M) + TTT (Ours) 0.4153 ± 0.00003 0.4323 0.3376 0.4168 0.3702 0.5195 TranceptEVE S* (Notin et al., 2022b) 0.4519 0.4750 0.3957 0.4426 0.4491 0.4973 GEMME* (Laine et al., 2019) 0.4547 0.4820 0.3827 0.4382 0.4517 0.5187 TranceptEVE M* (Notin et al., 2022b) 0.4548 0.4792 0.3858 0.4525 0.4538 0.5025 TranceptEVE L* (Notin et al., 2022b) 0.4559 0.4866 0.3758 0.4574 0.4597 0.5003 SaProt (650M) (Su et al., 2023) 0.4569 0.4584 0.3785 0.4884 0.3670 0.5919 SaProt (650M) + TTT (Ours) 0.4583 ± 0.00001 0.4593 0.3790 0.4883 0.3754 0.5896
Table 1: Test-time training (TTT) improves protein fitness prediction. The right section of the table presents performance averaged across individual proteins and then across different protein phenotypes, as classified in the ProteinGym benchmark (Notin et al., 2024). The middle column shows the final performance, averaged across all five phenotype classes. In total, ProteinGym contains 2.5 million mutations across 217 proteins. Standard deviations are calculated over 5 random seeds and, for brevity, omitted in the right panel, where the maximum standard deviation does not exceed 0.0004. Methods marked with an asterisk (“*”) are the other top-5 methods in ProteinGym, and the metrics are reproduced from the leaderboard (https://proteingym.org/benchmarks).
4.2 Protein structure prediction
Protein structure prediction, also known as protein folding, is the task of predicting 3D coordinates of protein atoms given the amino acid sequence. Protein folding is arguably one of the most successful applications of machine learning in the life sciences (Jumper et al., 2021; Lin et al., 2023; Abramson et al., 2024). However, even state-of-the-art protein folding methods struggle to generalize to entirely novel proteins (Kryshtafovych et al., 2023). In this work, we focus on the ESMFold (Lin et al., 2023) and ESM3 (Hayes et al., 2024) models, demonstrating how their performance on challenging protein structures can be boosted by utilizing test-time training.
Evaluation setup.
To evaluate the performance of TTT, we use CAMEO, a standard benchmark for protein folding. We use the validation and test folds from Lin et al. (2023), focusing only on challenging targets by filtering them according to standard measures of prediction confidence based on pLDDT and perplexity (Section A.2).
Given a protein sequence, the goal of protein folding is to predict 3D coordinates of the protein atoms. To assess the quality of the predicted protein structures with respect to the ground truth structures, we use two standard metrics: TM-score (Zhang & Skolnick, 2004) and LDDT (Mariani et al., 2013). TM-score measures the quality of the global 3D alignment of the target and predicted protein structures, while LDDT is an alignment-free method based on local distance difference tests.
As baseline methods, we use techniques alternative to TTT for improving the performance of the pre-trained base models. In particular, the ESMFold paper proposes randomly masking 15% of amino acids in a protein sequence, allowing for sampling multiple protein structure predictions from the regression ESMFold model (Lin et al., 2023). For each sequence, we sample a number of predictions equal to the total number of TTT steps and refer to this baseline as ESMFold + MP (Masked Prediction). As a baseline for ESM3, we use chain-of-thought iterative decoding, referred to as ESM3 + CoT, proposed in the ESM3 paper (Hayes et al., 2024).
Results.
[…]
[] TM-score $\uparrow$ LDDT $\uparrow$ ESM3 (Hayes et al., 2024) 0.3480 ± 0.0057 0.3723 ± 0.0055 ESM3 + CoT (Hayes et al., 2024) 0.3677 ± 0.0088 0.3835 ± 0.0024 ESM3 + TTT (Ours) 0.3954 ± 0.0067 0.4214 ± 0.0054 ESMFold (Lin et al., 2023) 0.4649 0.5194 ESMFold + MP (Lin et al., 2023) 0.4862 ± 0.0043 0.5375 ± 0.0070 ESMFold + TTT (Ours) 0.5047 ± 0.0132 0.5478 ± 0.0058
Table 2: Test-time training (TTT) improves protein structure prediction. The metrics are averaged across the 18 challenging targets in the CAMEO test set, and standard deviations correspond to 5 random seeds. CoT and MP stand for the chain of though and masked prediction baselines.
Test-time training (TTT) consistently improves the performance of both the ESMFold and ESM3 models, outperforming the masked prediction (ESMFold + MP) and chain-of-thought (ESM3 + CoT) baselines, as shown in Table 2. Of the 18 most challenging CAMEO test proteins, ESMFold and ESM3 significantly improved the prediction of 7 and 6 structures, respectively, while only slightly disrupting the prediction of 2 and 1 structures, respectively (Figure 9 in Section B.1). Most notably, TTT enables accurate structure prediction for targets that are poorly predicted with original base models. For instance, Figure 1 presents a strongly improved structure predicted using ESMFold + TTT for the target that was part of the CASP14 competition and shown as an unsuccessful case in the original ESMFold publication (Lin et al. (2023), Fig. 2E). Another example is shown in Figure 3, where TTT refined the structure prediction from a low-quality prediction (TM-score $=0.29$) to a nearly perfectly folded protein (TM-score $=0.92$).
4.3 Protein function prediction
Protein function prediction is essential for understanding biological processes and guiding bioengineering efforts (Yu et al., 2023; Radivojac & et al., 2013). However, predicting protein function is challenging due to its vague definition and the scarcity of labeled datasets (Stärk et al., 2021; Mikhael et al., 2024; Samusevich et al., 2024). Therefore, enhancing generalization through TTT is particularly relevant in this context. While improved structure prediction through TTT can already enhance function prediction (Song et al., 2024), we further demonstrate that TTT can facilitate function prediction directly. Specifically, we focus on two tasks: subcellular localization, which predicts a protein’s location within a cell (Stärk et al., 2021), and substrate classification for terpene synthases (TPS)—enzymes responsible for the production of terpenoids, the largest class of natural products, and known for their complex biochemistry (Christianson, 2017; Samusevich et al., 2024). Unlike protein fitness or structure prediction, both tasks involve classification (TPS substrate classification being multi-label), providing a complementary evaluation of TTT. We apply TTT to TerpeneMiner (Samusevich et al., 2024) for TPS detection and Light attention (Stärk et al., 2021) for subcellular localization, demonstrating consistent improvements in both tasks.
Evaulation setup.
Figure 5: Test-time training (TTT) enables the correct substrate classification for a terpene synthase (TPS) enzyme. With progressive test-time training steps of TerpeneMiner + TTT, the probability of the initially misclassified substrate (red) decreases, while the probability of the true substrates (green) increases. The bar plots also display the predicted probabilities for other substrates with non-zero values (grey).
For the terpene substrate classification, we use the largest available dataset of characterized TPS from Samusevich et al. (2024) and repurpose the original cross-validation schema. In the case of protein localization prediction, we use a standard DeepLoc dataset (Almagro Armenteros et al., 2017) as a validation set and setHard from (Stärk et al., 2021) as a test set.
Given a protein, the goal of function prediction is to correctly classify it into one of the predefined functional annotations. We assess the quality of the TPS substrate prediction using standard multi-label classification metrics used in the TerpeneMiner paper (Samusevich et al., 2024): mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC). In the case of protein localization prediction, we similarly use the classification metrics from the original paper (Stärk et al., 2021): accuracy, multi-class Matthews correlation coefficient (MCC), and F1-score.
Results.
[…]
[] TPS substrate classification mAP $\uparrow$ AUROC $\uparrow$ TerpeneMiner (Samusevich et al., 2024) 0.805 0.948 TerpeneMiner + TTT (Ours) 0.811 ± 0.0011 0.950 ± 0.0002 Subcellular localization prediction Accuracy $\uparrow$ MCC $\uparrow$ F1-score $\uparrow$ Light attention (Stärk et al., 2021) 0.627 0.549 0.618 Light attention + TTT (Ours) 0.634 ± 0.004 0.557 ± 0.005 0.627 ± 0.004
Table 3: Test-time training (TTT) improves protein function prediction. For the terpene syntase (TPS) substrate classification task, the metrics are computed on the 512 TPS sequences based on the cross-validation schema of the TPS dataset (Samusevich et al., 2024). Subcellular localization prediction performance is reported for 432 protein sequences from the setHard test set (Stärk et al., 2021). The error bars show standard deviations across five random seeds.
TTT improves the performance of the base models on both protein function prediction tasks and across all considered metrics (Table 3). Figure 5 provides a qualitative result, where TTT fine-tuning iteratively refines the prediction of TerpeneMiner toward a correct TPS substrate class.
5 Discussion
In this work, we have developed test-time training (TTT) for proteins, enabling per-protein adaptation of machine learning models for enhanced generalization. TTT improves performance across models, their scales, and benchmarks, while primarily enhancing performance on challenging targets. Our results open up the field of self-supervised adaptation for proteins and provide a proof-of-concept for other biology-related domains. While our method demonstrated strong potential, several underexplored directions remain for future work. In particular, gaining a comprehensive understanding of the success and failure modes of TTT and its application to emerging foundation models could unlock its full potential in computational biology. Additionally, our findings motivate future investigation of broader adaptation frameworks in the context of proteins. Among these frameworks are, for example, domain adaptation, where models leverage both training and test data to adapt to new domains (Ganin & Lempitsky, 2015), or adaptive risk minimization, which employs meta-learning to derive models that can adapt in the face of domain shifts (Zhang et al., 2021).
Reproducibility Statement
Our efforts are focused on ensuring that this research is easily reproducible. The proposed test-time training (TTT) method will be released as a Python package, providing easy-to-use wrappers for the models adapted in this paper. Detailed explanations of the application of TTT to individual models and the construction of datasets are included in the appendix. Where applicable, we will also release the source code for dataset generation.
Acknowledgments
This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through projects e-INFRA CZ [ID:90254], ELIXIR [LM2023055], CETOCOEN Excellence CZ.02.1.01/0.0/0.0/17_043/0009632, ESFRI RECETOX RI LM2023069. This work was also supported by the European Union (ERC project FRONTIER no. 101097822) and the CETOCOEN EXCELLENCE Teaming project supported from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857560. This work was also supported by the Czech Science Foundation (GA CR) grant 21-11563M and by the European Union’s Horizon 2020 research and innovation programme under Marie Skłodowska-Curie grant agreement No. 891397. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
Appendix
In Appendix A, we provide further details on the experimental setup, including comprehensive descriptions of the models, datasets, and metrics used. Next, in Appendix B, we present additional results and their analysis. We discuss the distribution of TTT effects and demonstrate that TTT primarily improves performance on challenging targets. We also explore the impact of hyperparameters by showing the performance on validation sets.
Appendix A Experimental details
In this section, we describe the experimental details for the three downstream tasks considered in this work: protein fitness prediction (Section A.1), protein structure prediction (Section A.2), and protein function prediction (Section A.3). Each subsection describes the application of test-time training (TTT) to the respective models, along with details on the datasets, metrics, and models. Table 4 additionally summarizes the hyperparameters used for the application of TTT to individual models.
A.1 Protein fitness prediction
A.1.1 Datasets
ProteinGym.
ProteinGym is the standard benchmark for protein fitness prediction (Notin et al., 2024). The latest, second version of the dataset includes 217 deep mutation scanning experiments (DMSs) across different proteins. In total, the dataset contains 2.5 mutants with annotated ground-truth fitness. Since ProteinGym does not contain a data split for the zero-shot setup, employed in this work, we use the whole dataset as the test set.
MaveDB dataset.
Figure 6: Comparison of the standard ProteinGym dataset with the MaveDB dataset constructed in this work. A) MaveDB, mined from Esposito et al. (2019), includes novel assays even after filtering to ensure distinct proteins from the comprehensive ProteinGym dataset. This is largely because most MaveDB assays post-filtering date to 2024, whereas the latest assays in ProteinGym date to 2023. B, C, D) MaveDB is of sufficient quality for model evaluation. Representative baselines, ESM2 and SaProt with both 35 million and 650 million parameters, evaluated on ProteinGym generalize effectively to MaveDB, following a similar distribution of predictions. Panel D illustrates the random subset of 50 proteins used for hyperparameter tuning for fitness prediction. Each point in the plots represents one protein and shows the Spearman correlation averaged across all assays corresponding to the protein (typically one assay per protein). The box plots standardly depict quartiles, medians, and outliers.
To establish a validation set disjoint from ProteinGym (Notin et al., 2024), we mined MaveDB [^3] https://www.mavedb.org (Esposito et al., 2019). As of August 1, 2024, the database contains 1178 Multiplexed Assays of Variant Effects (MAVEs), where each assay corresponds to a single protein, measuring the experimental fitness of its variants. We applied quality control filters to remove potentially noisy data. Specifically, we ensured that the UniProt identifier (Consortium, 2023) is valid and has a predicted structure available in the AlphaFold DB (Varadi et al., 2022). We also excluded assays with fewer than 100 variants, as well as those where at least one mutation had a wrongly annotated wild type or where most mutations failed during parsing. Additionally, to ensure no overlap between datasets, we removed any assays whose UniProt identifier matched with those in ProteinGym, ensuring that the validation and test sets contain different proteins.
The described methodology resulted in the MaveDB dataset comprising 676 assays (out of 1178 in the entire MaveDB) with experimental fitness annotations. This corresponds to 483 unique protein sequences and 867 thousand mutations in total. The large size of the dataset, despite the comprehensiveness of ProteinGym containing 217 assays, can be attributed to the fact that many assays in MaveDB were released after the ProteinGym construction (Figure 6A). To ensure the quality of the constructed MaveDB dataset, we validated that representative baselines from ProteinGym generalize to the new assays, following a similar distribution of predictions (Figure 6B,C). Finally, for efficiently tuning hyper-parameters for fitness prediction models we sampled 50 random proteins (Figure 6D), corresponding to 83 assays and collectively 134 thousand variants.
A.1.2 Metrics
Protein fitness labels are not standardized and can vary across different proteins. Nevertheless, the ranking of mutations for a single protein, as defined by fitness labels, can be used to assess the mutation scoring capabilities of machine learning models. As a result, Spearman correlation is a standard metric for evaluation.
Spearman by phenotype.
When computing Spearman correlations, we follow the evaluation protocol proposed in ProteinGym (Notin et al., 2024). First, for each protein, we compute Spearman correlation scores between the predicted ranks of mutations and their corresponding labels. Then, we average the scores across five categories of assayed phenotypes, measuring the effects of introduced mutations: protein catalytic activity (“Activity”), binding affinity to a target (“Binding”), protein expression levels in a cell (“Expression”), organism growth rate (“Organismal Fitness”), and protein thermostability (“Stability”).
Avg. Spearman.
We refer to the mean score across the five phenotype categories as “Avg. Spearman”. We report the “Avg. Spearman” metric as the mean and standard deviation across five random seeds (Section 4.1, Table 5).
Spearman by MSA Depth.
Following (Notin et al., 2024), we split the performance by the depth of available multiple sequence alignment (MSA), i.e., the number of homologous sequences available, as provided in ProteinGym: “Low depth”, “Medium depth”, and “High depth”, and report the Spearman correlation for each subset individually (Table 6). Specifically, the MSA depth categories in ProteinGym are determined using the following thresholds from Hopf et al. (2017): “Low” is defined as $N_{eff}/L<1$, “Medium” as $1001<N_{eff}/L<$ 100, and “High” as $N_{eff}/L>100$, where $N_{eff}$ represents the normalized number of effective sequences in the MSA, and $L$ is the sequence length covered in the MSA.
A.1.3 Models
ESM2.
The ESM2 model is a bidirectional, BERT-like (Devlin, 2018) transformer trained on millions of protein sequences using masked modeling (Lin et al., 2023). The goal of protein fitness prediction is to predict the effects of mutations, and protein language models are often adapted to this task using zero-shot transfer via log odds ratio (Notin et al., 2024; Meier et al., 2021). Specifically, for a given single- or multi-point mutation, where certain amino acids $T$ are substituted from $x_{i}$ to $x^{m}_{i}$ for each $i\in T$, the fitness prediction via the log odds ratio is defined as:
$\sum_{i\in T}{\log{p(x^{m}_{i}|x_{\setminus i})}-\log{p(x_{i}|x_{\setminus i})}},$ (3)
where the sum iterates over mutated positions $i\in T$ with $p(x^{m}_{i}|x_{\setminus i})$ and $p(x_{i}|x_{\setminus i})$ denoting the predicted probabilities of the mutated amino acid and the original one (i.e., wild type), respectively. The conditionals $x_{\setminus i}$ indicate that the input sequence to the model has the position $i$ masked. In this setup, the native (unmutated) sequence, where $T=\emptyset$, has a predicted fitness of 0. Mutations with negative values represent favorable mutations, while positive values correspond to disruptive mutations. We follow the ProteinGym benchmark [^4] https://github.com/OATML-Markslab/ProteinGym and use this formula (Notin et al., 2024) to evaluate the fitness prediction capabilities of ESM2. We use the implementation of ESM2 from ProteinGym.
ESM2 + TTT.
ESM2 can be straightforwardly enhanced with test-time training. Specifically, we treat the transformer encoder as the backbone $f$, and the language modeling head, which projects token embeddings to amino acid probabilities, as the pre-training head $g$. The log odds ratio given by Equation 3 serves as the task-specific head $h$, which in this case involves the pre-training head $g$ that predicts log probabilities. Overall, we apply TTT to the pre-trained ESM2 model and, after a pre-defined number of self-supervised fine-tuning steps, score mutations using Equation 3. During TTT we fine-tune all parameters in $g\circ f$ end-to-end except for token and position embeddings.
SaProt.
We also experiment with the state-of-the-art fitness prediction model, SaProt (Su et al., 2023). SaProt builds off the ESM2 model but incorporates structural information from predicted protein structures. Specifically, SaProt uses the same transformer architecture but expands its vocabulary by combining the 20 standard amino acid tokens with 20 structural tokens from the 3Di vocabulary, increasing the total alphabet size to 400. The 3Di tokens capture the geometry of the protein backbone and are generated using VQ-VAE (Razavi et al., 2019), which projects continuous geometric information into discrete tokens and was trained as part of the Foldseek method (van Kempen et al., 2022).
Since SaProt is also a protein language model, it also uses Equation 3 to score variants. However, please note that SaProt, as implemented in ProteinGym (Notin et al., 2024), uses a slightly different version of the log odds ratio. In SaProt, the conditions in the log probabilities in Equation 3 are replaced with $x_{\setminus T}$ instead of $x_{\setminus i}$, not assuming the independence of substitutions. During TTT, we only mask sequential information and leave the structural part of the tokens unchanged, reflecting the original pre-training setup. We use the implementation of SaProt from ProteinGym4.
SaProt + TTT.
Since the architecture of SaProt is based on ESM2, the TTT components $f$, $g$, and $h$ remain the same. It means that test-time training can be applied to the model in the same way as in the case of ESM2 + TTT discussed above.
A.2 Protein structure prediction
A.2.1 Datasets
CAMEO dataset.
To evaluate the capabilities of TTT on protein folding, we employ the CAMEO validation and test sets as described in Lin et al. (2023). Specifically, the validation set was obtained by querying the CAMEO (Continuous Automated Model Evaluation) web server [^5] https://www.cameo3d.org/modeling (Robin et al., 2021) for entries between August 2021 and January 2022, while the CAMEO test set consists of entries from April 1, 2022, to June 25, 2022. Most of the entries in the CAMEO sets are predicted with high accuracy and confidence (Lin et al., 2023). Therefore, we subselected the challenging validation and test sets where TTT is relevant.
Specifically, we applied two criteria: (1) preserving entries with ESMFold pLDDT scores below 70 to filter out high-confidence predictions (Jumper et al., 2021), and (2) selecting entries with ESM2 perplexity scores greater than or equal to 6, ensuring that the predictions are challenging due to poor sequence understanding rather than other factors. Additionally, most structures with perplexity scores below 6 are already associated with high-confidence predictions (Figure S5 in Lin et al. (2023)). After filtering, the resulting challenging validation and test sets consist of 27 (out of 378) and 18 (out of 194) targets, respectively. The vast majority of the remaining structures have accurate ESMFold structure predictions.
A.2.2 Metrics
To assess the quality of the predicted protein structures with respect to the ground truth structures, we use two standard metrics averaged across the test dataset: TM-score (Zhang & Skolnick, 2004) and LDDT (Mariani et al., 2013).
TM-score.
The TM-score (Template Modeling score) is a metric used to assess the quality of the global 3D alignment between the predicted and target protein structures. It evaluates the structural similarity by comparing the distance between corresponding residues after superposition. The TM-score ranges from 0 to 1, where higher values indicate better alignment.
LDDT.
The Local Distance Difference Test (LDDT) is an alignment-free metric used to assess the accuracy of predicted protein structures. Unlike global metrics, LDDT focuses on local structural differences by measuring the deviation in distances between atom pairs in the predicted structure compared to the target structure. It is particularly useful for evaluating the accuracy of local regions, such as secondary structure elements. LDDT scores range from 0 to 100, with higher values indicating better local structural agreement.
A.2.3 Models
ESMFold.
The ESMFold architecture comprises two key components: a protein language model, ESM2, which, given a protein sequence, generates embeddings for individual amino acids, and a folding block that, using these embeddings and the sequence, predicts the protein 3D structure along with per-amino-acid confidence scores, known as pLDDT scores. In our experiments, we use the esmfold_v0 model from the publicly available ESMFold checkpoints [^6] https://github.com/facebookresearch/esm/blob/main/esm/esmfold/v1/pretrained.py. Please note that we use esmfold_v0 and not esmfold_v1 to avoid data leakage with respect to the CAMEO test set.
ESMFold + TTT.
Since ESM2 backbone of ESMFold was pre-trained in a self-supervised masked modeling regime, the application of TTT to ESMFold is straightforward. We treat ESM2 as the backbone $f$, the language modeling head predicting amino acid classes from their embeddings as the self-supervised head $g$, and the folding trunk along with the structure modules as the downstream task head $h$. After each TTT step, we run $h\circ f$ to compute the pLDDT scores, which allows us to estimate the optimal number of TTT steps for each protein based on the highest pLDDT score.
Since the backbone $f$ is given by the ESM2 model containing 3 billion parameters, we apply LoRA (Hu et al., 2021) to all matrices involved in self-attention. This enables fine-tuning ESMFold + TTT on a single GPU.
ESMFold + ME.
Since ESMFold is a regression model, it only predicts one solution and does not have a straightforward mechanism of sampling multiple structure predictions. Nevertheless, the authors of ESMFold propose a way to sample multiple candidates (Section A.3.2 in Lin et al. (2023)). To sample more solutions, the masking prediction (ME) method randomly masks 15% (same ratio as during masked language modeling pre-training) of the amino acids embeddings before passing them to the structure prediction block. Selecting the solution with the highest pLDDT may lead to improved predicted structure. Since sampling multiple solutions with ESMFold + ME and selecting the best one via pLDDT is analogous to ESMFold + TTT, we employ the former as a baseline, running the method for the same number of step.
ESM3.
Unlike ESMFold, ESM3 is a fully multiple-track, BERT-like model (Devlin, 2018), pre-trained to unmask both protein sequence and structure tokens simultaneously (along with the function tokens). The structure tokens in ESM3 are generated via a separately pre-trained VQ-VAE (Razavi et al., 2019) operating on the protein geometry. In our experiments, we use the smallest, publicly available version of the ESM3 model (ESM3_sm_open_v0) [^7] https://github.com/evolutionaryscale/esm.
ESM3 + TTT.
[…]
[] Refer to caption
Figure 7: Average running time of ESM3 + TTT (50 steps) on the CAMEO validation set using a single NVIDIA A100 GPU.
We treat the transformer encoder of ESM3 as $f$, the language modeling head decoding amino acid classes as $g$, and the VQ-VAE decoder, which maps structure tokens to the 3D protein structure, as $h$. During the TTT steps, we train the model to unmask a protein sequence while keeping the structural track fully padded. During the inference, we provide the model with a protein sequence and run it to unmask the structural tokens, which are subsequently decoded with the VQ-VAE decoder. After each TTT step, we run $h\circ f$ to compute the pLDDT scores, which allows us to estimate the optimal number of TTT steps for each protein based on the highest pLDDT score. We choose the optimal hyperparameters by maximizing the difference in TM-score after and before applying TTT across the validation dataset.
Despite the fact that the model contains 1.4 billion parameters, even without using LoRA, ESM3 + TTT can be fine-tuned on a single NVIDIA A100 GPU. Therefore, we do not employ LoRA for fine-tuning ESM3. Figure 7 shows that ESM3 + TTT can be loaded in VRAM and run for 50 steps in just 10 seconds, with an additional 10 seconds overhead for pLDDT estimation. For comparison, the base ESM3 model can be run on the same hardware in 1 second.
ESM3 + CoT.
To improve the generalization and protein-specific performance of ESM3, the original ESM3 paper employs a chain of thought (CoT) procedure. The procedure unfolds in $n$ steps as follows. At each step, $1/n$ of the masked tokens with the lowest entropy after softmax on logits are unmasked. Then, the partially unmasked sequence is fed back into the model, and the process repeats until the entire sequence is unmasked. In our experiments, we set $n=8$, which is the default value provided in the official GitHub repository.
A.3 Protein function prediction
A.3.1 Datasets
TPS dataset.
For the evaluation of terpene substrate classification, we use the largest available dataset of characterized TPS enzymes from Samusevich et al. (2024) and repurpose the original 5-fold cross-validation schema. We focus on the most challenging TPS sequences, defined as those predicted by the TPS detector, proposed by the dataset authors, with confidence scores below 0.8. This filtering results in 104, 98, 113, 100, 97 examples in the individual folds.
setHard.
For the test evaluation of subcellular location prediction, we use the setHard dataset constructed by Stärk et al. (2021). The dataset was redundancy-reduced, both within itself and relative to all proteins in DeepLoc (Almagro Armenteros et al. (2017); next paragraph), a standard dataset used for training and validating machine learning models. The setHard dataset contains 490 protein sequences, each annotated with one of ten subcellular location classes, such as “Cytoplasm” or “Nucleus”. Since we use ESM-1b (Rives et al., 2021) in our experiments with the dataset, we further filter the data to 432 sequences that do not exceed a length of 1022 amino acids. This step, consistent with Stärk et al. (2021), ensures that ESM-1b can generate embeddings for all proteins.
DeepLoc.
For hyperparameter tuning in the subcellular location prediction task, we use the test set from the DeepLoc dataset (Almagro Armenteros et al., 2017). Similar to setHard, DeepLoc assigns labels from one of ten subcellular location classes. The dataset contains 2768 proteins, which we further filter to 2457 sequences that do not exceed a length of 1022 amino acids, ensuring compatibility with the embedding capabilities of ESM-1b. Since setHard was constructed to be independent of DeepLoc, setHard provides a leakage-free source of data for validation.
A.3.2 Metrics
mAP, AUROC.
The TPS substrate prediction problem is a 12-class multi-label classification task over possible TPS substrates. Therefore, we assess the quality of the predictions using standard multi-label classification metrics such as mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC) averaged across individual classes. These metrics were used in the original TerpeneMiner paper (Samusevich et al., 2024). We report the performance by averaging the metric values concatenated across all validation folds from the 5-fold cross-validation schema.
Accuracy, MCC, F1-score.
To evaluate the performance of subcellular location prediction methods, we use standard classification metrics as employed in Stärk et al. (2021). Accuracy standardly measures the ratio of correctly classified proteins, while Matthew’s correlation coefficient for multiple classes (MCC) serves as an alternative to the Pearson correlation coefficient for classification tasks (Gorodkin, 2004). The F1-score, the harmonic mean of precision and recall, evaluates performance from a retrieval perspective, balancing the trade-off between false positives and false negatives.
A.3.3 Models
TerpeneMiner.
TerpeneMiner is a state-of-the-art method for the classification of terpene synthase (TPS) substrates (Samusevich et al., 2024). The model consists of two parallel tracks. Given a protein sequence, TerpeneMiner first computes its ESM-1v embedding (Meier et al., 2021) and a vector of similarities to the functional domains of proteins from the training dataset, based on unsupervised domain segmentation of AlphaFold2-predicted structures (Jumper et al., 2021). The ESM-1v embedding and the similarity vector are then concatenated and processed by a separately trained random forest, which predicts TPS substrate class probabilities.
In our experiments, we use the “PLM only” version of the model, which leverages only ESM-1v embeddings (PLM stands for protein language model). This version exhibits a minor performance decrease compared to the full model but exactly follows a Y-shaped architecture, allowing us to validate the effectiveness of test-time training for predicting TPS substrates. We use the implementation of TerpeneMiner available at the official GitHub page  [^8] https://github.com/pluskal-lab/TerpeneMiner.
TerpeneMiner + TTT.
When applying TTT to TerpeneMiner, we treat the frozen ESM-1v model as a backbone $f$, its language modeling head as a self-supervised head $g$, and the random forest classifying TPS substrates as a downstream supervised head $h$.
Light Attention.
We use Light attention (Stärk et al., 2021) as a representative baseline for subcellular location prediction. Light attention leverages protein embeddings from a language model, which in our case is ESM-1b (Rives et al., 2021). The model processes per-residue embeddings via a softmax-weighted aggregation mechanism, referred to as light attention, which operates with linear complexity relative to sequence length and enables richer aggregation of per-residue information, as opposed to standard mean pooling. We re-train the model using ESM-1b embeddings on the DeepLoc dataset (Almagro Armenteros et al., 2017) using the code from the official GitHub page [^9] https://github.com/HannesStark/protein-localization.
Light attention + TTT.
When applying TTT to Light attention, we treat the frozen ESM-1b as the backbone $f$, the language modeling head of ESM-1b as the self-supervised head $g$, and the Light attention block as the fine-tuning head $h$.
[…]
[] Learning rate Batch size Grad. acc. steps TTT steps LoRA rank $r$ LoRa $\alpha$ Fitness prediction ESM2 (35M) + TTT 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32, 64 30 - - ESM2 (650M) + TTT 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - - SaProt (35M) + TTT 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - - SaProt (650M) + TTT 4e-5, 4e-4, 4e-3 2, 4 4, 8, 16, 32 30 - - Structure prediction ESMFold + TTT 4e-4 4 4, 8, 32, 64 30 (max pLDDT) 4, 8, 32 8, 16, 32 ESM3 + TTT 1e-4, 4e-4, 1e-3 2 1, 4, 16 50 (max pLDDT) - - Function prediction TerpeneMiner + TTT 4e-4, 1e-3 2 2, 4, 8 30 - - Light attention + TTT 4e-4, 1e-3, 3e-3 2 2, 4 30 - -
Table 4: Hyperparameters used for adapting TTT to individual models. The optimal hyperparameters were estimated using validation datasets corresponding to each of the considered tasks: Fitness prediction, Structure prediction, and Function prediction. Comma-separated lists show the values used for hyperparameter grid search, while the final values selected for computing the test results are highlighted in bold. Low-rank adaptation (LoRA) was only used with ESMFold, containing 3 billion parameters in the ESM2 backbone.
Appendix B Extended results
In this section, we provide additional results on test sets (Section B.1) and discuss validation performance (Section B.2).
B.1 Detailed test performance
In this section, we provide details on the test performance. Specifically, Table 5 shows that test-time training (TTT) primarily enhances performance on challenging targets, characterized by a low number of similar proteins in sequence databases, as measured by MSA depth. Additionally, we provide an example illustrating how TTT substantially improves the correlation between ESM2-predicted fitness and ground-truth stability by better identifying disruptive mutations within the protein core (Figure 8).
Next, Figure 9 shows the distribution of TTT effects: in many cases, TTT has minimal impact on performance; often, it leads to substantial improvements; and in rare cases TTT results in a decrease in performance. This positions TTT as a method for enhancing prediction accuracy, while a comprehensive analysis of its failure modes remains an important direction for future research. While we demonstrate these effects using a protein folding example, we observe a similar distribution of TTT impact across the tasks.
We also observe that the overall trend of TTT generally leads to improved performance, with robust consistency across random seeds. However, the progression of the performance curve can be rugged, particularly in classification tasks, where substantial changes in the underlying representations are required to shift the top-predicted class in the discrete probability distribution (Figure 10).
B.2 Validation performance
This section discusses the performance of test-time training (TTT) on validation data. Table 6 illustrates the validation performance of all tested methods for fitness prediction on our newly constructed MaveDB dataset. TTT enhances the performance of all the methods.
The primary focus of the section is hyperparameter tuning. Table 4 provides the grid of hyperparameters explored for each model and its size. Figure 11 demonstrates the trend of hyperparameter tuning with optimal hyperparameter combination balancing underfitting and overfitting to a single test protein. While most hyperparameter configurations lead to overall improvements when using TTT, poorly chosen hyperparameters can have detrimental effects due to rapid overfitting. However, with a reliable predicted confidence measure, such as pLDDT, the appropriate TTT step can be selected to mitigate overfitting. Figure 12 demonstrates that when using ESM3 + TTT with pLDDT-based step selection for protein folding, all hyperparameter configurations result in improved performance compared to the base ESM3 model.
Figure 8: Example of test-time training (TTT) applied to fitness prediction. Fitness predictions from ESM2 (650M) show poor correlation with experimental fitness values in the ProteinGym test set measured by the stability assay “UBR5_HUMAN_Tsuboyama_2023_1I2T” (Tsuboyama et al., 2023) (left). ESM2 + TTT achieves significantly higher correlation, likely due to improved detection of disruptive mutations in the protein core that impact protein stability (middle). The ground-truth fitness data aligns with the TTT-enhanced model, showing that residues crucial for stability (i.e., having negative mean fitness) are concentrated in the protein core (right). Residue colors represent the mean fitness upon all single-point substitutions (with the exception of several missing mutations in the ground-truth data), with red indicating residues where mutations have detrimental effects on average.
[…]
[] Avg. Spearman $\uparrow$ Spearman by MSA depth $\uparrow$ Low depth Medium depth High depth ESM2 (35M) (Lin et al., 2023) 0.3211 0.2394 0.2707 0.451 ESM2 (35M) + TTT (Ours) 0.3407 ± 0.00014 0.2445 0.3144 0.4598 SaProt (35M) (Su et al., 2023) 0.4062 0.3234 0.3921 0.5057 SaProt (35M) + TTT (Ours) 0.4106 ± 0.00004 0.3253 0.3972 0.5091 ESM2 (650M) (Lin et al., 2023) 0.4139 0.3346 0.4063 0.5153 ESM2 (650M) + TTT (Ours) 0.4153 ± 0.00003 0.3363 0.4126 0.5075 SaProt (650M) (Su et al., 2023) 0.4569 0.3947 0.4502 0.5448 SaProt (650M) + TTT (Ours) 0.4583 ± 0.00001 0.3954 0.4501 0.5439
Table 5: Test-time training (TTT) performance on ProteinGym depending on MSA depth. MSA depth reflects the number of available proteins similar to the target protein and, when using large protein language models, can be interpreted as a measure of the representation of similar proteins in the training data (Section A.1.2). TTT primarily improves performance on difficult targets, with low MSA depth. Standard deviations are calculated over 5 random seeds but are omitted in the right panel for brevity, where the maximum standard deviation does not exceed 0.0004.
[…]
[] Avg. Spearman $\uparrow$ ESM2 (35M) (Lin et al., 2023) 0.4458 ESM2 (35M) + TTT (Ours) 0.4593 ESM2 (650M) (Lin et al., 2023) 0.4568 ESM2 (650M) + TTT (Ours) 0.4604 SaProt (650M) (Su et al., 2023) 0.4926 SaProt (650M) + TTT (Ours) 0.4926 SaProt (35M) (Su et al., 2023) 0.5251 SaProt (35M) + TTT (Ours) 0.5271
Table 6: Performance of test-time training (TTT) on the MaveDB dataset. In this work, we use our newly constructed MaveDB benchmark as a validation fold for tuning the hyper-parameters of TTT for fitness prediction. For computational efficiency, we only select a subset of 50 proteins (Section A.1.1) and do not run TTT across multiple random seeds to estimate standard deviations. The performance shown was calculated by first aggregating correlations per assay, and then per protein (some assays correspond to the same protein).
Figure 9: Per-protein performance of ESMFold + TTT and ESM3 + TTT on the CAMEO test set. The y-axis shows the change in TM-score after applying test-time training (TTT), with higher values indicating improvement. The x-axis represents performance across five random seeds. The red dashed line marks no change in TM-score (TM-score difference $=0$), and the pink band represents minor changes in TM-score ($-0.05<$ TM-score difference $<0.05$), which we do not consider significant. Each point in the swarm plot corresponds to a single protein from the CAMEO test set. On average, applying TTT to ESMFold improves the structure predictions for 7 out of 18 proteins, with 2 showing degradation. The rest of the proteins are not significantly affected. Similarly, applying TTT to ESM3 results in 6 improvements out of 18 proteins, with 1 case of degradation.
Figure 10: Test performance of TerpeneMiner + TTT across fine-tuning steps. The performance is averaged across all 512 proteins in the dataset, with error bars representing the standard deviation across 5 random seeds.
Figure 11: Dependence on hyperparameters in test-time training for fitness prediction. Each plot shows the progression of Spearman correlation (green) increasing alongside a decrease in perplexity (pink) for each TTT step, averaged across all assays in the MaveDB validation dataset. The model used is ESM2 (35M) + TTT and the grid displays the progression for all hyperparameter combinations. As the learning rate increases (left to right) and the effective batch size grows, given by the increased number of gradient accumulation steps (top to bottom), the model reaches peak performance more quickly and begins overfitting to a test protein. The optimal hyperparameter combination (learning rate = 4e-4, gradient accumulation steps = 16) lies near the center of the grid, balancing between underfitting and overfitting to a test protein.
Figure 12: Hyperparameter search for protein structure prediction with ESM3 + TTT. We conducted a comprehensive grid search based on three key hyperparameters: learning rate (denoted as “lr”), number of gradient accumulation steps (denoted as “grad_steps”; with the batch size of two), and masking strategy (denoted as “mask”). We explored two learning rates, 4e-4 and 1e-3, three gradient accumulation step values of 1, 4, and 16, and five different masking strategies: uniform sampling of 0.05, 0.5, and 1.0 fractions of amino acids, as well as the beta30 and betalinear30 distributions proposed in the ESM3 paper (Hayes et al., 2024). Each row in the table presents the mean TM-score and LDDT metrics with standard deviations across five random seeds on the CAMEO validation fold. The last row, denoted as “No TTT”, shows the performance of ESM3 without TTT. The results indicate that ESM3 + TTT is robust to the choice of hyperparameters and consistently outperforms the base model across all configurations. We selected the configuration from the last row to compute the results on the test fold. For the hyperparameter search, we used 30 TTT steps instead of 50 to reduce computation time.

### v2

@@
-Training on test proteins improves fitness, structure, and function prediction
+One protein is all you need
 Abstract
-Data scarcity and distribution shifts often hinder the ability of machine learning models to generalize when applied to proteins and other biological data.
-Self-supervised pre-training on large datasets is a common method to enhance generalization.
-However, striving to perform well on all possible proteins can limit model’s capacity to excel on any specific one, even though practitioners are often most interested in accurate predictions for the individual protein they study.
-To address this limitation, we propose an orthogonal approach to achieve generalization.
-Building on the prevalence of self-supervised pre-training, we introduce a method for self-supervised fine-tuning at test time, allowing models to adapt to the test protein of interest on the fly and without requiring any additional data.
-We study our test-time training (TTT) method through the lens of perplexity minimization and show that it consistently enhances generalization across different models, their scales, and datasets.
-Notably, our method leads to new state-of-the-art results on the standard benchmark for protein fitness prediction, improves protein structure prediction for challenging targets, and enhances function prediction accuracy.
+Generalization beyond training data remains a central challenge in machine learning for biology.
+A common way to enhance generalization is self-supervised pre-training on large datasets.
+However, aiming to perform well on all possible proteins can limit a model’s capacity to excel on any specific one, whereas experimentalists typically need accurate predictions for individual proteins they study, often not covered in training data.
+To address this limitation, we propose a method that enables self-supervised customization of protein language models to one target protein at a time, on the fly, and without assuming any additional data.
+We show that our Protein Test-Time Training (ProteinTTT) method consistently enhances generalization across different models, their sizes, and datasets.
+ProteinTTT improves structure prediction for challenging targets, achieves new state-of-the-art results on protein fitness prediction, and enhances function prediction on two tasks.
+Through two challenging case studies, we also show that customization via ProteinTTT achieves more accurate antibody–antigen loop modeling and enhances 19% of structures in the Big Fantastic Virus Database, delivering improved predictions where general-purpose AlphaFold2 and ESMFold struggle.
 1 Introduction
-[…]
-[] Refer to caption
-Figure 1: Example of test-time training (TTT) applied to protein folding.
-ESMFold poorly predicts the structure of the CASP14 target T1074 (shown in white) because the underlying language model ESM2 poorly fits the sequence, as indicated by the high perplexity (Fig.
+Figure 1: Example of protein structure prediction after single-protein model customization via ProteinTTT.
+ESMFold poorly predicts the structure of the CASP14 target T1074 (white) because the underlying language model ESM2 poorly fits the sequence, as indicated by the high perplexity (left and Fig.
 2E in Lin et al.
-(2023) and the left panel here).
-Self-supervised test-time training of ESM2 on the single sequence of T1074 directly minimizes the perplexity, leading to improved structure prediction (better TM-score alignment and higher pLDDT predicted confidence).
-The same test-time training approach is also broadly applicable to other tasks, such as protein fitness and function prediction.
+(2023)).
+Self-supervised test-time customization of ESM2 to the single sequence of T1074 reduces the perplexity, resulting in improved structure prediction (right).
 A comprehensive understanding of protein structure, function, and fitness is essential for advancing research in the life sciences (Subramaniam & Kleywegt, 2022; Tyers & Mann, 2003; Papkou et al., 2023).
-While machine learning models have demonstrated remarkable potential in protein research, they are typically optimized for achieving the best average performance across large datasets (Jumper et al., 2021; Watson et al., 2023; Yang et al., 2024; Kouba et al., 2023).
-However, biologists often focus their research on individual proteins or protein complexes involved for example in metabolic disorders (Ashcroft et al., 2023; Gunn & Neher, 2023), oncogenic signalling (Hoxhaj & Manning, 2020; Keckesova et al., 2017), neurodegeneration (Gulen et al., 2023; oh Seo et al., 2023), and other biological phenomena (Gu et al., 2022).
+While machine learning models have shown remarkable potential in protein research, they are typically optimized for achieving the best average performance across large datasets (Jumper et al., 2021; Watson et al., 2023; Kouba et al., 2023).
+However, biologists often focus their research on individual proteins or protein complexes involved in, for example, metabolic disorders (Ashcroft et al., 2023; Gunn & Neher, 2023), oncogenic signaling (Hoxhaj & Manning, 2020; Keckesova et al., 2017), neurodegeneration (Gulen et al., 2023; oh Seo et al., 2023), and other biological phenomena (Gu et al., 2022).
 In these scenarios, detailed insights into a single protein can lead to significant scientific advances.
-Nonetheless, general machine learning models for proteins often struggle to generalize to individual case studies due to data scarcity (Bushuiev et al., 2023; Chen & Gong, 2022) and distribution shifts (Tagasovska et al., 2024; Feng et al., 2024).
-Bridging the gap between broad, dataset-wide optimizations and the precision required for studying single proteins in practical applications remains a critical challenge in integrating machine learning into biological research (Sapoval et al., 2022).
-By contrast, in other application domains of machine learning, such as computer vision and natural language processing, customization and adaptation approaches have emerged as powerful tools to improve model performance in specific contexts (Ruiz et al., 2023; Hardt & Sun, 2023).
-Drawing inspiration from the test-time training (TTT) approach developed in computer vision to mitigate distribution shifts (Sun et al., 2020; Gandelsman et al., 2022), in this work we propose the TTT approach for proteins.
-Our method enables adapting protein models to one protein at a time, on the fly, and without the need for additional data.
-Given a model that has been pre-trained using masked language modeling, our method minimizes the perplexity of the model on a given test protein through self-supervised fine-tuning, which, in turn, results in improved downstream performance without updating the downstream task head.
-The prevalence of masked modeling in protein machine learning makes our method broadly applicable to various downstream tasks.
-Empirically, we demonstrate its effectiveness across three key challenges in protein machine learning.
-First, TTT achieves state-of-the-art results on the ProteinGym dataset (Notin et al., 2024), a well-established benchmark for protein fitness prediction.
-Second, TTT enhances protein structure predictions with ESMFold (Lin et al., 2023) and ESM3 (Hayes et al., 2024) on challenging targets.
-Third, the application of TTT to protein function predictors results in improved classification of terpene synthase (TPS) substrates and protein subcellular localization.
-In summary, the key contributions of this work are three-fold:
-  1. 1.
-    Motivated by the generalization challenges and distribution shifts prevalent in protein machine learning, we introduce a new test-time training (TTT) method [^1] https://github.com/anton-bushuiev/ProteinTTT that enables models to adapt to individual proteins on the fly and without requiring additional data.
-  2. 2.
-    We establish a link between our TTT approach and perplexity minimization, providing an insight into why this approach enhances model effectiveness.
-  3. 3.
-    We empirically validate TTT, achieving state-of-the-art results in protein fitness prediction, improving the protein structure prediction capabilities of well-established folding models, and enhancing protein function predictions in the tasks of terpene synthase substrate classification and protein localization prediction.
+However, general machine learning models for proteins often struggle to generalize to practically interesting individual cases due to data scarcity (Bushuiev et al., 2023; Chen & Gong, 2022) and distribution shifts (Škrinjar et al., 2025; Tagasovska et al., 2024; Feng et al., 2024).
+Bridging the gap between broad, dataset-wide optimization and precision needed to study single proteins of practical interest remains a key challenge in integrating machine learning into biological research (Sapoval et al., 2022).
+This challenge is particularly acute in computational biology, where accurate predictions for individual proteins are essential to guide resource-intensive wet-lab experiments, in contrast to domains such as natural language processing or computer vision, where models are typically expected to flexibly handle diverse prompts from many users in real time (Brown, 2020; Ramesh et al., 2021).
+To address this challenge, we propose a test-time approach for generalization to one protein at a time, effectively enabling more accurate predictions for individual targets, particularly those poorly represented in training data.
+Our Protein Test-Time Training (ProteinTTT) method customizes protein language models (PLMs) to individual proteins on the fly and without assuming additional data.
+Our approach is based on a simple yet powerful premise: if a language model is less perplexed (surprised) by a protein sequence–or if it “understands” its unique patterns better–it will generate a more accurate representation for predicting its structure and function.
+Given a model pre-trained via masked language modeling, our method effectively minimizes perplexity on a target protein or its multiple sequence alignment (MSA) through self-supervised customization, improving downstream performance without updating the downstream task head.
+The widespread use of masked modeling as a pre-training paradigm makes ProteinTTT broadly applicable in computational biology.
+In summary, this work demonstrates the surprising effectiveness of protein model customization and lays the foundation for exploring other test-time strategies and broader biological applications.
+The key contributions are: (1) We introduce ProteinTTT, to the best of our knowledge the first customization method in machine learning for biology.
+We provide a user-friendly and easily extensible implementation  [^1] https://github.com/anton-bushuiev/ProteinTTT and provide insights into the effectiveness of protein model customization by linking it to perplexity minimization.
+(2) We empirically validate ProteinTTT, showing improvements in protein structure prediction with well-established models, achieving state-of-the-art results in protein fitness prediction, and enhancing protein function prediction on terpene synthase substrate classification and protein localization prediction.
+(3) We demonstrate the practical utility of focusing on one protein at a time through two challenging case studies.
+ProteinTTT enables more accurate prediction of antibody–antigen loops and improves 19% of structures in the Big Fantastic Virus Database, delivering accurate predictions where general-purpose AlphaFold2 and ESMFold struggle.
 2 Background and related work
-In this section, we present the context and related work that highlight the feasibility and broad applicability of test-time training (TTT) in the domain of machine learning on proteins.
-The widespread adoption of Y-shaped architectures relying on masked modeling enables the development of a general method for adapting protein models at test time via masking-based self-supervised fine-tuning.
+The broad adoption of Y-shaped architectures relying on masked modeling enables the development of a general method for customizing protein models at test time via masking-based self-supervision.
 The Y-shaped paradigm of learning.
-In machine learning applied to biology, architectures often follow a Y-shaped paradigm (Gandelsman et al., 2022), consisting of a backbone feature extractor $f$, a self-supervised head $g$, and an alternative fine-tuning head $h$.
-During training, $g\circ f$ is first pre-trained, after which the pre-trained backbone $f$ is reused to fine-tune $h\circ f$ toward a downstream task.
+In machine learning applied to proteins, architectures often follow a Y-shaped paradigm (Gandelsman et al., 2022), consisting of a backbone feature extractor $f$ operating on protein tokens $x$, a self-supervised head $g$, and an alternative fine-tuning head $h$.
+During training, $g\circ f$ is first pre-trained, and the pre-trained backbone $f$ is then reused to fine-tune $h\circ f$ toward a downstream task.
 Here, $\circ$ denotes a composition of two machine learning modules (e.g., $g$ is applied on top of $f$ in $g\circ f$).
 At test time, the final model $h\circ f$ is fixed.
-Generalization is achieved by leveraging the rich knowledge encoded in the backbone $f$ and the task-specific priors acquired in the fine-tuning head $h$.
-This paradigm enables overcoming data scarcity during fine-tuning and underlies breakthrough approaches in protein structure prediction (Lin et al., 2023), protein design (Watson et al., 2023), protein function prediction (Yu et al., 2023), and other protein-related tasks (Hayes et al., 2024).
+Generalization is achieved by leveraging the rich knowledge encoded in the backbone $f$ and the task-specific priors embedded in the fine-tuning head $h$.
+This paradigm enables overcoming data scarcity during fine-tuning and underlies breakthrough approaches in protein structure prediction (Lin et al., 2023), protein design (Watson et al., 2023), protein function prediction (Yu et al., 2023), and other tasks (Hayes et al., 2024).
 The backbone $f$ is typically a large neural network pre-trained in a self-supervised way on a large dataset using a smaller pre-training projection head $g$ (Hayes et al., 2024).
 The fine-tuning head $h$, however, depends on the application.
-In some cases, $h$ is a large neural network, repurposing the pre-trained model entirely (Watson et al., 2023; Lin et al., 2023); in others, $h$ is a minimal projection with few parameters (Cheng et al., 2023), or even without any parameters at all (i.e., a zero-shot setup, (Meier et al., 2021; Dutton et al., 2024)).
-In some cases, the fine-tuning head $h$ may also be a machine learning algorithm other than a neural network (Samusevich et al., 2024).
+In some cases, $h$ is a large neural network, repurposing the pre-trained model entirely (Watson et al., 2023; Lin et al., 2023); in others, $h$ is a minimal projection with few parameters (Cheng et al., 2023), or even without any parameters at all (i.e., a zero-shot setup; Meier et al.
+(2021); Dutton et al.
+(2024)).
+The fine-tuning head $h$ can also be a machine learning algorithm other than a neural network (Samusevich et al., 2025).
 Masked modeling.
 While the objective of fine-tuning $h\circ f$ is determined by the downstream application, the choice of pre-training objective for $g\circ f$ is less straightforward.
-Nevertheless, most methods employ various forms of masked modeling, i.e., optimizing the model weights to accurately reconstruct missing parts of proteins, regardless of the downstream application.
-Masked modeling pre-traning underpins models for protein structure (Lin et al., 2023) and function (Samusevich et al., 2024) prediction, as well as for protein design (Hayes et al., 2024).
-For example, in AlphaFold2, most of the loss function weight is put onto masked modeling of multiple sequence alignments (MSAs) (Jumper et al., 2021), and the model has been effectively fine-tuned for various tasks beyond structure prediction (Jing et al., 2024; Cheng et al., 2023; Motmaen et al., 2023).
-Masked modeling is a dominant pre-training objective not only across different tasks but also across various protein representations.
-Sequence models applied to proteins are typically pre-trained to predict randomly masked amino acids in a random or autoregressive manner (Lin et al., 2023; Rao et al., 2021; Elnaggar et al., 2023; Madani et al., 2023; Ferruz et al., 2022; Rives et al., 2021; Rao et al., 2020).
-Models utilizing graph neural networks or 3D convolutions on protein structures are also commonly pre-trained to fill in missing structural fragments (Dieckhaus et al., 2024; Diaz et al., 2023; Bushuiev et al., 2023; Hsu et al., 2022; Shroff et al., 2020).
-The most recent approaches combine both sequential and structural information under masked modeling (Hayes et al., 2024; Su et al., 2023; Heinzinger et al., 2023).
-Model adaptation.
-In many scenarios, machine learning models for proteins benefit from being adapted to a specific protein of interest.
-This adaptation is commonly achieved in two ways: either via additional input features or via protein-specific fine-tuning.
-Multiple sequence alignments (MSAs) containing sequences similar to the target protein provide a common way of supplying a model with protein-specific features (Abramson et al., 2024; Jumper et al., 2021; Rao et al., 2021).
-Another approach for injecting protein-specific knowledge into the model is standard supervised fine-tuning (i.e., via the $h\circ f$ track) on protein-specific data (Notin et al., 2024; Kirjner et al., 2023; Rao et al., 2019).
-An alternative is self-supervised fine-tuning (i.e., via the $g\circ f$ track) on proteins from the MSA (Notin et al., 2022b; Frazer et al., 2021; Alley et al., 2019) or on proteins sharing another property with the target protein, such as common family (Sevgen et al., 2023) or class (Samusevich et al., 2024).
-However, constructing MSAs is time-consuming (Fang et al., 2023), and similar proteins may not be available for many targets (Durairaj et al., 2023; Lin et al., 2023).
-Here, we propose an extreme case of self-supervised fine-tuning: learning from a single target protein, without the need for any additional data.
-To the best of our knowledge, this approach has not been employed in the field of machine learning applied to biology; however, similar methods have been developed in computer vision (Chi et al., 2024; Wang et al., 2023; Xiao et al., 2022; Karani et al., 2021) and natural language processing (Hardt & Sun, 2023; Ben-David et al., 2022; Banerjee et al., 2021).
-The paradigm of test-time training (TTT), developed to mitigate distribution shifts in computer vision applications (Gandelsman et al., 2022; Sun et al., 2020), is a main inspiration for our work.
-Here, we demonstrate that TTT is highly relevant for machine learning on proteins even without the presence of explicit distribution shift.
-We investigate the link of TTT to perplexity minimization and show that TTT improves performance on several downstream tasks.
-3 Test-time training (TTT) on proteins
-As discussed in the previous section, many machine learning models for proteins employ Y-shaped architectures, consisting of a backbone $f$ with a self-supervised head $g$ and a supervised head $h$.
-This design facilitates the use of self-supervised fine-tuning across various tasks and models.
-Notably, most of these models leverage masked modeling as a pre-training objective, which enables the introduction of a broadly applicable test-time training (TTT) method based on masking.
-Our method adapts models to specific test proteins through masked modeling (Figure 2).
-In this section, we first formally define the proposed TTT approach (Section 3.1), followed by its application to a range of well-established models (Section 3.2).
-Finally, we provide an insight into the effectiveness of our method by linking it to perplexity minimization (Section 3.3).
-3.1 Self-supervised fine-tuning on test proteins
-Figure 2: Overview of our test-time training (TTT) for proteins.
-Test-time training for proteins builds on the prevalence of Y-shaped architectures relying on masked modeling (i.e., self-supervised masking-based pre-training of $g\circ f$ followed by supervised fine-tuning of $h\circ f$, sharing the backbone $f$).
-Given a single test protein $x$, TTT adapts the backbone $f$ to the protein using self-supervised fine-tuning.
-This adaptation leads to better generalization for the downstream task, such as protein fitness, structure, or function prediction.
+Nevertheless, the dominant paradigm for protein pre-training is masked modeling, which optimizes model weights to reconstruct missing protein parts.
+This objective has proven effective across diverse tasks (Heinzinger & Rost, 2025; Schmirler et al., 2024), including structure (Lin et al., 2023; Jumper et al., 2021), fitness (Meier et al., 2021; Su et al., 2023), and function prediction (Samusevich et al., 2025; Yu et al., 2023; Elnaggar et al., 2021), as well as protein design (Hayes et al., 2024), and has been successfully applied to various protein representations such as sequences (Hayes et al., 2024; Elnaggar et al., 2023), graphs (Dieckhaus et al., 2024; Bushuiev et al., 2023), and voxels (Diaz et al., 2023).
+Model customization.
+Several studies have shown that machine learning models for proteins benefit from being fine-tuned on protein-specific  (Notin et al., 2024; Kirjner et al., 2023; Rao et al., 2019) or protein family-specific (Sevgen et al., 2023; Samusevich et al., 2025) data.
+However, collecting additional data may be resource-intensive, and for many targets, relevant datasets or proteins may be limited or not available (Durairaj et al., 2023; Kim et al., 2025).
+In this paper, we propose a versatile method enabling customizing PLMs for a single target protein or its MSA in a self-supervised manner, on the fly, and without assuming any additional data.
+Customization methods have been developed in computer vision (Chi et al., 2024; Wang et al., 2023; Xiao et al., 2022; Karani et al., 2021) and natural language processing (Hübotter et al., 2024; Hardt & Sun, 2023; Ben-David et al., 2022; Banerjee et al., 2021).
+The paradigm of test-time training (TTT), developed to mitigate distribution shifts in computer vision applications (Gandelsman et al., 2022; Sun et al., 2020), is the main inspiration for our work.
+We demonstrate that customization via test-time training enhances the accuracy of PLMs across a wide range of downstream tasks even without the presence of explicit distribution shifts.
+3 Protein model customization with ProteinTTT
+In this section, we describe the proposed Protein Test-Time Training (ProteinTTT) approach (Section˜3.1), followed by its applications to a range of well-established models and datasets (Section˜3.2).
+3.1 Self-supervised customization to a target protein
 At test time, we assume a Y-shaped model with a backbone $f$ that has been pre-trained via the self-supervised track $g\circ f$, followed by task-specific fine-tuning through the supervised track $h\circ f$.
-The goal of test-time training (TTT) is to adapt the backbone $f$ to a single test example $x$ before performing test-time inference on a downstream task via the supervised track.
-To achieve this, we first fine-tune the backbone $f$ using the self-supervised track $g\circ f$ on the single example $x$.
-This step customizes the backbone $f$ to the test sample $x$, and, as demonstrated in Section 4, enhances the generalization of $h\circ f$ without modifying the weights of the task-specific head $h$.
-Figure 2 illustrates our method.
-Although the concept of TTT is relatively simple, it involves several important design choices, such as selecting the optimizer and efficiently fine-tuning large backbones, which we describe in the following paragraphs.
-Training objective.
-We fine-tune $g\circ f$ on a test sample $x$ via minimizing the masked language modeling objective (Devlin, 2018; Rives et al., 2021):
-$\mathcal{L}(x)=\mathbb{E}_{M}\Biggl{[}\sum_{i\in M}{-\log{p(x_{i}|x_{\setminus M})}}\Biggl{]},$ (1)
+The goal of customization with ProteinTTT is to adapt the backbone $f$ to a single protein $x$ before making a prediction on a downstream task via the supervised track $h\circ f$.
+To achieve this, we customize the backbone $f$ to the single example $x$:
+$\text{ProteinTTT}:(h\circ f(\cdot;\theta_{0}),x)\mapsto h\circ f(\cdot;\theta_{x})$ (1)
+where $\theta_{0}$ denotes pre-trained parameters and $\theta_{x}$ parameters optimized for the target protein $x$ using the self-supervised track $g\circ f$, while the supervised head $h$ remains frozen.
+Figure˜2a illustrates our customization approach, which is summarized in the following sections.
+Appendix˜B describes the extension of our method to customization using a MSA of a protein, rather than its single sequence.
+Customization training objective.
+We customize $g\circ f$ to a single target protein sequence $x$ via minimizing the masked language modeling objective (Devlin, 2018; Rives et al., 2021):
+$\mathcal{L}(x;\theta)=\mathbb{E}_{M\sim p_{\text{mask}}(M)}\Biggl[\sum_{i\in M}{-\log{p(x_{i}|x_{\setminus M};\theta)}}\Biggl],$ (2)
 where $x$ denotes a sequence of protein tokens (typically amino acid types), and $\mathbb{E}_{M}$ represents the expectation over randomly sampled masking positions $M$.
-The loss function $\mathcal{L}(x)$ maximizes the log-probabilities $\log{p(x_{i}|x_{\setminus M})}$ of the true tokens $x_{i}$ at the masked positions $i\in M$ in the partially masked sequence $x_{\setminus M}$.
-Please note that here we focus on bi-directional masked modeling models, which employ random masking, but the method can be straightforwardly extended to models employing autoregressive masking.
-In practice, $\mathbb{E}_{M}$ can follow different distributions, such as sampling a fixed proportion (e.g., 15%) of random amino acid tokens (Lin et al., 2023), or dynamically varying the number of sampled tokens based on another distribution (e.g., a beta distribution) (Hayes et al., 2024).
-During test-time training, we replicate the masking distribution used during the pre-training.
-If relevant, we also replicate other pre-training tricks, such as replacing 10% of masked tokens with random tokens and another 10% with the original tokens (Devlin, 2018; Lin et al., 2023; Su et al., 2023) or cropping sequences to random 1024-token fragments (Lin et al., 2023; Su et al., 2023).
+The objective function $\mathcal{L}(x;\theta)$ maximizes the log-probabilities $\log{p(x_{i}|x_{\setminus M};\theta)}\ \dot{=}\ g(f(x_{\setminus M};\theta))_{i}$ of the true (i.e., wild-type) tokens $x_{i}$ at the masked positions $i\in M$ in the partially masked sequence $x_{\setminus M}$, where $\theta$ denotes the parameters of the backbone $f$, and $g$ is the masked language modeling head.
+Please note that here we focus on bi-directional masked modeling models, which employ random masking, but the method can be easily extended to models employing autoregressive masking.
+To ensure consistency between the customization and pre-training, ProteinTTT adopts the same masking and data preprocessing strategies used during pre-training.
+Specifically, $p_{\text{mask}}(M)$ can follow different distributions, such as sampling a fixed proportion (e.g., 15%) of random amino acid tokens (Lin et al., 2023), or dynamically varying the number of sampled tokens based on another distribution (e.g., a beta distribution; Hayes et al.
+(2024)).
+During the customization, we replicate the masking distribution used during the pre-training.
+We also replicate other pre-training practices, such as replacing 10% of masked tokens with random tokens and another 10% with the original tokens (Devlin, 2018; Lin et al., 2023; Su et al., 2023) or cropping sequences to random 1024-token fragments (Lin et al., 2023; Su et al., 2023).
 Optimization.
-We minimize the loss defined in Equation 1 using stochastic gradient descent (SGD) with zero momentum and zero weight decay (Ruder, 2016).
-While a more straightforward option might be to use the optimizer state from the final pre-training step, this approach is often impractical because the optimizer parameters are usually not provided with the pre-trained model weights (Hayes et al., 2024; Lin et al., 2023).
+Figure 2: Overview of protein language model (PLM) customization with ProteinTTT.
+(a) Given a protein sequence of interest $x$ and a pretrained PLM $f(\cdot;\theta_{0})$, ProteinTTT yields a customized version of the PLM $f(\cdot;\theta_{x})$ for that sequence.
+Customization is achieved by fine-tuning (fire icon) the pretrained parameters $\theta_{0}$ via masked language modeling solely on the input sequence for $T$ steps, selecting the optimal parameters $\theta_{x}$ using a confidence function $c$.
+This procedure adapts the model specifically to the input sequence, improving its internal representation as measured by model perplexity.
+(b) Once customized, the PLM can be used with pretrained task-specific heads, such as structure, fitness, or function prediction modules, $h_{1}$, $h_{2}$, and $h_{3}$, respectively, without modifying their parameters (snowflake icon).
+For example, the ESM2 PLM can be customized and then used with the pretrained ESMFold structure prediction head without modifying its 1.4-billion task-specific parameters, resulting in improved structure prediction for the given sequence (e.g., Figure˜1).
+Since customization with ProteinTTT does not assume more than a single protein, early stopping on validation data is not feasible.
+To address this, we first fine-tune the pre-trained parameters $\theta_{0}$ of a backbone $f$ for a fixed number of steps $T$, yielding parameters $\Theta=\{\theta_{0},\theta_{1},\dots,\theta_{T}\}$.
+The final customized parameters $\theta_{x}$ are selected as $\operatorname*{arg\,max}_{\theta\in\Theta}c(h(f(x;\theta)))$ where $c$ is a confidence function.
+If $c$ is not available, we set $\theta_{x}=\theta_{T}$.
+Section˜G.2 discusses how using pLDDT as the confidence function $c$ for structure prediction makes ProteinTTT robust to hyperparameter selection and how the number of steps $T$ can be fixed (e.g., $T=30$) while optimizing learning rate and batch size effectively.
+Before customizing for the next target protein, the parameters are reset to $\theta_{0}$.
+To make ProteinTTT easily applicable to large-scale models (e.g., the 3B-parameter ESM2 backbone), we leverage low-rank adaptation (LoRA; Hu et al.
+(2021)) and gradient accumulation during customization.
+Additionally, to improve the stability and predictability of customization, we use stochastic gradient descent (SGD; Ruder (2016)) instead of the commonly used Adam optimizer (Kingma & Ba, 2015), following (Gandelsman et al., 2022).
+Further details are provided in Appendix˜E.
+Figure 3: Customization with ProteinTTT improves protein structure prediction by reducing protein sequence perplexity.
+ESMFold fails to predict the structure of chain B from PDB entry 7EBL in the CAMEO validation set, as shown at customization step 0, where the perplexity is high and the TM-score is low.
+By applying customization with ProteinTTT for the single target sequence, the model iteratively improves the structure prediction quality, as demonstrated by the increasing TM-score, associated with reduced perplexity.
+At customization step 7, the predicted structure achieves the highest TM-score, as well as the highest predicted confidence metric pLDDT, enabling the selection of this step as the final prediction by the customized ESMFold + ProteinTTT.
+3.2 Inference on downstream tasks
+Once the backbone $f$ is adapted to a target protein via self-supervised customization, it can be used in conjunction with a pre-trained downstream head $h$, as $h\circ f$.
+The key idea of customization with ProteinTTT is not to update the head $h$, but instead to leverage improved representations from $f$ (Figure˜2b).
+Appendix˜A provides a justification for why these customized representations generally enhance performance on downstream tasks by linking ProteinTTT to perplexity minimization.
+Since Y-shaped architectures are prevalent in protein machine learning, ProteinTTT can be straightforwardly applied to numerous tasks.
+In this work, we consider three standard problems: protein structure, fitness, and function prediction, and apply our method to corresponding well-established models.
+For structure prediction, we apply ProteinTTT to ESMFold (Figure˜3, Lin et al.
+(2023), HelixFold-Single (Fang et al., 2023), and ESM3 (Hayes et al., 2024); for fitness prediction, we use ESM2 (Lin et al., 2023), SaProt (Su et al., 2023), ProSST (Li et al., 2024), and MSA Transformer (Rao et al., 2021); and for function prediction, we apply ProteinTTT to ESM-1v-based (Meier et al., 2021) EnzymeExplorer (Samusevich et al., 2025) and ESM-1b-based (Rives et al., 2021) Light attention (Stärk et al., 2021).
+In all models we consider, $f$ is a Transformer encoder that takes protein tokens as input, and $g$ is a masked language modeling head (a layer mapping token embeddings to amino acid types).
+The downstream task heads $h$ vary strongly across tasks.
+For structure prediction, $h$ is a protein structure predictor: in ESMFold and HelixFold-Single, it is an AlphaFold2-inspired module (Jumper et al., 2021), while in ESM3, it is a VQ-VAE structure decoder (Razavi et al., 2019).
+For fitness prediction, $h$ outputs a single score per sequence; ESM2, SaProt, and ProSST perform zero-shot inference using $h\circ f$ via log odds from $g$, with $h$ functioning as a simple adaptation of $g$ without introducing extra parameters.
+The function predictors are classification models: in EnzymeExplorer (Samusevich et al., 2025), $h$ is a random forest that outputs substrate probabilities, and in Light attention (Stärk et al., 2021), $h$ is a light attention module predicting protein localization classes within a cell.
+4 Experiments
+In this section, we evaluate ProteinTTT on three well-established downstream tasks in protein machine learning: structure (Section˜4.1), fitness (Section˜4.2), and function (Section˜4.3) prediction.
+4.1 Protein structure prediction
+[…]
+[…]
+Protein structure prediction is the task of predicting 3D coordinates of protein atoms given the amino acid sequence.
+It is arguably one of the best-established problems in computational biology (Jumper et al., 2021; Lin et al., 2023; Abramson et al., 2024).
+Evaluation setup.
+To evaluate the performance of ProteinTTT, we employ CAMEO, a standard benchmark for protein folding.
+We use the validation and test folds from Lin et al.
+(2023), focusing only on targets with low-confidence predictions from the base ESMFold, as determined by pLDDT and perplexity (Section˜E.1).
+We use the standard TM-score (Zhang & Skolnick, 2004) and LDDT (Mariani et al., 2013) metrics to evaluate global and local structure prediction quality, respectively.
+As baseline methods, we use techniques alternative to ProteinTTT for improving the performance of the pre-trained base models.
+In particular, the ESMFold paper proposes randomly masking 15% of amino acids in a protein sequence before inference, allowing for sampling multiple protein structure predictions from the regression ESMFold model (Lin et al., 2023).
+For each sequence, we sample a number of predictions equal to the total number of ProteinTTT steps and refer to this baseline as ESMFold + MP (Masked Prediction).
+As a baseline for ESM3, we use chain-of-thought iterative decoding, referred to as ESM3 + CoT, proposed in the ESM3 paper (Hayes et al., 2024).
+Results.
+Customization with ProteinTTT consistently improves the performance of all the tested methods, ESMFold, HelixFold-Single, and ESM3, outperforming the masked prediction (ESMFold + MP) and chain-of-thought (ESM3 + CoT) baselines, as shown in Table˜2.
+Among the 18 challenging CAMEO test proteins, ProteinTTT significantly improved the prediction of 7, 5, and 6 structures from ESMFold, HelixFold-Single, and ESM3, respectively, while only slightly disrupting the prediction of 2, 1, and 1 structures, respectively (Figure˜A6).
+Most notably, ProteinTTT enables accurate structure prediction for targets that are poorly predicted with the original models.
+For instance, Figure˜1 presents a strongly improved structure predicted using ESMFold + ProteinTTT for the target that was part of the CASP14 competition and shown as an unsuccessful case in the original ESMFold publication (Lin et al.
+(2023), Fig.
+2E).
+Another example is shown in Figure˜3, where ProteinTTT refined the structure prediction from a low-quality prediction (TM-score =0.29=0.29) to a nearly perfectly folded protein (TM-score =0.92=0.92).
+Figure˜A4 shows that ESMFold + ProteinTTT maintains computational efficiency of ESMFold, being an order of magnitude faster than AlphaFold2.
+Figure˜A11 additionally demonstrates the robustness of ESM3 + ProteinTTT to the choice of hyperparameters.
+4.2 Protein Fitness Prediction
+[…]
+[…]
+The task of protein fitness prediction is to accurately order mutations of a protein based on their disruptive/favorable effects on protein functioning.
+Evaluation Setup.
+We evaluate the models using ProteinGym, the state-of-the-art fitness prediction benchmark (Notin et al., 2024), focusing on its well-established zero-shot setup.
+Since the zero-shot setup only provides a test set without any data split, we also validate ProteinTTT on independent data.
+To achieve this, we create a new fitness prediction dataset mined from MaveDB, a public repository containing datasets from Multiplexed Assays of Variant Effect (MAVEs) (Esposito et al., 2019).
+Following ProteinGym, we measure performance on both datasets using Spearman correlation between predicted and experimental fitness values.
+Results.
+ProteinTTTconsistently enhances fitness prediction performance of all the tested models across varying model scales (35M and 650M parameters for both ESM2 and SaProt; 110M for ProSST) and both datasets, i.e., test ProteinGym (Table˜4) and validation MaveDB (Table˜A5).
+Notably, ProSST + ProteinTTT sets a new state of the art on the ProteinGym benchmark (Spearman correlation coefficients calculated for individual deep mutational scanning experiments (DMSs) have statistically significant difference according to a paired t-test with p < 0.05).
+We observe that ProteinTTT primarily improves performance for proteins with low MSA depth (i.e., the number of available homologous sequences), suggesting that single-sequence customization enhances predictions for proteins with fewer similar sequences in the training data (Table˜A4).
+The fact that ProteinTTT more effectively improves the performance of smaller ESM2 and SaProt models compared to their larger variants may be a result of the benchmark performance being saturated for larger models, consistent with a recent observation (Notin, 2025).
+We provide a qualitative example showing how ESM2 (650M) + ProteinTTT significantly improves fitness prediction by capturing residues critical for protein stability (Figure˜A5).
+We also demonstrate that customization can be combined with evolutionary information from MSA to further boost fitness prediction (Appendix˜B).
+4.3 Protein function prediction
+Finally, we demonstrate a proof of concept for customization in the context of protein function prediction.
+We experiment with two tasks: predicting protein location within a cell (Stärk et al., 2021), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (Samusevich et al., 2025).
+Appendix˜C shows that per-protein customization with ProteinTTT consistently enhances the performance of representative models on both tasks.
+5 Case studies
+ProteinTTTcan be readily incorporated into structure, fitness, or function prediction pipelines by adding several lines of code (Appendix˜D).
+Here, we demonstrate two challenging structure prediction case studies: improving modeling of antibody–antigen loops (Section˜5.1) and expanding known structures of viral proteins (Section˜5.2).
+5.1 Modeling antibody–antigen loops
+Figure 4: ProteinTTT improves modeling of antibody–antigen loops.
+(a) Average LDDT on the antibody complementarity-determining regions (CDRs, 175 structures) and antigens (814 structures) from the SAbDab dataset with ESMFold pLDDT < 70.
+Error bars indicate 95% confidence intervals estimated from 1000 bootstrap samples.
+(b) Example of improved structure prediction for CDRs in the 8K2W entry.
+The CDR regions H1, H2, and H3, i.e., the parts of the antibody that bind to the antigen, are highlighted with spheres, while black lines show the alignment error between the ground-truth CDR structure (white) and the predictions (colored).
+Accurately predicting structures of antibodies (e.g., human defensive proteins) and antigens (e.g., viral proteins) enables rational design of new therapeutics (Bennett et al., 2025).
+However, the presence of highly variable loop regions makes modeling of these interactions a long-standing challenge.
+Here, we show that ProteinTTT substantially improves structure prediction for these loop-formed complementarity-determining regions (CDRs) of antibodies, i.e., the parts that bind antigens, as well as for antigens themselves, on the well-established SAbDab dataset (Dunbar et al., 2014).
+We take the structures from SAbDab that are not predicted well by ESMFold (pLDDT < 70) and show that ProteinTTT improves the LDDT score for 115 of 175 antibody CDR substructures (66%) and 487 of 814 antigen chains (60%).
+As shown in Figure˜4a, ESMFold + ProteinTTT achieves significantly higher average LDDT scores compared to general-purpose ESMFold.
+Figure˜4b illustrates how ProteinTTT enables accurate prediction of all three CDRs in an antibody chain, providing an improved understanding of its binding interface with the corresponding antigen.
+5.2 Expanding known structures of viral proteins
+Figure 5: ProteinTTT expands the Big Fantastic Virus Database (BFVD).
+(a) ProteinTTT (light green) substantially improves the performance of ESMFold (yellow) on viral proteins, yielding better structures (pink) for 19% of BFVD entries compared to the original predictions by AlphaFold2 (green).
+(b) Improvements in pLDDT for ESMFold after ProteinTTT correspond to improvements in LDDT, as benchmarked against BFVD AlphaFold2 structures with pLDDT >> 90.
+(c) ProteinTTT provides the largest pLDDT improvements (y-axis) for the most out-of-distribution proteins, i.e., those with the smallest MSAs (left on the x-axis) from the Logan database.
+(d) Structural comparison for BFVD entry UPI000641889E against the PDB structure 2N2J (100% sequence identity) shows that ESMFold + ProteinTTT yields a prediction closest to the ground truth (gray), as also measured by LDDT.
+(e–g) Additional examples of high-quality viral structures (as measured by pLDDT) predicted with ESMFold + ProteinTTT but not with ESMFold or AlphaFold2.
+Higher pLDDT values are better.
+Predicting the structures of viral proteins is vital for vaccine development, antiviral design, and understanding infection (Bravi, 2024).
+Nevertheless, it remains challenging due to the high mutation rate, which often leaves viral proteins without close homologs or experimental structures in databases (Kim et al., 2025).
+Here, we demonstrate that per-protein customized predictions with ESMFold + ProteinTTT improve viral protein structure prediction, substantially expanding the Big Fantastic Virus Database—the repository of viral protein structures (Kim et al., 2025).
+Among all the entries in BFVD, predicted with AlphaFold2 through ColabFold (Mirdita et al., 2022) using MSAs constructed from Logan (Chikhi et al., 2024), only 55% have high-quality structure predictions (pLDDT > 70).
+We apply ESMFold and ESMFold + ProteinTTT to the BFVD entries to expand the database with higher-quality structures.
+This is achieved by applying all three methods to the specific protein and taking the predicted structure with the highest pLDDT.
+While ESMFold manages to improve the predicted structure (as measured by pLDDT) for 10% of the BFVD proteins, ESMFold + ProteinTTT leads to an improvement for 19% of the dataset entries, substantially increasing the quality of known viral protein structures (Figure˜5a).
+We validate that the improved pLDDT confidence values from ESMFold + ProteinTTT correlate with the quality of the predicted structures, as measured by LDDT against reference AlphaFold2 structures having pLDDT > 90 (Pearson = 0.875; Figure˜A9).
+Notably, the largest improvements in pLDDT align with the largest improvements in LDDT (Figure˜5b).
+We find that the benefit of customization saturates with the number of homologs available for a protein, indicating that ProteinTTT is most effective for challenging, out-of-distribution proteins (Figure˜5c).
+Finally, Figure˜5d–g shows examples where ProteinTTT enables high-confidence structure predictions in cases where general-purpose, uncustomized AlphaFold2 and ESMFold struggle.
+6 Discussion
+We introduce ProteinTTT, a method for customizing protein language models to individual targets.
+ProteinTTT consistently improves performance across various models, their scales, and downstream tasks.
+It excels on challenging, out-of-distribution examples where general models often fail.
+We demonstrate its practical value through two case studies: enhancing the structural prediction of difficult antibody-antigen loops and improving 19% of low-confidence viral protein structures in the Big Fantastic Virus Database.
+Our work establishes per-protein customization as a powerful and practical tool for biological research.
+Acknowledgments
+We thank Milot Mirdita for discussions and feedback about this work.
+This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through projects e-INFRA CZ [ID:90254], ELIXIR [LM2023055], CETOCOEN Excellence CZ.02.1.01/0.0/0.0/17_043/0009632, ESFRI RECETOX RI LM2023069.
+This work was also supported by the CETOCOEN EXCELLENCE Teaming project supported from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857560.
+This work was also supported by the European Union (CLARA (No.
+101136607), ERC FRONTIER (no.
+101097822), ELIAS (No.
+101120237)) and the Technology Agency of the Czech Republic under the NCC Programme from the state budget (No.
+RETEMED TN02000122) and under the NRP from the EU RRF (No.
+TEREP TN02000122/001N).
+This work was also supported by the Czech Science Foundation (GA CR) grant 21-11563M and by the European Union’s Horizon Europe program (ERC, TerpenCode, 101170268).
+Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council.
+Neither the European Union nor the granting authority can be held responsible for them.
+M.S.
+acknowledges support by the National Research Foundation of Korea grants (RS-2020-NR049543, RS-2021-NR061659 and RS-2021-NR056571, RS-2024-00396026), Creative-Pioneering Researchers Program and Novo Nordisk Foundation (NNF24SA0092560).
+Appendix
+Contents
+appendix.Aappendix.Bappendix.Cappendix.Dappendix.Esubsection.E.1subsubsection.E.1.1subsubsection.E.1.2subsubsection.E.1.3subsection.E.2subsubsection.E.2.1subsubsection.E.2.2subsubsection.E.2.3subsection.E.3subsubsection.E.3.1subsubsection.E.3.2subsubsection.E.3.3appendix.Fsubsection.F.1subsection.F.2appendix.Gsubsection.G.1subsection.G.2subsection.G.3
+Appendix A Justification of customization via perplexity minimization
+Figure A1: Quality of protein structure prediction, as measured by TM-score, correlates with perplexity of the underlying language model on the challenging targets from the CAMEO validation set.
+Higher TM-scores are associated with lower perplexity, indicating that better predictions are linked to lower uncertainty in the language model’s understanding of the protein sequence.
+While the paradigm of test-time customization has been investigated in other domains, the reasons behind its surprising effectiveness are not completely clear (Liu et al., 2021; Zhao et al., 2023).
+Here, we offer a potential justification for the effectiveness of ProteinTTT by linking it to perplexity minimization.
+Perplexity has traditionally been used in natural language processing to evaluate how well models comprehend sentences (Brown, 2020; Chelba et al., 2013).
+Protein language modeling has adopted this metric to assess how effectively models “understand” amino acid sequences (Hayes et al., 2024; Lin et al., 2023).
+For bidirectional, random masking language models, which are the focus of this study, we consider the following definition of perplexity [^2] Please note that this is an approximation of perplexity, which is computationally intractable for bidirectional models, and is often referred to as pseudo-perplexity (Lin et al., 2023; Salazar et al., 2019).:
+$\text{Perplexity}(x)=\exp\Biggl(\frac{1}{|x|}\sum_{i=1}^{|x|}{-\log{p(x_{i}|x_{\setminus i};\theta)}}\Biggl),$ (3)
+where $|x|$ is the length of the input protein sequence $x$ and $p(x_{i}|x_{\setminus i};\theta)$ represents the probability that the model correctly predicts the token $x_{i}$ at position $i$ when it is masked on the input $x_{\setminus i}$.
+Perplexity ranges from 1 to infinity (the lower, the better), providing an intuitive measure of how well a model fits, on average, tokens in a given sequence.
+A perplexity value of 1 indicates that the model perfectly fits the sequence, accurately predicting all the true tokens.
+Several studies have shown that lower perplexity on held-out protein sequences (calculated through the self-supervised track $g\circ f$) correlates with better performance on downstream tasks (via the supervised track $h\circ f$), such as predicting protein contacts (Rao et al., 2020), structure (Lin et al., 2023), or fitness (Kantroo et al., 2024).
+To give an example, we analyze the correlation between perplexity and structure prediction quality (Figure˜A1; see Section˜4.1 for experimental details).
+A notable correlation suggests that reducing a model’s perplexity on a single target sample $x$ (applied independently to all test samples) can lead to improved predictions on the downstream task (Figure˜3; Figure˜A10).
+Since we assume only a single target example $x$, the minimization of the masked language modeling loss $\mathcal{L}(x;\theta)$ (Equation˜2) on this example is directly linked to minimizing the perplexity Perplexity(x)(x) (Equation˜3).
+For instance, in the case of a single masked position (i.e., $|M|=1$), the loss is equal to the logarithm of perplexity.
+More generally, it can be shown formally that by minimizing the masked language modeling objective, the model learns to approximate the conditional marginals of the language (of proteins), including the leave-one-out probabilities evaluated in perplexity (Hennigen & Kim, 2023).
+As a result, applying self-supervised test-time customization on $x$ through $g\circ f$ enhances the representation of the target protein in the backbone $f$, leading to improved downstream performance via the fine-tuning track $h\circ f$.
+Appendix B Customization with multiple sequence alignment (MSA)
+Customization training objective.
+Since many target proteins may not have homologous sequences (Rao et al., 2021) and finding such homologs may be time-consuming (Lin et al., 2023), the ProteinTTT customization objective (Equation˜2) only assumes a single target sequence for customization.
+However, we also extend the loss function to the case when a multiple sequence alignment (MSA) is available:
+$\mathcal{L}_{\text{MSA}}(x;\theta)=\mathbb{E}_{x^{\prime}\sim p_{\text{MSA}}(x^{\prime}|x)}\big[\mathcal{L}(x^{\prime};\theta)\big],$ (4)
+[…]
+Results for fitness prediction.
+Table A1: ProteinTTT can be used with MSA when available. Please see Table˜4 for evaluation details.
+Method Avg. Spearman $\uparrow$
+ESM2 (Lin et al., 2023)
+0.4139
+[…]
+0.4299 ± 0.00099
+MSA Transformer (Rao et al., 2021)
+0.4319
+MSA Transformer + ProteinTTT (Ours)
+0.4326 ± 0.00003
+[…]
+Appendix C Customization for protein function prediction
+Protein function prediction is essential for understanding biological processes and guiding bioengineering, but is challenging due to its vague definition and limited data (Yu et al., 2023; Radivojac & et al., 2013; Stärk et al., 2021; Mikhael et al., 2024; Samusevich et al., 2025).
+While improved structure prediction with ProteinTTT (Section˜4.1) can already enhance function prediction (Song et al., 2024), we also evaluate our customization method directly on two function classification tasks: subcellular localization, predicting protein location within a cell (Stärk et al., 2021), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (Christianson, 2017; Samusevich et al., 2025).
+Using ProteinTTT with EnzymeExplorer (Samusevich et al., 2025) for TPS detection and Light attention (Stärk et al., 2021) for subcellular localization, we achieve consistent performance gains.
+Evaluation setup.
+Figure A2: Customization with ProteinTTT enables the correct substrate classification for a terpene synthase (TPS) enzyme.
+With progressive customization steps of EnzymeExplorer + ProteinTTT, the probability of the initially misclassified substrate (red) decreases, while the probability of the true substrates (green) increases.
+The bar plots also display the predicted probabilities for other substrates with non-zero values (grey).
+For the terpene substrate classification, we use the largest available dataset of characterized TPS from Samusevich et al.
+(2025) and reuse the original cross-validation schema.
+In the case of protein localization prediction, we use a standard DeepLoc dataset (Almagro Armenteros et al., 2017) as a validation set and setHard from Stärk et al.
+(2021) as the test set.
+Given a protein, the goal of function prediction is to correctly classify it into one of the predefined functional annotations.
+We assess the quality of the TPS substrate prediction using standard multi-label classification metrics used in the EnzymeExplorer paper (Samusevich et al., 2025): mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC).
+In the case of protein localization prediction, we similarly use the classification metrics from the original paper (Stärk et al., 2021): accuracy, multi-class Matthews correlation coefficient (MCC), and F1-score.
+Results.
+Table A2: Customization with ProteinTTT improves protein function prediction.
+For the terpene syntase (TPS) substrate classification task, the metrics are computed on the 512 TPS sequences based on the cross-validation schema of the TPS dataset (Samusevich et al., 2025).
+Subcellular localization prediction performance is reported for 432 protein sequences from the setHard test set (Stärk et al., 2021).
+The error bars show standard deviations across five random seeds.
+TPS substrate classification
+Method
+mAP $\uparrow$ AUROC $\uparrow$
+EnzymeExplorer
+(Samusevich et al., 2025)
+0.805 0.948
+EnzymeExplorer + ProteinTTT
+0.811 ± 0.0011 0.950 ± 0.0002
+Subcellular localization prediction
+Method
+Accuracy $\uparrow$ MCC $\uparrow$ F1-score $\uparrow$
+Light attention
+(Stärk et al., 2021)
+0.627 0.549 0.618
+Light attention + ProteinTTT
+0.634 ± 0.004 0.557 ± 0.005 0.627 ± 0.004
+Customization with ProteinTTT improves model performance on both of the protein function prediction tasks and across all considered metrics (Table˜A2).
+Figure˜A2 provides a qualitative result, where customization with ProteinTTT iteratively refines the prediction of EnzymeExplorer toward a correct TPS substrate class.
+We hypothesize that improvement with customization is more challenging in classification tasks, as opposed to regression problems, because a larger change in the latent space is required to shift the top-class probability.
+Appendix D Implementation details
+Infrastructure.
+All experiments with ProteinTTT are conducted on machines equipped with a single NVIDIA A100 40GB GPU, an 8-core AMD processor, and 128 GB of physical memory.
+Source code.
+We provide a user-friendly and easily extensible PyTorch (Paszke, 2019) implementation of ProteinTTT, available as the proteinttt Python package  [^3] https://github.com/anton-bushuiev/ProteinTTT .
+We provide two Python code snippets ˜1 and ˜2 to demonstrate the implementation of inference and customization with ProteinTTT, respectively.
+˜1 demonstrates how inference with ESMFold can be enhanced with ProteinTTT by adding just a few lines of code to enable customization.
+Next, ˜2 shows how ProteinTTT can be easily implemented for a PLM of interest by inheriting from the abstract TTTModule class.
+To integrate ProteinTTT within a model (e.g., ESM2), the user needs to implement methods that define the model’s vocabulary, an interface for predicting logits, and a specification of which modules need to be fine-tuned or remain frozen.
+The rest, i.e., the test-time training logic itself, is implemented within the unified TTTModule class.
+Listing 1: Incorporation of ProteinTTT into an ESMFold structure prediction pipeline using the proteinttt package.
+1import esm
+2from proteinttt.models.esmfold import ESMFoldTTT, DEFAULT_ESMFOLD_TTT_CFG
+3
+4# Set protein sequence
+5sequence = (
+6 "GIHLGELGLLPSTVLAIGYFENLVNIICESLNMLPKLEVSGKEYKKFKFTIVIPKDLDANIKKRAKIY"
+7 "FKQKSLIEIEIPTSSRNYPIHIQFDENSTDDILHLYDMPTTIGGIDKAIEMFMRKGHIGKTDQQKLLE"
+8 "ERELRNFKTTLENLIATDAFAKEMVEVIIEE"
+9)
+10
+11# Load model
+12model = esm.pretrained.esmfold_v1()
+13model = model.eval().cuda()
+14
+15predict_structure(model, sequence)
+16# pLDDT: 38.43025
+17
+18# ============================ ProteinTTT ===============================
+19# Customize model to sequence
+20model = ESMFoldTTT.ttt_from_pretrained(
+21 model, ttt_cfg=DEFAULT_ESMFOLD_TTT_CFG, esmfold_config=model.cfg
+22)
+23model.ttt(sequence)
+24# =======================================================================
+25
+26predict_structure(model, sequence)
+27# pLDDT: 78.69619
+28
+29# ============================ ProteinTTT ===============================
+30# Reset model to original state (after this model.ttt can be called with
+31# another protein)
+32model.ttt_reset()
+33# =======================================================================
+Optimization.
+We minimize the loss defined in Equation˜2 using stochastic gradient descent (SGD) with zero momentum and zero weight decay (Ruder, 2016).
+While a more straightforward option might be to use the optimizer state from the final pre-training step, this approach is often impractical because the optimizer parameters are usually not provided with the pre-trained model (Hayes et al., 2024; Lin et al., 2023).
 Moreover, many models are pre-trained using the Adam optimizer (Kingma & Ba, 2015) or its variants (Loshchilov & Hutter, 2019).
-However, it has been shown that Adam results in less predictable behavior of test-time training (TTT) compared to the SGD optimizer, possibly due to its more exploratory behavior (Gandelsman et al., 2022).
-Because each TTT experiment assumes only one test example available, we are not able to halt the training using early stopping on any validation sample.
-Therefore, for each choice of task-specific $f$ and $h$, we tune the optimal number of TTT steps using the entire validation set beforehand or rely on available performance estimates (e.g., pLDDT in the case of protein structure prediction; Section 4.2) to select the optimal number of optimization steps.
-Fine-tuning large models.
-We aim for test-time training to be applicable on the fly, i.e., without the need for any pre-computation and on a single GPU with a minimum computational overhead.
-This poses two challenges when using large models.
+However, it was shown that Adam results in less predictable behavior of test-time training compared to the SGD optimizer, possibly due to its more exploratory behavior (Gandelsman et al., 2022).
+Customizing large models.
+We aim for customization to be applicable on the fly, i.e., without the need for any pre-computation and on a single GPU with a minimum computational overhead.
+Since state-of-the-art models for many protein-oriented tasks are typically large, with up to billions of parameters, our aim presents two key challenges.
 First, when using pre-trained Transformers on a single GPU, even for the forward pass, the batch size is typically limited to only several samples due to the quadratic complexity of the inference (Vaswani, 2017).
@@
 To address the first challenge, we perform forward and backward passes through a small number of training examples and accumulate gradients to simulate updates with any batch size.
-We address the second challenge by employing low-rank adaptation (LoRA, Hu et al.
+We address the second challenge by employing low-rank adaptation (LoRA; Hu et al.
 (2021)), which in practice enables fine-tuning of any model for which a forward pass on a single sample is feasible, due to a low number of trainable parameters.
-Figure 3: Test-time training (TTT) improves protein structure prediction by reducing protein sequence perplexity.
-ESMFold fails to predict the structure of chain B from PDB entry 7EBL in the CAMEO validation set, as shown at TTT step 0, where the perplexity is high and the TM-score is low.
-By applying TTT on the single target sequence, the model iteratively improves the structure prediction quality, as demonstrated by the increasing TM-score, associated with reduced perplexity.
-At step 7, the predicted structure achieves the highest TM-score, as well as the highest predicted confidence metric pLDDT, enabling the selection of this step as the final prediction by ESMFold + TTT.
-3.2 Inference on downstream tasks
-Once the backbone $f$ is adapted to a test protein via self-supervised fine-tuning, it can be used in conjunction with a pre-trained downstream head $h$, as $h\circ f$.
-The key idea of TTT is not to update the head $h$ during test time, but rather to leverage improved input representations from $f$.
-Since Y-shaped architectures are prevalent in protein machine learning, TTT can be straightforwardly applied to numerous tasks in protein research.
-In this work, we address three primary challenges: protein fitness, structure, and function prediction, applying TTT to corresponding well-established models.
-For fitness prediction, we apply TTT to ESM2 (Lin et al., 2023) and SaProt (Su et al., 2023); for folding, we apply it to ESMFold (Lin et al., 2023) and ESM3 (Hayes et al., 2024); and for function prediction, we apply TTT to ESM-1v-based (Meier et al., 2021) TerpeneMiner (Samusevich et al., 2024) and ESM-1b-based (Rives et al., 2021) Light attention (Stärk et al., 2021).
-In all the models we consider, $f$ is a transformer encoder that takes a protein sequence as input (except for SaProt, which also uses structural tokens), while $g$ is a masked language modeling head (a layer mapping token embeddings to amino acid types).
-The downstream task heads $h$, however, vary significantly across tasks.
-For fitness prediction, $h$ outputs a single value for a mutated sequence, measuring how well the protein supports an organism’s functioning.
-Both ESM2 and SaProt use zero-shot inference from $g\circ f$ via log odds, where $h$ is a simple adaptation of $g$ without any additional parameters.
-For structure prediction, $h$ is a protein structure decoder: in ESMFold, it is an AlphaFold2-like structure prediction module (Jumper et al., 2021), while in ESM3, it is a VQ-VAE decoder (Razavi et al., 2019).
-The function predictors are classification models: in TerpeneMiner (Samusevich et al., 2024), $h$ is a random forest that outputs substrate probabilities, and in Light attention (Stärk et al., 2021), $h$ is a light attention module predicting localization class probabilities.
-Detailed descriptions of the models and their TTT adaptation are provided in Appendix A.
-3.3 Link between test-time training and perplexity minimization
-[…]
-[] Refer to caption
-Figure 4: The quality of protein structure prediction, as measured by TM-score, correlates with the perplexity of the underlying language model on the challenging targets from the CAMEO validation set.
-Higher TM-scores are associated with lower perplexity, indicating that better predictions are linked to lower uncertainty in the language model’s understanding of the protein sequence.
-While the approach of test-time training has been extensively investigated in computer vision and other domains, the reasons behind its effectiveness remain unclear (Liu et al., 2021; Zhao et al., 2023).
-Here, we offer a potential explanation by linking test-time training to perplexity minimization within the context of protein sequence modeling.
-Perplexity has traditionally been used in natural language processing to evaluate how well models comprehend test sentences (Brown, 2020; Chelba et al., 2013).
-Protein language modeling has adopted this metric to assess how effectively models understand amino acid sequences (Hayes et al., 2024; Lin et al., 2023).
-For bidirectional, random masking language models, which are the focus of this study, we consider the following definition of perplexity  [^2] Please note that this is an approximation of perplexity, which is computationally intractable for bidirectional models, and is often referred to as pseudo-perplexity (Lin et al., 2023; Salazar et al., 2019).:
-$\text{Perplexity}(x)=\exp\Biggl{(}\frac{1}{|x|}\sum_{i=1}^{|x|}{-\log{p(x_{i}|x_{\setminus i})}}\Biggl{)},$ (2)
-where $|x|$ is the length of the input protein sequence $x$ and $p(x_{i}|x_{\setminus i})$ represents the probability that the model correctly predicts the token $x_{i}$ at position $i$ when it is masked on the input $x_{\setminus i}$.
-Perplexity ranges from 1 to infinity (the lower the better), providing an intuitive measure of how well a model understands, on average, positions within a given sequence.
-A perplexity value of 1 indicates that the model perfectly understands the sequence, accurately predicting all the true tokens.
-Several studies have shown that lower perplexity on held-out protein sequences (calculated through the self-supervised track $g\circ f$) correlates with better performance on downstream tasks (via the supervised track $h\circ f$), such as predicting protein contacts (Rao et al., 2020), structure (Lin et al., 2023), or fitness (Kantroo et al., 2024).
-To provide an example, we analyze the correlation between perplexity and structure prediction performance (Figure 4).
-A strong correlation suggests that reducing a model’s perplexity on a single test sample $x$ can lead to improved performance on the downstream task (Figure 3; Figure 11).
-Since we consider only a single test example $x$, the minimization of the masked language modeling loss $\mathcal{L}(x)$ (Equation 1) on this example is directly linked to minimizing the perplexity Perplexity$(x)$ (Equation 2).
-For instance, in the case of a single masked position (i.e., $|M|=1$), the loss is equal to the logarithm of perplexity.
-More generally, it can be shown formally that by minimizing the masked language modeling objective, one learns to approximate the conditional marginals of the language (of proteins), including the leave-one-out probabilities evaluated in perplexity (Hennigen & Kim, 2023).
-As a result, applying test-time training (TTT) through $g\circ f$ enhances the representation of the test protein in the backbone $f$, leading to improved downstream performance via the fine-tuning track $h\circ f$.
-4 Experiments
-Building on the broad applicability of our test-time training (TTT) approach, we apply it to three example downstream tasks in protein machine learning: fitness, structure, and function prediction.
-The experimental setup and results for each task are presented in the following subsections.
-4.1 Protein Fitness Prediction
-Protein fitness refers to the ability of a protein to efficiently perform its biological function, which is determined by its structure, stability, and interactions with other molecules.
-Predicting protein fitness allows researchers to understand how mutations affect protein function, aiding in protein engineering (Notin et al., 2024).
-In this paper, we demonstrate that applying test-time training (TTT) to representative models, such as ESM2 (Lin et al., 2023) and SaProt (Su et al., 2023), enhances their protein fitness prediction capabilities.
-ESM2 is a protein language model trained on protein sequences, while SaProt is an extension of ESM2 that incorporates 3D information via additional structural tokens encoding structures predicted by AlphaFold2 (Jumper et al., 2021).
-Evaluation Setup.
-We evaluate the models using ProteinGym, the standard benchmark for fitness prediction (Notin et al., 2024), focusing specifically on its zero-shot setup.
-Since the zero-shot setup only provides a test set without any data split, we aim to validate TTT on independent data.
-To achieve this, we create a new fitness prediction dataset mined from MaveDB, a public repository containing datasets from Multiplexed Assays of Variant Effect (MAVEs) (Esposito et al., 2019).
-The quality of the new dataset is validated by confirming that both ESM2 and SaProt generalize well to the new data, achieving comparable performance (Appendix A).
-Given a protein and its variants, fitness prediction models output one real value per variant to estimate fitness.
-ProteinGym uses Spearman correlation between predicted and experimentally measured fitness values as the main evaluation metric for assessing the capabilities of models to score mutations.
-The correlation is first calculated for each protein and then aggregated per types of measured fitness: activity, binding, expression, organismal fitness, and stability.
-The final Spearman correlation metric is obtained by averaging across these five categories.
-We adopt this metric in our benchmarking.
-In our evaluation, we also include other top-performing baselines on the ProteinGym benchmark: TraceptEVE (Notin et al., 2022b) and GEMME (Laine et al., 2019).
-TraceptEVE combines language model Tranception (Notin et al., 2022a) with the protein-specific variational autoencoder, EVE, capturing the evolutionary information via MSAs (Frazer et al., 2021).
-GEMME is a statistical method deriving fitness predictions from evolutionary trees.
-Results.
-Test-time training (TTT) consistently enhances the protein fitness prediction performance of both ESM2 and SaProt models across varying model scales (35M and 650M parameters) and both datasets, test ProteinGym (Section 4.1 left) and validation MaveDB (Table 6 in Section B.2).
-Notably, SaProt (650M) + TTT sets a new state-of-the-art on the ProteinGym benchmark, achieving a 40% higher improvement compared to the previous leaderboard update (SaProt (650M) against TranceptEVE L).
-When examining performance across different phenotype categories, TTT yields improvements specifically in the categories where the baseline performance is weakest: “Organismal Fitness”, “Binding”, and “Activity” (Section 4.1 right).
-This improvement indicates the ability of TTT to enhance predictions on challenging targets.
-Additionally, we observe an inverse correlation between the degree of TTT enhancement and the depth of the MSA (i.e., the number of available homologous sequences) available for each test protein, suggesting that TTT primarily improves predictions for proteins with fewer similar sequences available in the training data (Table 5 in Section B.1).
-[…]
-[] Avg.
-Spearman $\uparrow$ Spearman by phenotype $\uparrow$ Activity Binding Expression Organismal Stability Fitness ESM2 (35M) (Lin et al., 2023) 0.3211 0.3137 0.2907 0.3435 0.2184 0.4392 ESM2 (35M) + TTT (Ours) 0.3407 ± 0.00014 0.3407 0.2942 0.3550 0.2403 0.4733 SaProt (35M) (Su et al., 2023) 0.4062 0.3721 0.3568 0.4390 0.2879 0.5749 SaProt (35M) + TTT (Ours) 0.4106 ± 0.00004 0.3783 0.3569 0.4430 0.2955 0.5795 ESM2 (650M) (Lin et al., 2023) 0.4139 0.4254 0.3366 0.4151 0.3691 0.5233 ESM2 (650M) + TTT (Ours) 0.4153 ± 0.00003 0.4323 0.3376 0.4168 0.3702 0.5195 TranceptEVE S* (Notin et al., 2022b) 0.4519 0.4750 0.3957 0.4426 0.4491 0.4973 GEMME* (Laine et al., 2019) 0.4547 0.4820 0.3827 0.4382 0.4517 0.5187 TranceptEVE M* (Notin et al., 2022b) 0.4548 0.4792 0.3858 0.4525 0.4538 0.5025 TranceptEVE L* (Notin et al., 2022b) 0.4559 0.4866 0.3758 0.4574 0.4597 0.5003 SaProt (650M) (Su et al., 2023) 0.4569 0.4584 0.3785 0.4884 0.3670 0.5919 SaProt (650M) + TTT (Ours) 0.4583 ± 0.00001 0.4593 0.3790 0.4883 0.3754 0.5896
-Table 1: Test-time training (TTT) improves protein fitness prediction.
-The right section of the table presents performance averaged across individual proteins and then across different protein phenotypes, as classified in the ProteinGym benchmark (Notin et al., 2024).
-The middle column shows the final performance, averaged across all five phenotype classes.
-In total, ProteinGym contains 2.5 million mutations across 217 proteins.
-Standard deviations are calculated over 5 random seeds and, for brevity, omitted in the right panel, where the maximum standard deviation does not exceed 0.0004.
-Methods marked with an asterisk (“*”) are the other top-5 methods in ProteinGym, and the metrics are reproduced from the leaderboard (https://proteingym.org/benchmarks).
-4.2 Protein structure prediction
-Protein structure prediction, also known as protein folding, is the task of predicting 3D coordinates of protein atoms given the amino acid sequence.
-Protein folding is arguably one of the most successful applications of machine learning in the life sciences (Jumper et al., 2021; Lin et al., 2023; Abramson et al., 2024).
-However, even state-of-the-art protein folding methods struggle to generalize to entirely novel proteins (Kryshtafovych et al., 2023).
-In this work, we focus on the ESMFold (Lin et al., 2023) and ESM3 (Hayes et al., 2024) models, demonstrating how their performance on challenging protein structures can be boosted by utilizing test-time training.
-Evaluation setup.
-To evaluate the performance of TTT, we use CAMEO, a standard benchmark for protein folding.
-We use the validation and test folds from Lin et al.
-(2023), focusing only on challenging targets by filtering them according to standard measures of prediction confidence based on pLDDT and perplexity (Section A.2).
-Given a protein sequence, the goal of protein folding is to predict 3D coordinates of the protein atoms.
-To assess the quality of the predicted protein structures with respect to the ground truth structures, we use two standard metrics: TM-score (Zhang & Skolnick, 2004) and LDDT (Mariani et al., 2013).
-TM-score measures the quality of the global 3D alignment of the target and predicted protein structures, while LDDT is an alignment-free method based on local distance difference tests.
-As baseline methods, we use techniques alternative to TTT for improving the performance of the pre-trained base models.
-In particular, the ESMFold paper proposes randomly masking 15% of amino acids in a protein sequence, allowing for sampling multiple protein structure predictions from the regression ESMFold model (Lin et al., 2023).
-For each sequence, we sample a number of predictions equal to the total number of TTT steps and refer to this baseline as ESMFold + MP (Masked Prediction).
-As a baseline for ESM3, we use chain-of-thought iterative decoding, referred to as ESM3 + CoT, proposed in the ESM3 paper (Hayes et al., 2024).
-Results.
-[…]
-[] TM-score $\uparrow$ LDDT $\uparrow$ ESM3 (Hayes et al., 2024) 0.3480 ± 0.0057 0.3723 ± 0.0055 ESM3 + CoT (Hayes et al., 2024) 0.3677 ± 0.0088 0.3835 ± 0.0024 ESM3 + TTT (Ours) 0.3954 ± 0.0067 0.4214 ± 0.0054 ESMFold (Lin et al., 2023) 0.4649 0.5194 ESMFold + MP (Lin et al., 2023) 0.4862 ± 0.0043 0.5375 ± 0.0070 ESMFold + TTT (Ours) 0.5047 ± 0.0132 0.5478 ± 0.0058
-Table 2: Test-time training (TTT) improves protein structure prediction.
-The metrics are averaged across the 18 challenging targets in the CAMEO test set, and standard deviations correspond to 5 random seeds.
-CoT and MP stand for the chain of though and masked prediction baselines.
-Test-time training (TTT) consistently improves the performance of both the ESMFold and ESM3 models, outperforming the masked prediction (ESMFold + MP) and chain-of-thought (ESM3 + CoT) baselines, as shown in Table 2.
-Of the 18 most challenging CAMEO test proteins, ESMFold and ESM3 significantly improved the prediction of 7 and 6 structures, respectively, while only slightly disrupting the prediction of 2 and 1 structures, respectively (Figure 9 in Section B.1).
-Most notably, TTT enables accurate structure prediction for targets that are poorly predicted with original base models.
-For instance, Figure 1 presents a strongly improved structure predicted using ESMFold + TTT for the target that was part of the CASP14 competition and shown as an unsuccessful case in the original ESMFold publication (Lin et al.
-(2023), Fig.
-2E).
-Another example is shown in Figure 3, where TTT refined the structure prediction from a low-quality prediction (TM-score $=0.29$) to a nearly perfectly folded protein (TM-score $=0.92$).
-4.3 Protein function prediction
-Protein function prediction is essential for understanding biological processes and guiding bioengineering efforts (Yu et al., 2023; Radivojac & et al., 2013).
-However, predicting protein function is challenging due to its vague definition and the scarcity of labeled datasets (Stärk et al., 2021; Mikhael et al., 2024; Samusevich et al., 2024).
-Therefore, enhancing generalization through TTT is particularly relevant in this context.
-While improved structure prediction through TTT can already enhance function prediction (Song et al., 2024), we further demonstrate that TTT can facilitate function prediction directly.
-Specifically, we focus on two tasks: subcellular localization, which predicts a protein’s location within a cell (Stärk et al., 2021), and substrate classification for terpene synthases (TPS)—enzymes responsible for the production of terpenoids, the largest class of natural products, and known for their complex biochemistry (Christianson, 2017; Samusevich et al., 2024).
-Unlike protein fitness or structure prediction, both tasks involve classification (TPS substrate classification being multi-label), providing a complementary evaluation of TTT.
-We apply TTT to TerpeneMiner (Samusevich et al., 2024) for TPS detection and Light attention (Stärk et al., 2021) for subcellular localization, demonstrating consistent improvements in both tasks.
-Evaulation setup.
-Figure 5: Test-time training (TTT) enables the correct substrate classification for a terpene synthase (TPS) enzyme.
-With progressive test-time training steps of TerpeneMiner + TTT, the probability of the initially misclassified substrate (red) decreases, while the probability of the true substrates (green) increases.
-The bar plots also display the predicted probabilities for other substrates with non-zero values (grey).
-For the terpene substrate classification, we use the largest available dataset of characterized TPS from Samusevich et al.
-(2024) and repurpose the original cross-validation schema.
-In the case of protein localization prediction, we use a standard DeepLoc dataset (Almagro Armenteros et al., 2017) as a validation set and setHard from (Stärk et al., 2021) as a test set.
-Given a protein, the goal of function prediction is to correctly classify it into one of the predefined functional annotations.
-We assess the quality of the TPS substrate prediction using standard multi-label classification metrics used in the TerpeneMiner paper (Samusevich et al., 2024): mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC).
-In the case of protein localization prediction, we similarly use the classification metrics from the original paper (Stärk et al., 2021): accuracy, multi-class Matthews correlation coefficient (MCC), and F1-score.
-Results.
-[…]
-[] TPS substrate classification mAP $\uparrow$ AUROC $\uparrow$ TerpeneMiner (Samusevich et al., 2024) 0.805 0.948 TerpeneMiner + TTT (Ours) 0.811 ± 0.0011 0.950 ± 0.0002 Subcellular localization prediction Accuracy $\uparrow$ MCC $\uparrow$ F1-score $\uparrow$ Light attention (Stärk et al., 2021) 0.627 0.549 0.618 Light attention + TTT (Ours) 0.634 ± 0.004 0.557 ± 0.005 0.627 ± 0.004
-Table 3: Test-time training (TTT) improves protein function prediction.
-For the terpene syntase (TPS) substrate classification task, the metrics are computed on the 512 TPS sequences based on the cross-validation schema of the TPS dataset (Samusevich et al., 2024).
-Subcellular localization prediction performance is reported for 432 protein sequences from the setHard test set (Stärk et al., 2021).
-The error bars show standard deviations across five random seeds.
-TTT improves the performance of the base models on both protein function prediction tasks and across all considered metrics (Table 3).
-Figure 5 provides a qualitative result, where TTT fine-tuning iteratively refines the prediction of TerpeneMiner toward a correct TPS substrate class.
-5 Discussion
-In this work, we have developed test-time training (TTT) for proteins, enabling per-protein adaptation of machine learning models for enhanced generalization.
-TTT improves performance across models, their scales, and benchmarks, while primarily enhancing performance on challenging targets.
-Our results open up the field of self-supervised adaptation for proteins and provide a proof-of-concept for other biology-related domains.
-While our method demonstrated strong potential, several underexplored directions remain for future work.
-In particular, gaining a comprehensive understanding of the success and failure modes of TTT and its application to emerging foundation models could unlock its full potential in computational biology.
-Additionally, our findings motivate future investigation of broader adaptation frameworks in the context of proteins.
-Among these frameworks are, for example, domain adaptation, where models leverage both training and test data to adapt to new domains (Ganin & Lempitsky, 2015), or adaptive risk minimization, which employs meta-learning to derive models that can adapt in the face of domain shifts (Zhang et al., 2021).
-Reproducibility Statement
-Our efforts are focused on ensuring that this research is easily reproducible.
-The proposed test-time training (TTT) method will be released as a Python package, providing easy-to-use wrappers for the models adapted in this paper.
-Detailed explanations of the application of TTT to individual models and the construction of datasets are included in the appendix.
-Where applicable, we will also release the source code for dataset generation.
-Acknowledgments
-This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through projects e-INFRA CZ [ID:90254], ELIXIR [LM2023055], CETOCOEN Excellence CZ.02.1.01/0.0/0.0/17_043/0009632, ESFRI RECETOX RI LM2023069.
-This work was also supported by the European Union (ERC project FRONTIER no.
-101097822) and the CETOCOEN EXCELLENCE Teaming project supported from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857560.
-This work was also supported by the Czech Science Foundation (GA CR) grant 21-11563M and by the European Union’s Horizon 2020 research and innovation programme under Marie Skłodowska-Curie grant agreement No.
-891397.
-Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council.
-Neither the European Union nor the granting authority can be held responsible for them.
-Appendix
-In Appendix A, we provide further details on the experimental setup, including comprehensive descriptions of the models, datasets, and metrics used.
-Next, in Appendix B, we present additional results and their analysis.
-We discuss the distribution of TTT effects and demonstrate that TTT primarily improves performance on challenging targets.
-We also explore the impact of hyperparameters by showing the performance on validation sets.
-Appendix A Experimental details
-In this section, we describe the experimental details for the three downstream tasks considered in this work: protein fitness prediction (Section A.1), protein structure prediction (Section A.2), and protein function prediction (Section A.3).
-Each subsection describes the application of test-time training (TTT) to the respective models, along with details on the datasets, metrics, and models.
-Table 4 additionally summarizes the hyperparameters used for the application of TTT to individual models.
-A.1 Protein fitness prediction
-A.1.1 Datasets
+Section˜G.3 details how ESMFold (Lin et al., 2023), with its 3B-parameter ESM2 backbone $f$, can be efficiently customized, retaining its speed advantage while enhancing performance.
+Listing 2: Implementation of ESM2 + ProteinTTT within the proteinttt package.
+1import torch
+2import esm
+3from esm.model.esm2 import ESM2
+4from proteinttt.base import TTTModule
+5
+6class ESM2TTT(TTTModule, ESM2):
+7 def __init__(self, ttt_cfg: TTTConfig, **kwargs):
+8 ESM2.__init__(self, **kwargs)
+9 TTTModule.__init__(self, ttt_cfg=ttt_cfg)
+10 self.ttt_alphabet = esm.Alphabet.from_architecture("ESM-1b")
+11 self.ttt_batch_converter = self.ttt_alphabet.get_batch_converter()
+12
+13 def _ttt_tokenize(self, seq: str, **kwargs):
+14 batch_labels, batch_strs, batch_tokens = self.ttt_batch_converter(
+15 [(None, seq)]
+16 )
+17 return batch_tokens
+18
+19 def _ttt_get_frozen_modules(self) -> list[torch.nn.Module]:
+20 return [self.embed_tokens]
+21
+22 def _ttt_mask_token(self, token: int) -> int:
+23 return self.ttt_alphabet.mask_idx
+24
+25 def _ttt_get_padding_token(self) -> int:
+26 return self.ttt_alphabet.padding_idx
+27
+28 def _ttt_token_to_str(self, token: int) -> str:
+29 return self.ttt_alphabet.all_toks[token]
+30
+31 def _ttt_get_all_tokens(self) -> list[int]:
+32 return [
+33 self.ttt_alphabet.tok_to_idx[t]
+34 for t in self.ttt_alphabet.all_toks
+35 ]
+36
+37 def _ttt_get_non_special_tokens(self) -> list[int]:
+38 return [
+39 self.ttt_alphabet.tok_to_idx[t]
+40 for t in self.ttt_alphabet.standard_toks
+41 ]
+42
+43 def _ttt_predict_logits(
+44 self, batch: torch.Tensor, start_indices: torch.Tensor = None
+45 ) -> torch.Tensor:
+46 return self(batch)["logits"]
+Appendix E Experimental details
+In this section, we describe the proposed benchmark suite for the three customization tasks considered in this work: protein structure prediction (Section˜E.1), protein fitness prediction (Section˜E.2), and protein function prediction (Section˜E.3).
+Each subsection describes the application of ProteinTTT to the respective models, along with details on the data, metrics, and models.
+Table˜A3 additionally summarizes the hyperparameters used for the application of ProteinTTT to individual models.
+E.1 Protein structure prediction
+E.1.1 Datasets
+CAMEO dataset.
+To evaluate the capabilities of ProteinTTT on protein structure prediction, we employ the CAMEO validation and test sets as described in Lin et al.
+(2023).
+Specifically, the validation set was obtained by querying the CAMEO (Continuous Automated Model Evaluation) web server [^4] https://www.cameo3d.org/modeling (Robin et al., 2021) for entries between August 2021 and January 2022, while the CAMEO test set consists of entries from April 1, 2022, to June 25, 2022.
+Most of the entries in the CAMEO sets are predicted with high accuracy and confidence (Lin et al., 2023).
+Therefore, we subselect the challenging validation and test sets where customization with ProteinTTT is suitable.
+Specifically, we apply two standard criteria: (1) preserving entries with ESMFold pLDDT scores below 70 to filter out high-confidence predictions (Jumper et al., 2021), and (2) selecting entries with ESM2 perplexity scores greater than or equal to 6, ensuring that the predictions are challenging due to poor sequence understanding rather than other factors.
+Additionally, most structures with perplexity scores below 6 are already associated with high-confidence predictions (Figure S5 in Lin et al.
+(2023)).
+After filtering, the resulting challenging validation and test sets consist of 27 (out of 378) and 18 (out of 194) targets, respectively.
+E.1.2 Metrics
+To assess the quality of the predicted protein structures with respect to the ground truth structures, we use two standard metrics averaged across the test dataset: TM-score (Zhang & Skolnick, 2004) and LDDT (Mariani et al., 2013).
+TM-score.
+The TM-score (Template Modeling score) is a metric used to assess the quality of the global 3D alignment between the predicted and target protein structures.
+It evaluates the structural similarity by comparing the distance between corresponding residues after superposition.
+The TM-score ranges from 0 to 1, where higher values indicate better alignment.
+LDDT.
+The Local Distance Difference Test (LDDT) is an alignment-free metric used to assess the accuracy of predicted protein structures.
+Unlike global metrics, LDDT focuses on local structural differences by measuring the deviation in distances between atom pairs in the predicted structure compared to the target structure.
+It is particularly useful for evaluating the accuracy of local regions, such as secondary structure elements.
+LDDT scores range from 0 to 100, with higher values indicating better local structural agreement.
+E.1.3 Models
+ESMFold.
+The ESMFold architecture comprises two key components: a protein language model, ESM2, which, given a protein sequence, generates embeddings for individual amino acids, and a folding block that, using these embeddings and the sequence, predicts the protein 3D structure along with per-amino-acid confidence scores, known as pLDDT scores.
+In our experiments, we use the esmfold_v0 model from the publicly available ESMFold checkpoints [^5] https://github.com/facebookresearch/esm/blob/main/esm/esmfold/v1/pretrained.py.
+Please note that we use esmfold_v0 and not esmfold_v1 to avoid data leakage with respect to the CAMEO test set.
+ESMFold + ProteinTTT.
+Since the ESM2 backbone of ESMFold was pre-trained in a self-supervised masked modeling regime, the application of ProteinTTT to ESMFold is straightforward.
+We treat ESM2 as the backbone $f$, the language modeling head predicting amino acid classes from their embeddings as the self-supervised head $g$, and the folding trunk along with the structure modules as the downstream task head $h$.
+After each ProteinTTT step, we run $h\circ f$ to compute the pLDDT scores, which allows us to estimate the optimal number of customization steps for each protein based on the highest pLDDT score.
+Since the backbone $f$ is given by the ESM2 model containing 3 billion parameters, we apply LoRA (Hu et al., 2021) to all matrices involved in self-attention.
+This enables fine-tuning ESMFold + ProteinTTT on a single GPU.
+ESMFold + ME.
+Since ESMFold is a regression model, it only predicts one solution and does not have a straightforward mechanism for sampling multiple structure predictions.
+Nevertheless, the authors of ESMFold propose a way to sample multiple candidates (Section A.3.2 in Lin et al.
+(2023)).
+To sample more predictions, the masking prediction (ME) method randomly masks 15% (same ratio as during masked language modeling pre-training) of the amino acids before passing them to the language model.
+Selecting the solution with the highest pLDDT may lead to improved predicted structure.
+Since sampling multiple solutions with ESMFold + ME and selecting the best one via pLDDT is analogous to ESMFold + ProteinTTT, we employ the former as a baseline, running the method for the same number of steps.
+ESM3.
+Unlike ESMFold, ESM3 is a fully multiple-track, BERT-like model (Devlin, 2018), pre-trained to unmask both protein sequence and structure tokens simultaneously (along with the function tokens).
+The structure tokens in ESM3 are generated via a separately pre-trained VQ-VAE (Razavi et al., 2019) operating on the protein geometry.
+In our experiments, we use the smallest, publicly available version of the ESM3 model (ESM3_sm_open_v0) [^6] https://github.com/evolutionaryscale/esm.
+ESM3 + ProteinTTT.
+We treat the Transformer encoder of ESM3 as $f$, the language modeling head decoding amino acid classes as $g$, and the VQ-VAE decoder, which maps structure tokens to the 3D protein structure, as $h$.
+During the customization steps, we train the model to unmask a protein sequence while keeping the structural track fully padded.
+During the inference, we provide the model with a protein sequence and run it to unmask the structural tokens, which are subsequently decoded with the VQ-VAE decoder.
+After each customization step, we run $h\circ f$ to compute the pLDDT scores, which allows us to estimate the optimal number of customization steps for each protein based on the highest pLDDT score.
+We choose the optimal hyperparameters by maximizing the difference in TM-score after and before applying ProteinTTT across the validation dataset.
+Despite the fact that the model contains 1.4 billion parameters, even without using LoRA, ESM3 + ProteinTTT can be fine-tuned on a single NVIDIA A100 GPU.
+Therefore, we do not employ LoRA for fine-tuning ESM3, while this can also be possible.
+ESM3 + CoT.
+To improve the generalization and protein-specific performance of ESM3, the original ESM3 paper employs a chain of thought (CoT) procedure.
+The procedure unfolds in $n$ steps as follows.
+At each step, $1/n$ of the masked tokens with the lowest entropy after softmax on logits are unmasked.
+Then, the partially unmasked sequence is fed back into the model, and the process repeats until the entire sequence is unmasked.
+In our experiments, we set $n=8$, which is the default value provided in the official GitHub repository.
+HelixFold-Single.
+HelixFold-Single is an MSA-free protein structure prediction model that combines representations from a pretrained protein language model with adapted AlphaFold2 geometric modules (EvoformerS and Structure) to directly predict atomic coordinates (Fang et al., 2023).
+We use the official implementation [^7] https://github.com/PaddlePaddle/PaddleHelix/tree/dev/apps/protein_folding/helixfold-single
+HelixFold-Single + ProteinTTT.
+HelixFold-Single shares the main concept with ESMFold, and we combine it with ProteinTTT in the same way as in ESMFold + ProteinTTT.
+E.2 Protein fitness prediction
+E.2.1 Datasets
 ProteinGym.
-ProteinGym is the standard benchmark for protein fitness prediction (Notin et al., 2024).
+ProteinGym [^8] https://github.com/OATML-Markslab/ProteinGym is the standard benchmark for protein fitness prediction (Notin et al., 2024).
 The latest, second version of the dataset includes 217 deep mutation scanning experiments (DMSs) across different proteins.
-In total, the dataset contains 2.5 mutants with annotated ground-truth fitness.
+We focus on the well-established zero-shot setup of the benchmark and do not experiment with the supervised setup, as it has not yet been fully incorporated into the official codebase at the time of this study.
+In total, the dataset contains 2.5M mutants with annotated ground-truth fitness.
 Since ProteinGym does not contain a data split for the zero-shot setup, employed in this work, we use the whole dataset as the test set.
 MaveDB dataset.
-Figure 6: Comparison of the standard ProteinGym dataset with the MaveDB dataset constructed in this work.
+Figure A3: Comparison of the standard ProteinGym dataset with the MaveDB dataset constructed in this work.
 A) MaveDB, mined from Esposito et al.
@@
 This corresponds to 483 unique protein sequences and 867 thousand mutations in total.
-The large size of the dataset, despite the comprehensiveness of ProteinGym containing 217 assays, can be attributed to the fact that many assays in MaveDB were released after the ProteinGym construction (Figure 6A).
-To ensure the quality of the constructed MaveDB dataset, we validated that representative baselines from ProteinGym generalize to the new assays, following a similar distribution of predictions (Figure 6B,C).
-Finally, for efficiently tuning hyper-parameters for fitness prediction models we sampled 50 random proteins (Figure 6D), corresponding to 83 assays and collectively 134 thousand variants.
-A.1.2 Metrics
+The large size of the dataset, despite the comprehensiveness of ProteinGym containing 217 assays, can be attributed to the fact that many assays in MaveDB were released after the ProteinGym construction (Figure˜A3A).
+To ensure the quality of the constructed MaveDB dataset, we validated that representative baselines from ProteinGym generalize to the new assays, following similar distributions of predictions (Figure˜A3B,C).
+Finally, for efficiently tuning hyperparameters for fitness prediction models, we sampled 50 proteins (Figure˜A3D), corresponding to 83 assays comprising 134 thousand variants.
+E.2.2 Metrics
 Protein fitness labels are not standardized and can vary across different proteins.
@@
 First, for each protein, we compute Spearman correlation scores between the predicted ranks of mutations and their corresponding labels.
-Then, we average the scores across five categories of assayed phenotypes, measuring the effects of introduced mutations: protein catalytic activity (“Activity”), binding affinity to a target (“Binding”), protein expression levels in a cell (“Expression”), organism growth rate (“Organismal Fitness”), and protein thermostability (“Stability”).
+Then, we average the scores across five categories of assayed phenotypes, measuring the effects of mutations: catalytic activity (“Activity”), binding affinity to a target (“Binding”), protein expression levels in a cell (“Expression”), organism growth rate (“Organismal Fitness”), and protein thermostability (“Stability”).
 Avg. Spearman.
@@
 We report the “Avg.
-Spearman” metric as the mean and standard deviation across five random seeds (Section 4.1, Table 5).
+Spearman” metric as the mean and standard deviation across five random seeds (Table˜4, Table˜A4).
 Spearman by MSA Depth.
-Following (Notin et al., 2024), we split the performance by the depth of available multiple sequence alignment (MSA), i.e., the number of homologous sequences available, as provided in ProteinGym: “Low depth”, “Medium depth”, and “High depth”, and report the Spearman correlation for each subset individually (Table 6).
+Following (Notin et al., 2024), we split the performance by the depth of available multiple sequence alignment (MSA), i.e., the number of homologous sequences available, as provided in ProteinGym: “Low depth”, “Medium depth”, and “High depth”, and report the Spearman correlation for each subset individually (Table˜A4).
 Specifically, the MSA depth categories in ProteinGym are determined using the following thresholds from Hopf et al.
-(2017): “Low” is defined as $N_{eff}/L<1$, “Medium” as $1001<N_{eff}/L<$ 100, and “High” as $N_{eff}/L>100$, where $N_{eff}$ represents the normalized number of effective sequences in the MSA, and $L$ is the sequence length covered in the MSA.
-A.1.3 Models
+(2017): “Low” is defined as $N_{eff}/L<1$, “Medium” as $1<N_{eff}/L<100$, and “High” as $N_{eff}/L>100$, where $N_{eff}$ represents the normalized number of effective sequences in the MSA, and $L$ is the sequence length covered in the MSA.
+E.2.3 Models
 ESM2.
 The ESM2 model is a bidirectional, BERT-like (Devlin, 2018) Transformer trained on millions of protein sequences using masked modeling (Lin et al., 2023).
-The goal of protein fitness prediction is to predict the effects of mutations, and protein language models are often adapted to this task using zero-shot transfer via log odds ratio (Notin et al., 2024; Meier et al., 2021).
+The goal of protein fitness prediction is to predict the effects of mutations, and PLMs are often adapted to this task using zero-shot transfer via log odds ratio (Notin et al., 2024; Meier et al., 2021).
 Specifically, for a given single- or multi-point mutation, where certain amino acids $T$ are substituted from $x_{i}$ to $x^{m}_{i}$ for each $i\in T$, the fitness prediction via the log odds ratio is defined as:
-$\sum_{i\in T}{\log{p(x^{m}_{i}|x_{\setminus i})}-\log{p(x_{i}|x_{\setminus i})}},$ (3)
+$\sum_{i\in T}{\Bigl(\log{p(x^{m}_{i}|x_{\setminus i})}-\log{p(x_{i}|x_{\setminus i})}\Bigr)},$ (5)
 where the sum iterates over mutated positions $i\in T$ with $p(x^{m}_{i}|x_{\setminus i})$ and $p(x_{i}|x_{\setminus i})$ denoting the predicted probabilities of the mutated amino acid and the original one (i.e., wild type), respectively.
@@
 Mutations with negative values represent favorable mutations, while positive values correspond to disruptive mutations.
-We follow the ProteinGym benchmark [^4] https://github.com/OATML-Markslab/ProteinGym and use this formula (Notin et al., 2024) to evaluate the fitness prediction capabilities of ESM2.
+We follow the ProteinGym benchmark and use this formula (Notin et al., 2024) to evaluate the fitness prediction capabilities of ESM2.
 We use the implementation of ESM2 from ProteinGym.
-ESM2 + TTT.
-ESM2 can be straightforwardly enhanced with test-time training.
-Specifically, we treat the transformer encoder as the backbone $f$, and the language modeling head, which projects token embeddings to amino acid probabilities, as the pre-training head $g$.
-The log odds ratio given by Equation 3 serves as the task-specific head $h$, which in this case involves the pre-training head $g$ that predicts log probabilities.
-Overall, we apply TTT to the pre-trained ESM2 model and, after a pre-defined number of self-supervised fine-tuning steps, score mutations using Equation 3.
-During TTT we fine-tune all parameters in $g\circ f$ end-to-end except for token and position embeddings.
+ESM2 + ProteinTTT.
+[…]
 SaProt.
-We also experiment with the state-of-the-art fitness prediction model, SaProt (Su et al., 2023).
+We also experiment with a structure-aware protein language model, SaProt (Su et al., 2023).
 SaProt builds off the ESM2 model but incorporates structural information from predicted protein structures.
@@
 The 3Di tokens capture the geometry of the protein backbone and are generated using VQ-VAE (Razavi et al., 2019), which projects continuous geometric information into discrete tokens and was trained as part of the Foldseek method (van Kempen et al., 2022).
-Since SaProt is also a protein language model, it also uses Equation 3 to score variants.
+Since SaProt is also a protein language model, it also uses Equation˜5 to score variants.
 However, please note that SaProt, as implemented in ProteinGym (Notin et al., 2024), uses a slightly different version of the log odds ratio.
-In SaProt, the conditions in the log probabilities in Equation 3 are replaced with $x_{\setminus T}$ instead of $x_{\setminus i}$, not assuming the independence of substitutions.
-During TTT, we only mask sequential information and leave the structural part of the tokens unchanged, reflecting the original pre-training setup.
-We use the implementation of SaProt from ProteinGym4.
-SaProt + TTT.
-Since the architecture of SaProt is based on ESM2, the TTT components $f$, $g$, and $h$ remain the same.
-It means that test-time training can be applied to the model in the same way as in the case of ESM2 + TTT discussed above.
-A.2 Protein structure prediction
-A.2.1 Datasets
-CAMEO dataset.
-To evaluate the capabilities of TTT on protein folding, we employ the CAMEO validation and test sets as described in Lin et al.
-(2023).
-Specifically, the validation set was obtained by querying the CAMEO (Continuous Automated Model Evaluation) web server [^5] https://www.cameo3d.org/modeling (Robin et al., 2021) for entries between August 2021 and January 2022, while the CAMEO test set consists of entries from April 1, 2022, to June 25, 2022.
-Most of the entries in the CAMEO sets are predicted with high accuracy and confidence (Lin et al., 2023).
-Therefore, we subselected the challenging validation and test sets where TTT is relevant.
-Specifically, we applied two criteria: (1) preserving entries with ESMFold pLDDT scores below 70 to filter out high-confidence predictions (Jumper et al., 2021), and (2) selecting entries with ESM2 perplexity scores greater than or equal to 6, ensuring that the predictions are challenging due to poor sequence understanding rather than other factors.
-Additionally, most structures with perplexity scores below 6 are already associated with high-confidence predictions (Figure S5 in Lin et al.
+In SaProt, the conditions in the log probabilities in Equation˜5 are replaced with $x_{\setminus T}$ instead of $x_{\setminus i}$, not assuming the independence of substitutions.
+During customization with ProteinTTT, we only mask sequential information and leave the structural part of the tokens unchanged, reflecting the original pre-training setup.
+We use the implementation of SaProt from ProteinGym8.
+SaProt + ProteinTTT.
+Since the architecture of SaProt is based on ESM2, the ProteinTTT components $f$, $g$, and $h$ remain the same.
+It means that customization can be applied to the model in the same way as in the case of ESM2 + ProteinTTT discussed above.
+ProSST.
+Finally, we experiment with the state-of-the-art fitness predictor, ProSST (Li et al., 2024).
+ProSST primarily improves upon SaProt (Su et al., 2023) by incorporating a larger vocabulary of structural tokens and employing disentangled attention mechanisms.
+Instead of relying on the 3Di alphabet optimized for protein structure search with Foldseek (van Kempen et al., 2022), Li et al.
+(2024) pre-train a new autoencoder to denoise corrupted protein backbones and cluster the resulting latent space using the $K$-means algorithm (Lloyd, 1982).
+Notably, optimal performance for fitness prediction is achieved with $K=2048$ tokens, compared to just 20 in the 3Di vocabulary used by SaProt.
+We adopt this model in our experiments.
+Additionally, disentangled attention in ProSST enhances information propagation between sequence and structure within its Transformer blocks, further improving prediction performance.
+The model has 110M parameters in total.
+ProSST, similarly to ESM2 and SaProt, is pre-trained using masked language modeling applied to protein sequence tokens.
+To score mutations on the ProteinGym benchmark (Notin et al., 2024), ProSST also uses the log-odds ratio, but in a slightly different way compared to ESM2 and SaProt.
+Specifically, ProSST performs a single forward pass to predict log probabilities, which are then used to score all mutations.
+Formally, this approach modifies the log probability condition in Equation˜5, replacing $x_{\setminus i}$ with $x$.
+ProSST + ProteinTTT.
+Similarly to ESM2 and SaProt, we treat the Transformer encoder in ProSST as the backbone $f$, the masked language modeling head as the pre-training head $g$, and the log-odds ratio formula as the task-specific head $h$.
+MSA Transformer.
+Finally, we experiment with MSA Transformer for fitness prediction (Rao et al., 2021).
+Similar to ESM2 (Lin et al., 2023), MSA Transformer is pre-trained on large protein sequence datasets; however, it is trained on multiple sequence alignments (MSAs) rather than individual sequences.
+Since MSA Transformer is also a protein language model, it can be used for fitness prediction in the same way as ESM2, as discussed above, by computing the log-odds ratio over the first sequence in the MSA in this case.
+We reproduce the results of MSA Transformer on the ProteinGym benchmark with two modifications: (1) we sample a weighted subset of 32 sequences from each MSA instead of 400, and (2) we use only one random seed instead of five for ensembling.
+These changes significantly reduce computational time while also slightly improving performance compared to the results reported in ProteinGym.
+This improvement may be explained by the fact that the performance of MSA Transformer saturates with increasing MSA input size (Figure 4 in Rao et al.
 (2021)).
-After filtering, the resulting challenging validation and test sets consist of 27 (out of 378) and 18 (out of 194) targets, respectively.
-The vast majority of the remaining structures have accurate ESMFold structure predictions.
-A.2.2 Metrics
-To assess the quality of the predicted protein structures with respect to the ground truth structures, we use two standard metrics averaged across the test dataset: TM-score (Zhang & Skolnick, 2004) and LDDT (Mariani et al., 2013).
-TM-score.
-The TM-score (Template Modeling score) is a metric used to assess the quality of the global 3D alignment between the predicted and target protein structures.
-It evaluates the structural similarity by comparing the distance between corresponding residues after superposition.
-The TM-score ranges from 0 to 1, where higher values indicate better alignment.
-LDDT.
-The Local Distance Difference Test (LDDT) is an alignment-free metric used to assess the accuracy of predicted protein structures.
-Unlike global metrics, LDDT focuses on local structural differences by measuring the deviation in distances between atom pairs in the predicted structure compared to the target structure.
-It is particularly useful for evaluating the accuracy of local regions, such as secondary structure elements.
-LDDT scores range from 0 to 100, with higher values indicating better local structural agreement.
-A.2.3 Models
-ESMFold.
-The ESMFold architecture comprises two key components: a protein language model, ESM2, which, given a protein sequence, generates embeddings for individual amino acids, and a folding block that, using these embeddings and the sequence, predicts the protein 3D structure along with per-amino-acid confidence scores, known as pLDDT scores.
-In our experiments, we use the esmfold_v0 model from the publicly available ESMFold checkpoints [^6] https://github.com/facebookresearch/esm/blob/main/esm/esmfold/v1/pretrained.py.
-Please note that we use esmfold_v0 and not esmfold_v1 to avoid data leakage with respect to the CAMEO test set.
-ESMFold + TTT.
-Since ESM2 backbone of ESMFold was pre-trained in a self-supervised masked modeling regime, the application of TTT to ESMFold is straightforward.
-We treat ESM2 as the backbone $f$, the language modeling head predicting amino acid classes from their embeddings as the self-supervised head $g$, and the folding trunk along with the structure modules as the downstream task head $h$.
-After each TTT step, we run $h\circ f$ to compute the pLDDT scores, which allows us to estimate the optimal number of TTT steps for each protein based on the highest pLDDT score.
-Since the backbone $f$ is given by the ESM2 model containing 3 billion parameters, we apply LoRA (Hu et al., 2021) to all matrices involved in self-attention.
-This enables fine-tuning ESMFold + TTT on a single GPU.
-ESMFold + ME.
-Since ESMFold is a regression model, it only predicts one solution and does not have a straightforward mechanism of sampling multiple structure predictions.
-Nevertheless, the authors of ESMFold propose a way to sample multiple candidates (Section A.3.2 in Lin et al.
-(2023)).
-To sample more solutions, the masking prediction (ME) method randomly masks 15% (same ratio as during masked language modeling pre-training) of the amino acids embeddings before passing them to the structure prediction block.
-Selecting the solution with the highest pLDDT may lead to improved predicted structure.
-Since sampling multiple solutions with ESMFold + ME and selecting the best one via pLDDT is analogous to ESMFold + TTT, we employ the former as a baseline, running the method for the same number of step.
-ESM3.
-Unlike ESMFold, ESM3 is a fully multiple-track, BERT-like model (Devlin, 2018), pre-trained to unmask both protein sequence and structure tokens simultaneously (along with the function tokens).
-The structure tokens in ESM3 are generated via a separately pre-trained VQ-VAE (Razavi et al., 2019) operating on the protein geometry.
-In our experiments, we use the smallest, publicly available version of the ESM3 model (ESM3_sm_open_v0) [^7] https://github.com/evolutionaryscale/esm.
-ESM3 + TTT.
+MSA Transformer + ProteinTTT.
 […]
-[] Refer to caption
-Figure 7: Average running time of ESM3 + TTT (50 steps) on the CAMEO validation set using a single NVIDIA A100 GPU.
-We treat the transformer encoder of ESM3 as $f$, the language modeling head decoding amino acid classes as $g$, and the VQ-VAE decoder, which maps structure tokens to the 3D protein structure, as $h$.
-During the TTT steps, we train the model to unmask a protein sequence while keeping the structural track fully padded.
-During the inference, we provide the model with a protein sequence and run it to unmask the structural tokens, which are subsequently decoded with the VQ-VAE decoder.
-After each TTT step, we run $h\circ f$ to compute the pLDDT scores, which allows us to estimate the optimal number of TTT steps for each protein based on the highest pLDDT score.
-We choose the optimal hyperparameters by maximizing the difference in TM-score after and before applying TTT across the validation dataset.
-Despite the fact that the model contains 1.4 billion parameters, even without using LoRA, ESM3 + TTT can be fine-tuned on a single NVIDIA A100 GPU.
-Therefore, we do not employ LoRA for fine-tuning ESM3.
-Figure 7 shows that ESM3 + TTT can be loaded in VRAM and run for 50 steps in just 10 seconds, with an additional 10 seconds overhead for pLDDT estimation.
-For comparison, the base ESM3 model can be run on the same hardware in 1 second.
-ESM3 + CoT.
-To improve the generalization and protein-specific performance of ESM3, the original ESM3 paper employs a chain of thought (CoT) procedure.
-The procedure unfolds in $n$ steps as follows.
-At each step, $1/n$ of the masked tokens with the lowest entropy after softmax on logits are unmasked.
-Then, the partially unmasked sequence is fed back into the model, and the process repeats until the entire sequence is unmasked.
-In our experiments, we set $n=8$, which is the default value provided in the official GitHub repository.
-A.3 Protein function prediction
-A.3.1 Datasets
+E.3 Protein function prediction
+E.3.1 Datasets
 TPS dataset.
@@
 Since setHard was constructed to be independent of DeepLoc, setHard provides a leakage-free source of data for validation.
-A.3.2 Metrics
+E.3.2 Metrics
 mAP, AUROC.
@@
 Therefore, we assess the quality of the predictions using standard multi-label classification metrics such as mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC) averaged across individual classes.
-These metrics were used in the original TerpeneMiner paper (Samusevich et al., 2024).
+These metrics were used in the original EnzymeExplorer paper (Samusevich et al., 2025).
 We report the performance by averaging the metric values concatenated across all validation folds from the 5-fold cross-validation schema.
@@
 The F1-score, the harmonic mean of precision and recall, evaluates performance from a retrieval perspective, balancing the trade-off between false positives and false negatives.
-A.3.3 Models
-TerpeneMiner.
-TerpeneMiner is a state-of-the-art method for the classification of terpene synthase (TPS) substrates (Samusevich et al., 2024).
+E.3.3 Models
+EnzymeExplorer.
+EnzymeExplorer is a state-of-the-art method for the classification of terpene synthase (TPS) substrates (Samusevich et al., 2025).
 The model consists of two parallel tracks.
-Given a protein sequence, TerpeneMiner first computes its ESM-1v embedding (Meier et al., 2021) and a vector of similarities to the functional domains of proteins from the training dataset, based on unsupervised domain segmentation of AlphaFold2-predicted structures (Jumper et al., 2021).
+Given a protein sequence, EnzymeExplorer first computes its ESM-1v embedding (Meier et al., 2021) and a vector of similarities to the functional domains of proteins from the training dataset, based on unsupervised domain segmentation of AlphaFold2-predicted structures (Jumper et al., 2021).
 The ESM-1v embedding and the similarity vector are then concatenated and processed by a separately trained random forest, which predicts TPS substrate class probabilities.
-In our experiments, we use the “PLM only” version of the model, which leverages only ESM-1v embeddings (PLM stands for protein language model).
-This version exhibits a minor performance decrease compared to the full model but exactly follows a Y-shaped architecture, allowing us to validate the effectiveness of test-time training for predicting TPS substrates.
-We use the implementation of TerpeneMiner available at the official GitHub page  [^8] https://github.com/pluskal-lab/TerpeneMiner.
-TerpeneMiner + TTT.
-When applying TTT to TerpeneMiner, we treat the frozen ESM-1v model as a backbone $f$, its language modeling head as a self-supervised head $g$, and the random forest classifying TPS substrates as a downstream supervised head $h$.
+In our experiments, we use the “PLM only” version of the model, which leverages only ESM-1v embeddings.
+This version exhibits a minor performance decrease compared to the full model but exactly follows a Y-shaped architecture, allowing us to validate the effectiveness of ProteinTTT for predicting TPS substrates.
+We use the implementation of EnzymeExplorer available at the official GitHub page  [^10] https://github.com/pluskal-lab/EnzymeExplorer.
+EnzymeExplorer + ProteinTTT.
+When applying ProteinTTT to EnzymeExplorer, we treat the frozen ESM-1v model as a backbone $f$, its language modeling head as a self-supervised head $g$, and the random forest classifying TPS substrates as a downstream supervised head $h$.
 Light Attention.
@@
 We re-train the model using ESM-1b embeddings on the DeepLoc dataset (Almagro Armenteros et al., 2017) using the code from the official GitHub page [^11] https://github.com/HannesStark/protein-localization.
-Light attention + TTT.
-When applying TTT to Light attention, we treat the frozen ESM-1b as the backbone $f$, the language modeling head of ESM-1b as the self-supervised head $g$, and the Light attention block as the fine-tuning head $h$.
-[…]
-[] Learning rate Batch size Grad.
-acc.
-steps TTT steps LoRA rank $r$ LoRa $\alpha$ Fitness prediction ESM2 (35M) + TTT 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32, 64 30 - - ESM2 (650M) + TTT 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - - SaProt (35M) + TTT 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - - SaProt (650M) + TTT 4e-5, 4e-4, 4e-3 2, 4 4, 8, 16, 32 30 - - Structure prediction ESMFold + TTT 4e-4 4 4, 8, 32, 64 30 (max pLDDT) 4, 8, 32 8, 16, 32 ESM3 + TTT 1e-4, 4e-4, 1e-3 2 1, 4, 16 50 (max pLDDT) - - Function prediction TerpeneMiner + TTT 4e-4, 1e-3 2 2, 4, 8 30 - - Light attention + TTT 4e-4, 1e-3, 3e-3 2 2, 4 30 - -
-Table 4: Hyperparameters used for adapting TTT to individual models.
+Light attention + ProteinTTT.
+When applying ProteinTTT to Light attention, we treat the frozen ESM-1b as the backbone $f$, the language modeling head of ESM-1b as the self-supervised head $g$, and the Light attention block as the fine-tuning head $h$.
+Table A3: Hyperparameters used for adapting ProteinTTT to individual models.
 The optimal hyperparameters were estimated using validation datasets corresponding to each of the considered tasks: Fitness prediction, Structure prediction, and Function prediction.
@@
 Low-rank adaptation (LoRA) was only used with ESMFold, containing 3 billion parameters in the ESM2 backbone.
-Appendix B Extended results
-In this section, we provide additional results on test sets (Section B.1) and discuss validation performance (Section B.2).
-B.1 Detailed test performance
+Please note that we did not tune the number of customization steps, as adjusting the learning rate and batch size effectively controls the expected performance under the fixed number of steps, as shown in Figure˜A10.
+Therefore, we used 30 steps in all our experiments.
+The only exception was ESM3 + ProteinTTT, where the number of steps was set to 50 during initial experiments with different models/tasks conducted in parallel before standardizing the number of steps to 30.
+Methods marked with an asterisk (“*”) used a slightly different calculation of the loss function.
+Specifically, the loss was propagated over all tokens, including special and non-masking tokens, while averaging the loss across all tokens simultaneously, rather than first averaging over sequences.
+This approach was used in the early stages of development, and we provide it in our codebase via loss_kind = ‘‘unnormalized_cross_entropy’’.
+Please note that MSA Transformer always uses 1 MSA in a batch and the “Batch size” represents the number of sequences in this MSA with the target sequence always present as the first one.
+Learning rate Batch size Grad. acc. steps Steps (Conf. func. $c$) LoRA rank $r$ LoRa $\alpha$
+Fitness prediction
+ESM2 (35M) + ProteinTTT * 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32, 64 30 - -
+ESM2 (650M) + ProteinTTT * 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - -
+SaProt (35M) + ProteinTTT * 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - -
+SaProt (650M) + ProteinTTT * 4e-5, 4e-4, 4e-3 2, 4 4, 8, 16, 32 30 - -
+ProSST (K=2048) + ProteinTTT * 1e-5, 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - -
+[…]
+MSA Transformer + ProteinTTT 1e-6, 3e-6, 1e-5, 3e-5, 1e-4 1, 4, 8, 16, 32 1, 2, 4, 8 30 - -
+Structure prediction
+ESMFold + ProteinTTT 4e-4 4 4, 8, 32, 64 30 (pLDDT) 4, 8, 32 8, 16, 32
+HelixFold-Single + ProteinTTT 4e-4, 1e-3 4, 8, 16 1 30 (pLDDT) - -
+ESM3 + ProteinTTT 1e-4, 4e-4, 1e-3 2 1, 4, 16 50 (pLDDT) - -
+Function prediction
+EnzymeExplorer + ProteinTTT 4e-4, 1e-3 2 2, 4, 8 30 - -
+Light attention + ProteinTTT 4e-4, 1e-3, 3e-3 2 2, 4 30 - -
+Appendix F Case study details
+F.1 Modeling antibody-antigen loops
+We download the SAbDab dataset from the official website [^12] https://opig.stats.ox.ac.uk/webapps/sabdab-sabpred/sabdab(Dunbar et al., 2014).
+We apply ProteinTTT to targets with low-confidence ESMFold predictions (pLDDT < 70) and remove sequences longer than 400 residues due to GPU memory limitations.
+This results in a final set of 175 antibody and 814 antigen chains.
+We predict the full structures using ESMFold+ProteinTTT (with the same hyperparameters tuned on the CAMEO validation set specified in Table˜A3) and compute LDDT scores against the corresponding PDB structures to assess local errors, which are particularly relevant for loop regions.
+For antibodies, we evaluate the complete structures, while for complementarity-determining regions (CDRs), we extract the CDR substructures as annotated in SAbDab according to Chothia numbering (Chothia & Lesk, 1987) and calculate LDDT on these regions.
+F.2 Expanding known structures of viral proteins
+We use BFVD version archived/2023_02_v2131313https://bfvd.steineggerlab.workers.dev.
+This version contains maximum-pLDDT structures from predictions generated by two strategies: (i) ColabFold (Mirdita et al., 2022) with MSAs constructed using Logan (Chikhi et al., 2024), and (ii) ColabFold with 12 additional recycle steps and MSAs constructed using Logan.
+In Figure˜5, we also report pLDDT values for BFVD version archived/2023_02_v1, where structures are simply obtained from ColabFold with MSAs from Logan, i.e., strategy (i).
+We re-predict structures using ESMFold and ESMFold+ProteinTTT for sequences with length < 450 due to GPU memory constraints.
+We use the same hyperparameters tuned on the CAMEO validation set, as specified in Table˜A3, with the exception of 20 instead of 30 steps for computational efficiency.
+Appendix G Extended results
+In this section, we provide additional results on test sets (Section˜G.1), discuss validation performance (Section˜G.2), and analyze the runtime performance of customization (Section˜G.3).
+G.1 Detailed test performance
 In this section, we provide details on the test performance.
-Specifically, Table 5 shows that test-time training (TTT) primarily enhances performance on challenging targets, characterized by a low number of similar proteins in sequence databases, as measured by MSA depth.
-Additionally, we provide an example illustrating how TTT substantially improves the correlation between ESM2-predicted fitness and ground-truth stability by better identifying disruptive mutations within the protein core (Figure 8).
-Next, Figure 9 shows the distribution of TTT effects: in many cases, TTT has minimal impact on performance; often, it leads to substantial improvements; and in rare cases TTT results in a decrease in performance.
-This positions TTT as a method for enhancing prediction accuracy, while a comprehensive analysis of its failure modes remains an important direction for future research.
-While we demonstrate these effects using a protein folding example, we observe a similar distribution of TTT impact across the tasks.
-We also observe that the overall trend of TTT generally leads to improved performance, with robust consistency across random seeds.
-However, the progression of the performance curve can be rugged, particularly in classification tasks, where substantial changes in the underlying representations are required to shift the top-predicted class in the discrete probability distribution (Figure 10).
-B.2 Validation performance
-This section discusses the performance of test-time training (TTT) on validation data.
-Table 6 illustrates the validation performance of all tested methods for fitness prediction on our newly constructed MaveDB dataset.
-TTT enhances the performance of all the methods.
-The primary focus of the section is hyperparameter tuning.
-Table 4 provides the grid of hyperparameters explored for each model and its size.
-Figure 11 demonstrates the trend of hyperparameter tuning with optimal hyperparameter combination balancing underfitting and overfitting to a single test protein.
-While most hyperparameter configurations lead to overall improvements when using TTT, poorly chosen hyperparameters can have detrimental effects due to rapid overfitting.
-However, with a reliable predicted confidence measure, such as pLDDT, the appropriate TTT step can be selected to mitigate overfitting.
-Figure 12 demonstrates that when using ESM3 + TTT with pLDDT-based step selection for protein folding, all hyperparameter configurations result in improved performance compared to the base ESM3 model.
-Figure 8: Example of test-time training (TTT) applied to fitness prediction.
+Specifically, Table˜A4 shows that customization with ProteinTTT primarily enhances performance on challenging targets, characterized by a low number of similar proteins in sequence databases, as measured by MSA depth.
+Additionally, we provide a qualitative example illustrating how ProteinTTT substantially improves the correlation between ESM2-predicted fitness and ground-truth stability by better identifying disruptive mutations in the protein core (Figure A5).
+Next, Figure˜A6 shows the distribution of ProteinTTT effects: in many cases, customization has minimal impact on performance; often, it leads to substantial improvements; and in rare cases, customization results in a decrease in performance.
+This positions ProteinTTT as a method for enhancing prediction accuracy, while a comprehensive analysis of its failure modes remains an important direction for future research.
+While we demonstrate these effects using a protein folding example, we observe a similar distribution of ProteinTTT impact across the tasks.
+We also observe that the overall trend of customization with ProteinTTT generally leads to improved performance, with robust consistency across random seeds.
+However, the progression of the performance curve can be rugged, particularly in classification tasks, where substantial changes in the underlying representations are required to shift the top-predicted class in the discrete probability distribution (Figure˜A8).
+G.2 Validation performance
+This section discusses the performance of ProteinTTT on validation data.
+Table˜A5 illustrates the validation performance of the tested methods for fitness prediction on our newly constructed MaveDB dataset.
+ProteinTTT enhances the performance of all the methods.
+Next, we discuss the hyperparameter optimization.
+Table˜A3 provides the grid of hyperparameters explored for each model and its size, as well as specifies the optimal hyperparameters suitable for downstream applications.
+Figure˜A10 demonstrates the trend of hyperparameter tuning with optimal hyperparameter combination balancing underfitting and overfitting to a single target protein.
+While most of reasonable hyperparameter configurations lead to overall improvements when using customization with ProteinTTT, poorly chosen hyperparameters can have detrimental effects due to rapid overfitting.
+However, with a reliable predicted confidence measure, such as pLDDT, the appropriate customization step for each protein can be selected to mitigate overfitting.
+Figure˜A11 demonstrates that when using ESM3 + ProteinTTT with pLDDT-based step selection for protein structure prediction, all hyperparameter configurations result in improved performance compared to the base ESM3 model.
+G.3 Runtime performance
+In this section, we demonstrate that customization with ProteinTTT can be done efficiently, with an acceptable computational overhead.
+Specifically, we show that ESMFold, known for being a faster alternative to more performant methods such as AlphaFold2 (Jumper et al., 2021) or AlphaFold3 (Abramson et al., 2024), still remains in the category of lightweight methods even with ProteinTTT customization (Figure˜A4).
+This observation highlights the practical utility of ProteinTTT.
+For example, ESMFold enabled structural characterization of large metagenomics data (>617 million metagenomic sequences), which would be infeasible with AlphaFold2 (Lin et al., 2023).
+Nevertheless, the original ESMFold has high confidence predictions only for 36% of sequences from the metagenomic database, while the other 392 million sequences remain with low or medium confidence predictions.
+At the same time, ESMFold + ProteinTTT enables more accurate predictions compared to the original ESMFold (Figure˜A6 suggests that ESMFold + ProteinTTT significantly improves predictions in almost 40% of challenging sequences).
+It means that applying ESMFold + ProteinTTT to these remaining sequences could significantly expand the metagenomic atlas characterized by ESMFold.
+Here, we illustrate this on a similar case study by applying ESMFold + ProteinTTT to more than 300 thousand viral proteins in BFVD (Section˜5.2)
+Table A4: ProteinTTT performance on ProteinGym depending on MSA depth.
+MSA depth reflects the number of available proteins similar to the target protein and, when using large protein language models, can be interpreted as a measure of the representation of similar proteins in the training data (Section˜E.2.2).
+Customization with ProteinTTT primarily improves performance on difficult targets, with low MSA depth.
+Standard deviations are calculated over 5 random seeds but are omitted in the right panel for brevity, where the maximum standard deviation does not exceed 0.0004.
+Avg. Spearman $\uparrow$
+Spearman by MSA depth $\uparrow$
+Low depth Medium depth High depth
+ESM2 (35M) (Lin et al., 2023) 0.3211 0.2394 0.2707 0.451
+ESM2 (35M) + ProteinTTT (Ours) 0.3407 ± 0.00014 0.2445 0.3144 0.4598
+SaProt (35M) (Su et al., 2023) 0.4062 0.3234 0.3921 0.5057
+SaProt (35M) + ProteinTTT (Ours) 0.4106 ± 0.00004 0.3253 0.3972 0.5091
+ESM2 (650M) (Lin et al., 2023) 0.4139 0.3346 0.4063 0.5153
+ESM2 (650M) + ProteinTTT (Ours) 0.4153 ± 0.00003 0.3363 0.4126 0.5075
+SaProt (650M) (Su et al., 2023) 0.4569 0.3947 0.4502 0.5448
+SaProt (650M) + ProteinTTT (Ours) 0.4583 ± 0.00001 0.3954 0.4501 0.5439
+ProSST (K=2048) (Li et al., 2024) 0.5068 0.4731 0.5107 0.5749
+ProSST (K=2048) + ProteinTTT (Ours) 0.5087 ± 0.00004 0.4809 0.5104 0.5750
+Table A5: Performance of ProteinTTT on the MaveDB dataset.
+In this work, we use our newly constructed MaveDB dataset as a validation fold for tuning the ProteinTTT hyper-parameters for fitness prediction.
+For computational efficiency, we only select a subset of 50 proteins (Section˜E.2.1) and do not run customization across multiple random seeds to estimate standard deviations.
+The performance shown was calculated by first aggregating correlations per assay, and then per protein (some assays correspond to the same protein).
+Avg. Spearman $\uparrow$
+ESM2 (35M) (Lin et al., 2023) 0.4458
+ESM2 (35M) + ProteinTTT (Ours) 0.4593
+ESM2 (650M) (Lin et al., 2023) 0.4568
+ESM2 (650M) + ProteinTTT (Ours) 0.4604
+SaProt (650M) (Su et al., 2023) 0.4926
+SaProt (650M) + ProteinTTT (Ours) 0.4926
+SaProt (35M) (Su et al., 2023) 0.5251
+SaProt (35M) + ProteinTTT (Ours) 0.5271
+ProSST (K=2048) (Li et al., 2024) 0.5444
+ProSST (K=2048) + ProteinTTT (Ours) 0.5462
+Figure A4: Running time of ESMFold + ProteinTTT.
+For ESMFold and its variants, the median and interquartile ranges of running times on the CAMEO test set are shown using a single NVIDIA A100 GPU.
+For AlphaFold2, we use estimates from Lin et al.
+(2023).
+Specifically, a forward pass through AlphaFold2 is approximately 60 times more computationally expensive than ESMFold (e.g., AlphaFold2 (no MSA; estimate): $2\times 60=120$ seconds), with additional MSA construction taking at least 10 minutes using standard pipelines (AlphaFold2 (estimate): $2\times 60+10\times 60=720$ seconds).
+ESMFold + ProteinTTT (30 steps) involves LoRA parameter updates, along with forward passes at each customization step to estimate pLDDT and select the structure with the highest predicted confidence.
+Disabling pLDDT significantly reduces computational overhead (ESMFold + ProteinTTT (no pLDDT) compared to ESMFold + ProteinTTT), but may require careful parameter tuning (Section˜G.2).
+Overall, ESMFold + ProteinTTT maintains the speed advantage of ESMFold, and is at least an order of magnitude faster than AlphaFold2.
+Figure A5: Example of protein fitness prediction upon single-sequence model customization with ProteinTTT.
 Fitness predictions from ESM2 (650M) show poor correlation with experimental fitness values in the ProteinGym test set measured by the stability assay “UBR5_HUMAN_Tsuboyama_2023_1I2T” (Tsuboyama et al., 2023) (left).
-ESM2 + TTT achieves significantly higher correlation, likely due to improved detection of disruptive mutations in the protein core that impact protein stability (middle).
-The ground-truth fitness data aligns with the TTT-enhanced model, showing that residues crucial for stability (i.e., having negative mean fitness) are concentrated in the protein core (right).
+ESM2 + ProteinTTT achieves significantly higher correlation, likely due to improved detection of disruptive mutations in the protein core that impact protein stability (middle).
+The ground-truth fitness data aligns with the customized model, showing that residues crucial for stability (i.e., having negative mean fitness) are concentrated in the protein core (right).
 Residue colors represent the mean fitness upon all single-point substitutions (with the exception of several missing mutations in the ground-truth data), with red indicating residues where mutations have detrimental effects on average.
-[…]
-[] Avg.
-Spearman $\uparrow$ Spearman by MSA depth $\uparrow$ Low depth Medium depth High depth ESM2 (35M) (Lin et al., 2023) 0.3211 0.2394 0.2707 0.451 ESM2 (35M) + TTT (Ours) 0.3407 ± 0.00014 0.2445 0.3144 0.4598 SaProt (35M) (Su et al., 2023) 0.4062 0.3234 0.3921 0.5057 SaProt (35M) + TTT (Ours) 0.4106 ± 0.00004 0.3253 0.3972 0.5091 ESM2 (650M) (Lin et al., 2023) 0.4139 0.3346 0.4063 0.5153 ESM2 (650M) + TTT (Ours) 0.4153 ± 0.00003 0.3363 0.4126 0.5075 SaProt (650M) (Su et al., 2023) 0.4569 0.3947 0.4502 0.5448 SaProt (650M) + TTT (Ours) 0.4583 ± 0.00001 0.3954 0.4501 0.5439
-Table 5: Test-time training (TTT) performance on ProteinGym depending on MSA depth.
-MSA depth reflects the number of available proteins similar to the target protein and, when using large protein language models, can be interpreted as a measure of the representation of similar proteins in the training data (Section A.1.2).
-TTT primarily improves performance on difficult targets, with low MSA depth.
-Standard deviations are calculated over 5 random seeds but are omitted in the right panel for brevity, where the maximum standard deviation does not exceed 0.0004.
-[…]
-[] Avg.
-Spearman $\uparrow$ ESM2 (35M) (Lin et al., 2023) 0.4458 ESM2 (35M) + TTT (Ours) 0.4593 ESM2 (650M) (Lin et al., 2023) 0.4568 ESM2 (650M) + TTT (Ours) 0.4604 SaProt (650M) (Su et al., 2023) 0.4926 SaProt (650M) + TTT (Ours) 0.4926 SaProt (35M) (Su et al., 2023) 0.5251 SaProt (35M) + TTT (Ours) 0.5271
-Table 6: Performance of test-time training (TTT) on the MaveDB dataset.
-In this work, we use our newly constructed MaveDB benchmark as a validation fold for tuning the hyper-parameters of TTT for fitness prediction.
-For computational efficiency, we only select a subset of 50 proteins (Section A.1.1) and do not run TTT across multiple random seeds to estimate standard deviations.
-The performance shown was calculated by first aggregating correlations per assay, and then per protein (some assays correspond to the same protein).
-Figure 9: Per-protein performance of ESMFold + TTT and ESM3 + TTT on the CAMEO test set.
-The y-axis shows the change in TM-score after applying test-time training (TTT), with higher values indicating improvement.
+Figure A6: Per-protein performance of ESMFold + ProteinTTT and ESM3 + ProteinTTT on the CAMEO test set.
+The y-axis shows the change in TM-score after applying customization with ProteinTTT, with higher values indicating improvement.
 The x-axis represents performance across five random seeds.
-The red dashed line marks no change in TM-score (TM-score difference $=0$), and the pink band represents minor changes in TM-score ($-0.05<$ TM-score difference $<0.05$), which we do not consider significant.
+The red dashed line marks no change in TM-score (TM-score difference =0=0), and the pink band represents minor changes in TM-score (0.05<-0.05< TM-score difference <0.05<0.05), which we do not consider significant.
 Each point in the swarm plot corresponds to a single protein from the CAMEO test set.
-On average, applying TTT to ESMFold improves the structure predictions for 7 out of 18 proteins, with 2 showing degradation.
+On average, applying ProteinTTT to ESMFold improves the structure predictions for 7 out of 18 proteins, with 2 showing degradation.
 The rest of the proteins are not significantly affected.
-Similarly, applying TTT to ESM3 results in 6 improvements out of 18 proteins, with 1 case of degradation.
-Figure 10: Test performance of TerpeneMiner + TTT across fine-tuning steps.
+Similarly, applying ProteinTTT to ESM3 results in 6 improvements out of 18 proteins, with 1 case of degradation.
+Figure A7: Test performance of ESMFold + ProteinTTT and ESM3 + ProteinTTT on the CAMEO test set depending on the total number of customization steps.
+The x-axis shows the averaged performance across all test proteins, with error bars representing the standard deviation across five random seeds.
+The y-axis metrics correspond to the structure with the highest pLDDT score up to the given step.
+While an increased number of ProteinTTT steps generally enhances performance, only a few steps (e.g., five) may suffice to achieve significant performance improvement.
+Figure A8: Test performance of EnzymeExplorer + ProteinTTT across customization steps.
 The performance is averaged across all 512 proteins in the dataset, with error bars representing the standard deviation across 5 random seeds.
-Figure 11: Dependence on hyperparameters in test-time training for fitness prediction.
-Each plot shows the progression of Spearman correlation (green) increasing alongside a decrease in perplexity (pink) for each TTT step, averaged across all assays in the MaveDB validation dataset.
-The model used is ESM2 (35M) + TTT and the grid displays the progression for all hyperparameter combinations.
-As the learning rate increases (left to right) and the effective batch size grows, given by the increased number of gradient accumulation steps (top to bottom), the model reaches peak performance more quickly and begins overfitting to a test protein.
-The optimal hyperparameter combination (learning rate = 4e-4, gradient accumulation steps = 16) lies near the center of the grid, balancing between underfitting and overfitting to a test protein.
-Figure 12: Hyperparameter search for protein structure prediction with ESM3 + TTT.
+Figure A9: ESMFold + ProteinTTT pLDDT correlates with ESMFold + ProteinTTT LDDT.
+The evaluation was performed on 17,582 AlphaFold2 reference structures from the BFVD database with pLDDT > 90.
+Here, $r=0.875$ denotes the Pearson correlation coefficient.
+Figure A10: Dependence on ProteinTTT hyperparameters for customized fitness prediction.
+Each plot shows the progression of Spearman correlation (green) increasing alongside a decrease in perplexity (pink) for each customization step, averaged across all assays in the MaveDB validation dataset.
+The model used is ESM2 (35M) + ProteinTTT, and the grid displays the combinations of different numbers of gradient accumulation steps (i.e., effective batch sizes; shown in rows, increasing from top to bottom) and learning rates (columns, increasing from left to right).
+As the learning rate increases and the number of gradient accumulation steps grows, the model reaches peak performance more quickly but begins to overfit to a target protein.
+The optimal hyperparameter combination (learning rate = 4e-4, gradient accumulation steps = 16) lies near the center of the grid, balancing between underfitting and overfitting to a target protein.
+Notably, the figure demonstrates that, although ProteinTTT involves three main hyperparameters (batch size, learning rate, and the number of steps), there are effectively only two degrees of freedom controlling the performance of the model.
+In other words, by keeping the number of steps constant (e.g., 30), the expected performance can be controlled by adjusting the learning rate and the batch size.
+Figure A11: Hyperparameter search for protein structure prediction with ESM3 + ProteinTTT.
 We conducted a comprehensive grid search based on three key hyperparameters: learning rate (denoted as “lr”), number of gradient accumulation steps (denoted as “grad_steps”; with the batch size of two), and masking strategy (denoted as “mask”).
-We explored two learning rates, 4e-4 and 1e-3, three gradient accumulation step values of 1, 4, and 16, and five different masking strategies: uniform sampling of 0.05, 0.5, and 1.0 fractions of amino acids, as well as the beta30 and betalinear30 distributions proposed in the ESM3 paper (Hayes et al., 2024).
+We explored two learning rates, 4e-4 and 1e-3, three gradient accumulation step values of 1, 4, and 16, and five different masking strategies: uniform sampling of 0.05, 0.5, and 1.0 fractions of amino acids, as well as the “beta30” and “betalinear30” distributions proposed in the ESM3 paper (Hayes et al., 2024).
 Each row in the table presents the mean TM-score and LDDT metrics with standard deviations across five random seeds on the CAMEO validation fold.
-The last row, denoted as “No TTT”, shows the performance of ESM3 without TTT.
-The results indicate that ESM3 + TTT is robust to the choice of hyperparameters and consistently outperforms the base model across all configurations.
-We selected the configuration from the last row to compute the results on the test fold.
-For the hyperparameter search, we used 30 TTT steps instead of 50 to reduce computation time.
+The last row, denoted as “No ProteinTTT”, shows the performance of ESM3 without customization.
+The results indicate that ESM3 + ProteinTTT is robust to the choice of hyperparameters and consistently outperforms the base model across all configurations.
+We selected the configuration from the last row (excluding “No ProteinTTT”) to compute the results on the test fold.
+For the hyperparameter search, we used 30 customization steps instead of 50 to reduce computation time.

### v3

@@
 ESMFold poorly predicts the structure of the CASP14 target T1074 (white) because the underlying language model ESM2 poorly fits the sequence, as indicated by the high perplexity (left and Fig.
-2E in Lin et al.
-(2023)).
+2E in lin2023evolutionary).
 Self-supervised test-time customization of ESM2 to the single sequence of T1074 reduces the perplexity, resulting in improved structure prediction (right).
-A comprehensive understanding of protein structure, function, and fitness is essential for advancing research in the life sciences (Subramaniam & Kleywegt, 2022; Tyers & Mann, 2003; Papkou et al., 2023).
-While machine learning models have shown remarkable potential in protein research, they are typically optimized for achieving the best average performance across large datasets (Jumper et al., 2021; Watson et al., 2023; Kouba et al., 2023).
-However, biologists often focus their research on individual proteins or protein complexes involved in, for example, metabolic disorders (Ashcroft et al., 2023; Gunn & Neher, 2023), oncogenic signaling (Hoxhaj & Manning, 2020; Keckesova et al., 2017), neurodegeneration (Gulen et al., 2023; oh Seo et al., 2023), and other biological phenomena (Gu et al., 2022).
+A comprehensive understanding of protein structure, function, and fitness is essential for advancing research in the life sciences (subramaniam2022paradigm; tyers2003from; papkou2023rugged).
+While machine learning models have shown remarkable potential in protein research, they are typically optimized for achieving the best average performance across large datasets (jumper2021highly; watson2023novo; kouba2023machine).
+However, biologists often focus their research on individual proteins or protein complexes involved in, for example, metabolic disorders (ashcroft2023glucokinase; gunn2023structure), oncogenic signaling (hoxhaj2020pi3k; keckesova2017lactb), neurodegeneration (gulen2023cgas; seo2023apoe), and other biological phenomena (gu2022sestrin).
 In these scenarios, detailed insights into a single protein can lead to significant scientific advances.
-However, general machine learning models for proteins often struggle to generalize to practically interesting individual cases due to data scarcity (Bushuiev et al., 2023; Chen & Gong, 2022) and distribution shifts (Škrinjar et al., 2025; Tagasovska et al., 2024; Feng et al., 2024).
-Bridging the gap between broad, dataset-wide optimization and precision needed to study single proteins of practical interest remains a key challenge in integrating machine learning into biological research (Sapoval et al., 2022).
-This challenge is particularly acute in computational biology, where accurate predictions for individual proteins are essential to guide resource-intensive wet-lab experiments, in contrast to domains such as natural language processing or computer vision, where models are typically expected to flexibly handle diverse prompts from many users in real time (Brown, 2020; Ramesh et al., 2021).
+However, general machine learning models for proteins often struggle to generalize to practically interesting individual cases due to data scarcity (bushuiev2023learning; chen2022hotprotein) and distribution shifts (vskrinjar2025have; tagasovska2024antibody; feng2024deep).
+Bridging the gap between broad, dataset-wide optimization and precision needed to study single proteins of practical interest remains a key challenge in integrating machine learning into biological research (sapoval2022current).
+This challenge is particularly acute in computational biology, where accurate predictions for individual proteins are essential to guide resource-intensive wet-lab experiments, in contrast to domains such as natural language processing or computer vision, where models are typically expected to flexibly handle diverse prompts from many users in real time (brown2020language; ramesh2021zero).
 To address this challenge, we propose a test-time approach for generalization to one protein at a time, effectively enabling more accurate predictions for individual targets, particularly those poorly represented in training data.
@@
 The Y-shaped paradigm of learning.
-In machine learning applied to proteins, architectures often follow a Y-shaped paradigm (Gandelsman et al., 2022), consisting of a backbone feature extractor $f$ operating on protein tokens $x$, a self-supervised head $g$, and an alternative fine-tuning head $h$.
+In machine learning applied to proteins, architectures often follow a Y-shaped paradigm (gandelsman2022test), consisting of a backbone feature extractor $f$ operating on protein tokens $x$, a self-supervised head $g$, and an alternative fine-tuning head $h$.
 During training, $g\circ f$ is first pre-trained, and the pre-trained backbone $f$ is then reused to fine-tune $h\circ f$ toward a downstream task.
@@
 Generalization is achieved by leveraging the rich knowledge encoded in the backbone $f$ and the task-specific priors embedded in the fine-tuning head $h$.
-This paradigm enables overcoming data scarcity during fine-tuning and underlies breakthrough approaches in protein structure prediction (Lin et al., 2023), protein design (Watson et al., 2023), protein function prediction (Yu et al., 2023), and other tasks (Hayes et al., 2024).
-The backbone $f$ is typically a large neural network pre-trained in a self-supervised way on a large dataset using a smaller pre-training projection head $g$ (Hayes et al., 2024).
+This paradigm enables overcoming data scarcity during fine-tuning and underlies breakthrough approaches in protein structure prediction (lin2023evolutionary), protein design (watson2023novo), protein function prediction (yu2023enzyme), and other tasks (hayes2024simulating).
+The backbone $f$ is typically a large neural network pre-trained in a self-supervised way on a large dataset using a smaller pre-training projection head $g$ (hayes2024simulating).
 The fine-tuning head $h$, however, depends on the application.
-In some cases, $h$ is a large neural network, repurposing the pre-trained model entirely (Watson et al., 2023; Lin et al., 2023); in others, $h$ is a minimal projection with few parameters (Cheng et al., 2023), or even without any parameters at all (i.e., a zero-shot setup; Meier et al.
-(2021); Dutton et al.
-(2024)).
-The fine-tuning head $h$ can also be a machine learning algorithm other than a neural network (Samusevich et al., 2025).
+In some cases, $h$ is a large neural network, repurposing the pre-trained model entirely (watson2023novo; lin2023evolutionary); in others, $h$ is a minimal projection with few parameters (cheng2023accurate), or even without any parameters at all (i.e., a zero-shot setup; meier2021language; dutton2024improving).
+The fine-tuning head $h$ can also be a machine learning algorithm other than a neural network (samusevich2024discovery).
 Masked modeling.
@@
 Nevertheless, the dominant paradigm for protein pre-training is masked modeling, which optimizes model weights to reconstruct missing protein parts.
-This objective has proven effective across diverse tasks (Heinzinger & Rost, 2025; Schmirler et al., 2024), including structure (Lin et al., 2023; Jumper et al., 2021), fitness (Meier et al., 2021; Su et al., 2023), and function prediction (Samusevich et al., 2025; Yu et al., 2023; Elnaggar et al., 2021), as well as protein design (Hayes et al., 2024), and has been successfully applied to various protein representations such as sequences (Hayes et al., 2024; Elnaggar et al., 2023), graphs (Dieckhaus et al., 2024; Bushuiev et al., 2023), and voxels (Diaz et al., 2023).
+This objective has proven effective across diverse tasks (heinzinger2025teaching; schmirler2024fine), including structure (lin2023evolutionary; jumper2021highly), fitness (meier2021language; su2023saprot), and function prediction (samusevich2024discovery; yu2023enzyme; elnaggar2021prottrans), as well as protein design (hsieh2025elucidating; hayes2024simulating; nijkamp2023progen2), and has been successfully applied to various protein representations such as sequences (hayes2024simulating; elnaggar2023ankh), graphs (dieckhaus2024transfer; bushuiev2023learning), and voxels (diaz2023stability).
 Model customization.
-Several studies have shown that machine learning models for proteins benefit from being fine-tuned on protein-specific  (Notin et al., 2024; Kirjner et al., 2023; Rao et al., 2019) or protein family-specific (Sevgen et al., 2023; Samusevich et al., 2025) data.
-However, collecting additional data may be resource-intensive, and for many targets, relevant datasets or proteins may be limited or not available (Durairaj et al., 2023; Kim et al., 2025).
+Several studies have shown that machine learning models for proteins benefit from being fine-tuned on protein-specific  (notin2024proteingym; kirjner2023improving; rao2019evaluating) or protein family-specific (sevgen2023prot; samusevich2024discovery) data.
+However, collecting additional data may be resource-intensive, and for many targets, relevant datasets or proteins may be limited or not available (durairaj2023uncovering; kim2025bfvd).
 In this paper, we propose a versatile method enabling customizing PLMs for a single target protein or its MSA in a self-supervised manner, on the fly, and without assuming any additional data.
-Customization methods have been developed in computer vision (Chi et al., 2024; Wang et al., 2023; Xiao et al., 2022; Karani et al., 2021) and natural language processing (Hübotter et al., 2024; Hardt & Sun, 2023; Ben-David et al., 2022; Banerjee et al., 2021).
-The paradigm of test-time training (TTT), developed to mitigate distribution shifts in computer vision applications (Gandelsman et al., 2022; Sun et al., 2020), is the main inspiration for our work.
+Customization methods have been developed in computer vision (chi2024adapting; wang2023test; xiao2022learning; karani2021test) and natural language processing (hubotter2024efficiently; hardt2023test; ben2022pada; banerjee2021self).
+The paradigm of test-time training (TTT), developed to mitigate distribution shifts in computer vision applications (gandelsman2022test; sun2020test), is the main inspiration for our work.
 We demonstrate that customization via test-time training enhances the accuracy of PLMs across a wide range of downstream tasks even without the presence of explicit distribution shifts.
@@
 Figure˜2a illustrates our customization approach, which is summarized in the following sections.
-Appendix˜B describes the extension of our method to customization using a MSA of a protein, rather than its single sequence.
+Appendix˜C describes the extension of our method to customization using a MSA of a protein, rather than its single sequence.
 Customization training objective.
-We customize $g\circ f$ to a single target protein sequence $x$ via minimizing the masked language modeling objective (Devlin, 2018; Rives et al., 2021):
+We customize $g\circ f$ to a single target protein sequence $x$ via minimizing the masked language modeling objective (devlin2018bert; rives2021biological):
 $\mathcal{L}(x;\theta)=\mathbb{E}_{M\sim p_{\text{mask}}(M)}\Biggl[\sum_{i\in M}{-\log{p(x_{i}|x_{\setminus M};\theta)}}\Biggl],$ (2)
@@
 The objective function $\mathcal{L}(x;\theta)$ maximizes the log-probabilities $\log{p(x_{i}|x_{\setminus M};\theta)}\ \dot{=}\ g(f(x_{\setminus M};\theta))_{i}$ of the true (i.e., wild-type) tokens $x_{i}$ at the masked positions $i\in M$ in the partially masked sequence $x_{\setminus M}$, where $\theta$ denotes the parameters of the backbone $f$, and $g$ is the masked language modeling head.
-Please note that here we focus on bi-directional masked modeling models, which employ random masking, but the method can be easily extended to models employing autoregressive masking.
-To ensure consistency between the customization and pre-training, ProteinTTT adopts the same masking and data preprocessing strategies used during pre-training.
-Specifically, $p_{\text{mask}}(M)$ can follow different distributions, such as sampling a fixed proportion (e.g., 15%) of random amino acid tokens (Lin et al., 2023), or dynamically varying the number of sampled tokens based on another distribution (e.g., a beta distribution; Hayes et al.
-(2024)).
-During the customization, we replicate the masking distribution used during the pre-training.
-We also replicate other pre-training practices, such as replacing 10% of masked tokens with random tokens and another 10% with the original tokens (Devlin, 2018; Lin et al., 2023; Su et al., 2023) or cropping sequences to random 1024-token fragments (Lin et al., 2023; Su et al., 2023).
+While we focus on classical bi-directional masked modeling, we also demonstrate that ProteinTTT can be similarly applied to autoregressive and discrete diffusion models (Appendix˜B).
+To ensure consistency between customization and pre-training, ProteinTTT adopts the same masking and preprocessing strategies used during pre-training.
+Specifically, $p_{\text{mask}}(M)$ can follow different distributions, such as sampling a fixed ratio (e.g., 15%) of random tokens (lin2023evolutionary), or dynamically varying the number of sampled tokens based on another distribution (e.g., a beta distribution; hayes2024simulating).
+During customization, we replicate the masking distribution used in pre-training.
+We also replicate other pre-training practices, such as replacing 10% of masked tokens with random tokens and another 10% with the original tokens (devlin2018bert; lin2023evolutionary; su2023saprot) or cropping sequences to random 1024-token fragments (lin2023evolutionary; su2023saprot).
 Optimization.
@@
 If $c$ is not available, we set $\theta_{x}=\theta_{T}$.
-Section˜G.2 discusses how using pLDDT as the confidence function $c$ for structure prediction makes ProteinTTT robust to hyperparameter selection and how the number of steps $T$ can be fixed (e.g., $T=30$) while optimizing learning rate and batch size effectively.
+Section˜H.2 discusses how using pLDDT as the confidence function $c$ for structure prediction makes ProteinTTT robust to hyperparameter selection and how the number of steps $T$ can be fixed (e.g., $T=30$) while optimizing learning rate and batch size effectively.
 Before customizing for the next target protein, the parameters are reset to $\theta_{0}$.
-To make ProteinTTT easily applicable to large-scale models (e.g., the 3B-parameter ESM2 backbone), we leverage low-rank adaptation (LoRA; Hu et al.
-(2021)) and gradient accumulation during customization.
-Additionally, to improve the stability and predictability of customization, we use stochastic gradient descent (SGD; Ruder (2016)) instead of the commonly used Adam optimizer (Kingma & Ba, 2015), following (Gandelsman et al., 2022).
-Further details are provided in Appendix˜E.
+To make ProteinTTT easily applicable to large-scale models (e.g., the 3B-parameter ESM2 backbone), we leverage low-rank adaptation (LoRA; hu2021lora) and gradient accumulation during customization.
+Additionally, to improve the stability and predictability of customization, we use stochastic gradient descent (SGD; ruder2016overview) instead of the commonly used Adam optimizer (kingma2015adam), following (gandelsman2022test).
+Further details are provided in Appendix˜F.
 Figure 3: Customization with ProteinTTT improves protein structure prediction by reducing protein sequence perplexity.
@@
 In this work, we consider three standard problems: protein structure, fitness, and function prediction, and apply our method to corresponding well-established models.
-For structure prediction, we apply ProteinTTT to ESMFold (Figure˜3, Lin et al.
-(2023), HelixFold-Single (Fang et al., 2023), and ESM3 (Hayes et al., 2024); for fitness prediction, we use ESM2 (Lin et al., 2023), SaProt (Su et al., 2023), ProSST (Li et al., 2024), and MSA Transformer (Rao et al., 2021); and for function prediction, we apply ProteinTTT to ESM-1v-based (Meier et al., 2021) EnzymeExplorer (Samusevich et al., 2025) and ESM-1b-based (Rives et al., 2021) Light attention (Stärk et al., 2021).
-In all models we consider, $f$ is a Transformer encoder that takes protein tokens as input, and $g$ is a masked language modeling head (a layer mapping token embeddings to amino acid types).
-The downstream task heads $h$ vary strongly across tasks.
-For structure prediction, $h$ is a protein structure predictor: in ESMFold and HelixFold-Single, it is an AlphaFold2-inspired module (Jumper et al., 2021), while in ESM3, it is a VQ-VAE structure decoder (Razavi et al., 2019).
-For fitness prediction, $h$ outputs a single score per sequence; ESM2, SaProt, and ProSST perform zero-shot inference using $h\circ f$ via log odds from $g$, with $h$ functioning as a simple adaptation of $g$ without introducing extra parameters.
-The function predictors are classification models: in EnzymeExplorer (Samusevich et al., 2025), $h$ is a random forest that outputs substrate probabilities, and in Light attention (Stärk et al., 2021), $h$ is a light attention module predicting protein localization classes within a cell.
+For structure prediction, we apply ProteinTTT to ESMFold (Figure˜3, lin2023evolutionary, HelixFold-Single (fang2023method), DPLM2 Bit-based (hsieh2025elucidating), and ESM3 (hayes2024simulating); for fitness prediction, we use ESM2 (lin2023evolutionary), SaProt (su2023saprot), ProSST (li2024prosst), MSA Transformer (rao2021msa), and ProGen2 (nijkamp2023progen2); and for function prediction, we apply ProteinTTT to ESM-1v-based (meier2021language) EnzymeExplorer (samusevich2024discovery) and ESM-1b-based (rives2021biological) Light attention (stark2021light).
+In all models we consider, $f$ is a Transformer encoder operating on protein tokens, and $g$ is a masked language modeling head mapping embeddings to amino acid types.
+The downstream head $h$, however, varies strongly by task.
+For structure prediction, $h$ is a structure predictor: AlphaFold2-inspired modules in ESMFold, HelixFold-Single and DPLM2 Bit-wise (jumper2021highly), and a VQ-VAE decoder in ESM3 (razavi2019generating).
+For fitness prediction, $h$ outputs a single score; all methods perform zero-shot inference using $h\circ f$ via log likelihoods from $g$, with $h$ acting as a simple, parameter-free adaptation of $g$.
+For function prediction, $h$ is a classifier: a random forest in EnzymeExplorer (samusevich2024discovery) and a light attention module in (stark2021light).
 4 Experiments
@@
 […]
-Protein structure prediction is the task of predicting 3D coordinates of protein atoms given the amino acid sequence.
-It is arguably one of the best-established problems in computational biology (Jumper et al., 2021; Lin et al., 2023; Abramson et al., 2024).
+Protein structure prediction is the task of predicting 3D atom coordinates from an amino acid sequence. It is arguably one of the best-established problems in computational biology (jumper2021highly).
 Evaluation setup.
 To evaluate the performance of ProteinTTT, we employ CAMEO, a standard benchmark for protein folding.
-We use the validation and test folds from Lin et al.
-(2023), focusing only on targets with low-confidence predictions from the base ESMFold, as determined by pLDDT and perplexity (Section˜E.1).
-We use the standard TM-score (Zhang & Skolnick, 2004) and LDDT (Mariani et al., 2013) metrics to evaluate global and local structure prediction quality, respectively.
+We use the validation and test folds from lin2023evolutionary, focusing only on targets with low-confidence predictions from the base ESMFold, as determined by pLDDT and perplexity (Section˜F.1).
+We use the standard TM-score (zhang2004scoring) and LDDT (mariani2013lddt) metrics to evaluate global and local structure prediction quality, respectively.
 As baseline methods, we use techniques alternative to ProteinTTT for improving the performance of the pre-trained base models.
-In particular, the ESMFold paper proposes randomly masking 15% of amino acids in a protein sequence before inference, allowing for sampling multiple protein structure predictions from the regression ESMFold model (Lin et al., 2023).
+In particular, the ESMFold paper proposes randomly masking 15% of amino acids in a protein sequence before inference, allowing for sampling multiple protein structure predictions from the regression ESMFold model (lin2023evolutionary).
 For each sequence, we sample a number of predictions equal to the total number of ProteinTTT steps and refer to this baseline as ESMFold + MP (Masked Prediction).
-As a baseline for ESM3, we use chain-of-thought iterative decoding, referred to as ESM3 + CoT, proposed in the ESM3 paper (Hayes et al., 2024).
+As a baseline for ESM3, we use chain-of-thought iterative decoding, referred to as ESM3 + CoT, proposed in the ESM3 paper (hayes2024simulating).
 Results.
 Customization with ProteinTTT consistently improves the performance of all the tested methods, ESMFold, HelixFold-Single, and ESM3, outperforming the masked prediction (ESMFold + MP) and chain-of-thought (ESM3 + CoT) baselines, as shown in Table˜2.
-Among the 18 challenging CAMEO test proteins, ProteinTTT significantly improved the prediction of 7, 5, and 6 structures from ESMFold, HelixFold-Single, and ESM3, respectively, while only slightly disrupting the prediction of 2, 1, and 1 structures, respectively (Figure˜A6).
+Among the 18 challenging CAMEO test proteins, ProteinTTT significantly improved the prediction of 7, 4, 5, and 6 structures from ESMFold, DPLM2 Bit-based, HelixFold-Single, and ESM3, respectively, while only moderately disrupting the prediction of 2, 1, 1, and 1 structures, respectively (Figure˜A7).
+Remarkably, ProteinTTT improves DPLM2 Bit-based despite the absence of a confidence function (no trained pLDDT head available) and despite the model being pretrained via discrete diffusion, while still using the same masked-modeling objective for customization as for the other methods.
 Most notably, ProteinTTT enables accurate structure prediction for targets that are poorly predicted with the original models.
-For instance, Figure˜1 presents a strongly improved structure predicted using ESMFold + ProteinTTT for the target that was part of the CASP14 competition and shown as an unsuccessful case in the original ESMFold publication (Lin et al.
-(2023), Fig.
+For instance, Figure˜1 presents a strongly improved structure predicted using ESMFold + ProteinTTT for the target that was part of the CASP14 competition and shown as an unsuccessful case in the original ESMFold publication (lin2023evolutionary, Fig.
 2E).
@@
 Figure˜A4 shows that ESMFold + ProteinTTT maintains computational efficiency of ESMFold, being an order of magnitude faster than AlphaFold2.
-Figure˜A11 additionally demonstrates the robustness of ESM3 + ProteinTTT to the choice of hyperparameters.
+Figure˜A10 additionally demonstrates the robustness of ESM3 + ProteinTTT to the choice of hyperparameters.
 4.2 Protein Fitness Prediction
@@
 Evaluation Setup.
-We evaluate the models using ProteinGym, the state-of-the-art fitness prediction benchmark (Notin et al., 2024), focusing on its well-established zero-shot setup.
+We evaluate the models using ProteinGym, the state-of-the-art fitness prediction benchmark (notin2024proteingym), focusing on its well-established zero-shot setup.
 Since the zero-shot setup only provides a test set without any data split, we also validate ProteinTTT on independent data.
-To achieve this, we create a new fitness prediction dataset mined from MaveDB, a public repository containing datasets from Multiplexed Assays of Variant Effect (MAVEs) (Esposito et al., 2019).
-Following ProteinGym, we measure performance on both datasets using Spearman correlation between predicted and experimental fitness values.
+To achieve this, we create a new fitness prediction dataset mined from MaveDB, a public repository of Multiplexed Assays of Variant Effect (MAVEs) (esposito2019mavedb).
+Following ProteinGym, we report Spearman correlation between predicted and experimental fitness values.
 Results.
@@
 We observe that ProteinTTT primarily improves performance for proteins with low MSA depth (i.e., the number of available homologous sequences), suggesting that single-sequence customization enhances predictions for proteins with fewer similar sequences in the training data (Table˜A4).
-The fact that ProteinTTT more effectively improves the performance of smaller ESM2 and SaProt models compared to their larger variants may be a result of the benchmark performance being saturated for larger models, consistent with a recent observation (Notin, 2025).
-We provide a qualitative example showing how ESM2 (650M) + ProteinTTT significantly improves fitness prediction by capturing residues critical for protein stability (Figure˜A5).
-We also demonstrate that customization can be combined with evolutionary information from MSA to further boost fitness prediction (Appendix˜B).
+The fact that ProteinTTT more effectively improves the performance of smaller ESM2 and SaProt models compared to their larger variants may be a result of the benchmark performance being saturated for larger models, consistent with a recent observation (notin2025scalingwall).
+We provide a qualitative example showing how ESM2 (650M) + ProteinTTT significantly improves fitness prediction by capturing residues critical for protein stability (Figure˜A6).
+We also demonstrate that customization can be combined with evolutionary information from MSA to further boost fitness prediction (Appendix˜C).
 4.3 Protein function prediction
 Finally, we demonstrate a proof of concept for customization in the context of protein function prediction.
-We experiment with two tasks: predicting protein location within a cell (Stärk et al., 2021), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (Samusevich et al., 2025).
-Appendix˜C shows that per-protein customization with ProteinTTT consistently enhances the performance of representative models on both tasks.
+We experiment with two tasks: predicting protein location within a cell (stark2021light), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (samusevich2024discovery).
+Appendix˜D shows that per-protein customization with ProteinTTT consistently enhances the performance of representative models on both tasks.
 5 Case studies
-ProteinTTTcan be readily incorporated into structure, fitness, or function prediction pipelines by adding several lines of code (Appendix˜D).
-Here, we demonstrate two challenging structure prediction case studies: improving modeling of antibody–antigen loops (Section˜5.1) and expanding known structures of viral proteins (Section˜5.2).
+ProteinTTTcan be incorporated into structure, fitness, or function prediction pipelines with a few lines of code (Appendix˜E).
+Here, we demonstrate two challenging case studies: improving modeling of antibody–antigen loops (Section˜5.1) and expanding known structures of viral proteins (Section˜5.2).
 5.1 Modeling antibody–antigen loops
@@
 The CDR regions H1, H2, and H3, i.e., the parts of the antibody that bind to the antigen, are highlighted with spheres, while black lines show the alignment error between the ground-truth CDR structure (white) and the predictions (colored).
-Accurately predicting structures of antibodies (e.g., human defensive proteins) and antigens (e.g., viral proteins) enables rational design of new therapeutics (Bennett et al., 2025).
+Accurately predicting structures of antibodies (e.g., human defensive proteins) and antigens (e.g., viral proteins) enables rational design of new therapeutics (bennett2025atomically).
 However, the presence of highly variable loop regions makes modeling of these interactions a long-standing challenge.
-Here, we show that ProteinTTT substantially improves structure prediction for these loop-formed complementarity-determining regions (CDRs) of antibodies, i.e., the parts that bind antigens, as well as for antigens themselves, on the well-established SAbDab dataset (Dunbar et al., 2014).
+Here, we show that ProteinTTT substantially improves structure prediction for these loop-formed complementarity-determining regions (CDRs) of antibodies, i.e., the parts that bind antigens, as well as for antigens themselves, on the well-established SAbDab dataset (dunbar2014sabdab).
 We take the structures from SAbDab that are not predicted well by ESMFold (pLDDT < 70) and show that ProteinTTT improves the LDDT score for 115 of 175 antibody CDR substructures (66%) and 487 of 814 antigen chains (60%).
-As shown in Figure˜4a, ESMFold + ProteinTTT achieves significantly higher average LDDT scores compared to general-purpose ESMFold.
+As shown in Figure˜4a, ESMFold + ProteinTTT achieves significantly higher average LDDT scores compared to general-purpose ESMFold (paired t-test p-value < 0.05).
 Figure˜4b illustrates how ProteinTTT enables accurate prediction of all three CDRs in an antibody chain, providing an improved understanding of its binding interface with the corresponding antigen.
@@
 Higher pLDDT values are better.
-Predicting the structures of viral proteins is vital for vaccine development, antiviral design, and understanding infection (Bravi, 2024).
-Nevertheless, it remains challenging due to the high mutation rate, which often leaves viral proteins without close homologs or experimental structures in databases (Kim et al., 2025).
-Here, we demonstrate that per-protein customized predictions with ESMFold + ProteinTTT improve viral protein structure prediction, substantially expanding the Big Fantastic Virus Database—the repository of viral protein structures (Kim et al., 2025).
-Among all the entries in BFVD, predicted with AlphaFold2 through ColabFold (Mirdita et al., 2022) using MSAs constructed from Logan (Chikhi et al., 2024), only 55% have high-quality structure predictions (pLDDT > 70).
+Predicting the structures of viral proteins is vital for vaccine development, antiviral design, and understanding infection (bravi2024development).
+Nevertheless, it remains challenging due to the high mutation rate, which often leaves viral proteins without close homologs or experimental structures in databases (kim2025bfvd).
+Here, we demonstrate that per-protein customized predictions with ESMFold + ProteinTTT improve viral protein structure prediction, substantially expanding the Big Fantastic Virus Database—the comprehensive repository of 351,242 viral protein structures (kim2025bfvd).
+Among all the entries in BFVD, predicted with AlphaFold2 through ColabFold (mirdita2022colabfold) using MSAs constructed from Logan (chikhi2024logan), only 55% have high-quality structure predictions (pLDDT > 70).
 We apply ESMFold and ESMFold + ProteinTTT to the BFVD entries to expand the database with higher-quality structures.
@@
 While ESMFold manages to improve the predicted structure (as measured by pLDDT) for 10% of the BFVD proteins, ESMFold + ProteinTTT leads to an improvement for 19% of the dataset entries, substantially increasing the quality of known viral protein structures (Figure˜5a).
-We validate that the improved pLDDT confidence values from ESMFold + ProteinTTT correlate with the quality of the predicted structures, as measured by LDDT against reference AlphaFold2 structures having pLDDT > 90 (Pearson = 0.875; Figure˜A9).
+We validate that the improved pLDDT confidence values from ESMFold + ProteinTTT correlate with the quality of the predicted structures, as measured by LDDT against reference AlphaFold2 structures having pLDDT > 90 (Pearson = 0.875; Figure˜A11).
 Notably, the largest improvements in pLDDT align with the largest improvements in LDDT (Figure˜5b).
@@
 This work was also supported by the CETOCOEN EXCELLENCE Teaming project supported from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 857560.
-This work was also supported by the European Union (CLARA (No.
-101136607), ERC FRONTIER (no.
+This work was also supported by the European Union, ERC FRONTIER (No.
 101097822), ELIAS (No.
@@
 TEREP TN02000122/001N).
+This work was also supported by the EU’s Horizon Europe Programme under the Grant agreement No.
+101136607 (CLARA Project), and was co-funded by the EU from the Operational Programme Jan Amos Komenský (OP JAK) (project “Center for Artificial Intelligence and Quantum Computing in System Brain Research, reg.
+no.
+CZ.02.01.01/00/23_029/0008437).
 This work was also supported by the Czech Science Foundation (GA CR) grant 21-11563M and by the European Union’s Horizon Europe program (ERC, TerpenCode, 101170268).
@@
 acknowledges support by the National Research Foundation of Korea grants (RS-2020-NR049543, RS-2021-NR061659 and RS-2021-NR056571, RS-2024-00396026), Creative-Pioneering Researchers Program and Novo Nordisk Foundation (NNF24SA0092560).
+This work was supported with funding from Google.org via the Google PhD Fellowship (R.B.).
 Appendix
 Contents
-appendix.Aappendix.Bappendix.Cappendix.Dappendix.Esubsection.E.1subsubsection.E.1.1subsubsection.E.1.2subsubsection.E.1.3subsection.E.2subsubsection.E.2.1subsubsection.E.2.2subsubsection.E.2.3subsection.E.3subsubsection.E.3.1subsubsection.E.3.2subsubsection.E.3.3appendix.Fsubsection.F.1subsection.F.2appendix.Gsubsection.G.1subsection.G.2subsection.G.3
+appendix.Aappendix.Bappendix.Cappendix.Dappendix.Eappendix.Fsubsection.F.1subsubsection.F.1.1subsubsection.F.1.2subsubsection.F.1.3subsection.F.2subsubsection.F.2.1subsubsection.F.2.2subsubsection.F.2.3subsection.F.3subsubsection.F.3.1subsubsection.F.3.2subsubsection.F.3.3appendix.Gsubsection.G.1subsection.G.2appendix.Hsubsection.H.1subsection.H.2subsection.H.3appendix.I
 Appendix A Justification of customization via perplexity minimization
@@
 Higher TM-scores are associated with lower perplexity, indicating that better predictions are linked to lower uncertainty in the language model’s understanding of the protein sequence.
-While the paradigm of test-time customization has been investigated in other domains, the reasons behind its surprising effectiveness are not completely clear (Liu et al., 2021; Zhao et al., 2023).
+While the paradigm of test-time customization has been investigated in other domains, the reasons behind its surprising effectiveness are not completely clear (liu2021ttt; zhao2023pitfalls).
 Here, we offer a potential justification for the effectiveness of ProteinTTT by linking it to perplexity minimization.
-Perplexity has traditionally been used in natural language processing to evaluate how well models comprehend sentences (Brown, 2020; Chelba et al., 2013).
-Protein language modeling has adopted this metric to assess how effectively models “understand” amino acid sequences (Hayes et al., 2024; Lin et al., 2023).
-For bidirectional, random masking language models, which are the focus of this study, we consider the following definition of perplexity [^2] Please note that this is an approximation of perplexity, which is computationally intractable for bidirectional models, and is often referred to as pseudo-perplexity (Lin et al., 2023; Salazar et al., 2019).:
+Perplexity has traditionally been used in natural language processing to evaluate how well models comprehend sentences (brown2020language; chelba2013one).
+Protein language modeling has adopted this metric to assess how effectively models “understand” amino acid sequences (hayes2024simulating; lin2023evolutionary).
+For bidirectional, random masking language models, which are the focus of this study, we consider the following definition of perplexity [^2] Please note that this is an approximation of perplexity, which is computationally intractable for bidirectional models, and is often referred to as pseudo-perplexity (lin2023evolutionary; salazar2019masked).:
 $\text{Perplexity}(x)=\exp\Biggl(\frac{1}{|x|}\sum_{i=1}^{|x|}{-\log{p(x_{i}|x_{\setminus i};\theta)}}\Biggl),$ (3)
@@
 A perplexity value of 1 indicates that the model perfectly fits the sequence, accurately predicting all the true tokens.
-Several studies have shown that lower perplexity on held-out protein sequences (calculated through the self-supervised track $g\circ f$) correlates with better performance on downstream tasks (via the supervised track $h\circ f$), such as predicting protein contacts (Rao et al., 2020), structure (Lin et al., 2023), or fitness (Kantroo et al., 2024).
+Several studies have shown that lower perplexity on held-out protein sequences (calculated through the self-supervised track $g\circ f$) correlates with better performance on downstream tasks (via the supervised track $h\circ f$), such as predicting protein contacts (rao2020transformer), structure (lin2023evolutionary), or fitness (kantroo2024pseudo).
 To give an example, we analyze the correlation between perplexity and structure prediction quality (Figure˜A1; see Section˜4.1 for experimental details).
-A notable correlation suggests that reducing a model’s perplexity on a single target sample $x$ (applied independently to all test samples) can lead to improved predictions on the downstream task (Figure˜3; Figure˜A10).
+A notable correlation suggests that reducing a model’s perplexity on a single target sample $x$ (applied independently to all test samples) can lead to improved predictions on the downstream task (Figure˜3; Figure˜A9).
 Since we assume only a single target example $x$, the minimization of the masked language modeling loss $\mathcal{L}(x;\theta)$ (Equation˜2) on this example is directly linked to minimizing the perplexity Perplexity(x)(x) (Equation˜3).
 For instance, in the case of a single masked position (i.e., $|M|=1$), the loss is equal to the logarithm of perplexity.
-More generally, it can be shown formally that by minimizing the masked language modeling objective, the model learns to approximate the conditional marginals of the language (of proteins), including the leave-one-out probabilities evaluated in perplexity (Hennigen & Kim, 2023).
+More generally, it can be shown formally that by minimizing the masked language modeling objective, the model learns to approximate the conditional marginals of the language (of proteins), including the leave-one-out probabilities evaluated in perplexity (hennigen2023deriving).
 As a result, applying self-supervised test-time customization on $x$ through $g\circ f$ enhances the representation of the target protein in the backbone $f$, leading to improved downstream performance via the fine-tuning track $h\circ f$.
-Appendix B Customization with multiple sequence alignment (MSA)
+Appendix B Customization beyond masked language modeling
+In this work, we primarily focus on protein language models pretrained with masked language modeling (MLM), where a fixed proportion of randomly selected tokens (e.g., 15%) are masked for training.
+To date, MLM has been the dominant paradigm in protein representation learning.
+Nevertheless, we also provide a proof of concept showing that ProteinTTT can be applied to autoregressive and discrete diffusion–based protein language models, with details provided in the corresponding paragraphs below.
+Furthermore, in Appendix˜I we discuss how ProteinTTT could be extended beyond protein language models.
+Autoregressive customization objective.
+To perform single-sequence customization in an autoregressive setting (i.e., customization of ProGen2 (nijkamp2023progen2)), we apply a standard teacher forcing procedure (vaswani2017attention) with a batch size of one.
+Specifically, each ProteinTTT step optimizes next token prediction across the whole sequence in parallel via the following loss function:
+$\mathcal{L}_{\text{AR}}(x;\theta)=\frac{1}{|x|}\sum_{i=1}^{|x|}-\log p(x_{i}\mid x_{<i};\theta),$ (4)
+where $x$ denotes a sequence of protein tokens, and $p(x_{i}\mid x_{<i};\theta)\doteq g(f(x_{<i};\theta))_{x_{i}}$ is the probability assigned by the model to the true token $x_{i}$ given all preceding tokens $x_{<i}$.
+Here, we use the notation consistent with Equation˜2.
+Discrete diffusion customization objective.
+Recently, discrete diffusion protein language models have emerged as an extension of MLM-based protein language models.
+Instead of masking a fixed ratio of tokens, discrete diffusion approaches vary the masking ratio during training according to a diffusion schedule (hsieh2025elucidating; wang2024dplm; wang2024diffusion; campbell2024generative; alamdari2023protein).
+This has been shown to improve representation learning and to enable sequence generation by starting from a fully masked sequence and gradually denoising it (wang2024diffusion).
+In this work, we experiment with the DPLM2 Bit-based discrete diffusion model (hsieh2025elucidating) for protein structure prediction.
+Interestingly, we find that using a standard MLM objective with a fixed 15% masking ratio for customization (Equation˜2) already improves performance.
+Exploring modifications of the customization objective tailored specifically to discrete diffusion models presents an exciting direction for future work.
+Appendix C Customization with multiple sequence alignment (MSA)
+Table A1: ProteinTTT can be used with MSA when available. Please see Table˜4 for evaluation details.
+Method Avg. Spearman $\uparrow$
+ESM2 (lin2023evolutionary) 0.4139
+[…]
+MSA Transformer (rao2021msa) 0.4319
+MSA Transformer + ProteinTTT (Ours) 0.4326 ± 0.00003
 Customization training objective.
-Since many target proteins may not have homologous sequences (Rao et al., 2021) and finding such homologs may be time-consuming (Lin et al., 2023), the ProteinTTT customization objective (Equation˜2) only assumes a single target sequence for customization.
+Since many target proteins may not have homologous sequences (rao2021msa) and finding such homologs may be time-consuming (lin2023evolutionary), the ProteinTTT customization objective (Equation˜2) only assumes a single target sequence for customization.
 However, we also extend the loss function to the case when a multiple sequence alignment (MSA) is available:
@@
 Results for fitness prediction.
-Table A1: ProteinTTT can be used with MSA when available. Please see Table˜4 for evaluation details.
-Method Avg. Spearman $\uparrow$
-ESM2 (Lin et al., 2023)
-0.4139
 […]
-0.4299 ± 0.00099
-MSA Transformer (Rao et al., 2021)
-0.4319
-MSA Transformer + ProteinTTT (Ours)
-0.4326 ± 0.00003
-[…]
-Appendix C Customization for protein function prediction
-Protein function prediction is essential for understanding biological processes and guiding bioengineering, but is challenging due to its vague definition and limited data (Yu et al., 2023; Radivojac & et al., 2013; Stärk et al., 2021; Mikhael et al., 2024; Samusevich et al., 2025).
-While improved structure prediction with ProteinTTT (Section˜4.1) can already enhance function prediction (Song et al., 2024), we also evaluate our customization method directly on two function classification tasks: subcellular localization, predicting protein location within a cell (Stärk et al., 2021), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (Christianson, 2017; Samusevich et al., 2025).
-Using ProteinTTT with EnzymeExplorer (Samusevich et al., 2025) for TPS detection and Light attention (Stärk et al., 2021) for subcellular localization, we achieve consistent performance gains.
-Evaluation setup.
+Appendix D Customization for protein function prediction
 Figure A2: Customization with ProteinTTT enables the correct substrate classification for a terpene synthase (TPS) enzyme.
@@
 The bar plots also display the predicted probabilities for other substrates with non-zero values (grey).
-For the terpene substrate classification, we use the largest available dataset of characterized TPS from Samusevich et al.
-(2025) and reuse the original cross-validation schema.
-In the case of protein localization prediction, we use a standard DeepLoc dataset (Almagro Armenteros et al., 2017) as a validation set and setHard from Stärk et al.
-(2021) as the test set.
+Protein function prediction is essential for understanding biological processes and guiding bioengineering, but is challenging due to its vague definition and limited data (yu2023enzyme; radivojac2013large; stark2021light; mikhael2024clipzyme; samusevich2024discovery).
+While improved structure prediction with ProteinTTT (Section˜4.1) can already enhance function prediction (song2024accurately), we also evaluate our customization method directly on two function classification tasks: subcellular localization, predicting protein location within a cell (stark2021light), and substrate classification for terpene synthases (TPS), enzymes producing the largest class of natural products (christianson2017structural; samusevich2024discovery).
+Using ProteinTTT with EnzymeExplorer (samusevich2024discovery) for TPS detection and Light attention (stark2021light) for subcellular localization, we achieve consistent performance gains.
+Evaluation setup.
+For the terpene substrate classification, we use the largest available dataset of characterized TPS from samusevich2024discovery and reuse the original cross-validation schema.
+In the case of protein localization prediction, we use a standard DeepLoc dataset (almagro2017deeploc) as a validation set and setHard from stark2021light as the test set.
 Given a protein, the goal of function prediction is to correctly classify it into one of the predefined functional annotations.
-We assess the quality of the TPS substrate prediction using standard multi-label classification metrics used in the EnzymeExplorer paper (Samusevich et al., 2025): mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC).
-In the case of protein localization prediction, we similarly use the classification metrics from the original paper (Stärk et al., 2021): accuracy, multi-class Matthews correlation coefficient (MCC), and F1-score.
+We assess the quality of the TPS substrate prediction using standard multi-label classification metrics used in the EnzymeExplorer paper (samusevich2024discovery): mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC).
+In the case of protein localization prediction, we similarly use the classification metrics from the original paper (stark2021light): accuracy, multi-class Matthews correlation coefficient (MCC), and F1-score.
 Results.
 Table A2: Customization with ProteinTTT improves protein function prediction.
-For the terpene syntase (TPS) substrate classification task, the metrics are computed on the 512 TPS sequences based on the cross-validation schema of the TPS dataset (Samusevich et al., 2025).
-Subcellular localization prediction performance is reported for 432 protein sequences from the setHard test set (Stärk et al., 2021).
+For the terpene syntase (TPS) substrate classification task, the metrics are computed on the 512 TPS sequences based on the cross-validation schema of the TPS dataset (samusevich2024discovery).
+Subcellular localization prediction performance is reported for 432 protein sequences from the setHard test set (stark2021light).
 The error bars show standard deviations across five random seeds.
 TPS substrate classification
-Method
-mAP $\uparrow$ AUROC $\uparrow$
-EnzymeExplorer
-(Samusevich et al., 2025)
-0.805 0.948
-EnzymeExplorer + ProteinTTT
-0.811 ± 0.0011 0.950 ± 0.0002
+Method mAP $\uparrow$ AUROC $\uparrow$
+EnzymeExplorer (samusevich2024discovery) 0.805 0.948
+EnzymeExplorer + ProteinTTT (Ours) 0.811 ± 0.0011 0.950 ± 0.0002
 Subcellular localization prediction
-Method
-Accuracy $\uparrow$ MCC $\uparrow$ F1-score $\uparrow$
-Light attention
-(Stärk et al., 2021)
-0.627 0.549 0.618
-Light attention + ProteinTTT
-0.634 ± 0.004 0.557 ± 0.005 0.627 ± 0.004
+Method Accuracy $\uparrow$ MCC $\uparrow$ F1-score $\uparrow$
+Light attention (stark2021light) 0.627 0.549 0.618
+Light attention + ProteinTTT (Ours) 0.634 ± 0.004 0.557 ± 0.005 0.627 ± 0.004
 Customization with ProteinTTT improves model performance on both of the protein function prediction tasks and across all considered metrics (Table˜A2).
@@
 We hypothesize that improvement with customization is more challenging in classification tasks, as opposed to regression problems, because a larger change in the latent space is required to shift the top-class probability.
-Appendix D Implementation details
-Infrastructure.
-All experiments with ProteinTTT are conducted on machines equipped with a single NVIDIA A100 40GB GPU, an 8-core AMD processor, and 128 GB of physical memory.
-Source code.
-We provide a user-friendly and easily extensible PyTorch (Paszke, 2019) implementation of ProteinTTT, available as the proteinttt Python package  [^3] https://github.com/anton-bushuiev/ProteinTTT .
-We provide two Python code snippets ˜1 and ˜2 to demonstrate the implementation of inference and customization with ProteinTTT, respectively.
-˜1 demonstrates how inference with ESMFold can be enhanced with ProteinTTT by adding just a few lines of code to enable customization.
-Next, ˜2 shows how ProteinTTT can be easily implemented for a PLM of interest by inheriting from the abstract TTTModule class.
-To integrate ProteinTTT within a model (e.g., ESM2), the user needs to implement methods that define the model’s vocabulary, an interface for predicting logits, and a specification of which modules need to be fine-tuned or remain frozen.
-The rest, i.e., the test-time training logic itself, is implemented within the unified TTTModule class.
-Listing 1: Incorporation of ProteinTTT into an ESMFold structure prediction pipeline using the proteinttt package.
+Appendix E Implementation details
+Code snippet 1: Incorporation of ProteinTTT into an ESMFold structure prediction pipeline using the proteinttt package.
 1import esm
@@
 33# =======================================================================
-Optimization.
-We minimize the loss defined in Equation˜2 using stochastic gradient descent (SGD) with zero momentum and zero weight decay (Ruder, 2016).
-While a more straightforward option might be to use the optimizer state from the final pre-training step, this approach is often impractical because the optimizer parameters are usually not provided with the pre-trained model (Hayes et al., 2024; Lin et al., 2023).
-Moreover, many models are pre-trained using the Adam optimizer (Kingma & Ba, 2015) or its variants (Loshchilov & Hutter, 2019).
-However, it was shown that Adam results in less predictable behavior of test-time training compared to the SGD optimizer, possibly due to its more exploratory behavior (Gandelsman et al., 2022).
-Customizing large models.
-We aim for customization to be applicable on the fly, i.e., without the need for any pre-computation and on a single GPU with a minimum computational overhead.
-Since state-of-the-art models for many protein-oriented tasks are typically large, with up to billions of parameters, our aim presents two key challenges.
-First, when using pre-trained Transformers on a single GPU, even for the forward pass, the batch size is typically limited to only several samples due to the quadratic complexity of the inference (Vaswani, 2017).
-Second, for the backward pass, even a batch size of one is not always feasible for large models.
-To address the first challenge, we perform forward and backward passes through a small number of training examples and accumulate gradients to simulate updates with any batch size.
-We address the second challenge by employing low-rank adaptation (LoRA; Hu et al.
-(2021)), which in practice enables fine-tuning of any model for which a forward pass on a single sample is feasible, due to a low number of trainable parameters.
-Section˜G.3 details how ESMFold (Lin et al., 2023), with its 3B-parameter ESM2 backbone $f$, can be efficiently customized, retaining its speed advantage while enhancing performance.
-Listing 2: Implementation of ESM2 + ProteinTTT within the proteinttt package.
+Code snippet 2: Implementation of ESM2 + ProteinTTT within the proteinttt package.
 1import torch
@@
 46 return self(batch)["logits"]
-Appendix E Experimental details
-In this section, we describe the proposed benchmark suite for the three customization tasks considered in this work: protein structure prediction (Section˜E.1), protein fitness prediction (Section˜E.2), and protein function prediction (Section˜E.3).
+Infrastructure.
+All experiments with ProteinTTT are conducted on machines equipped with a single NVIDIA A100 40GB GPU, an 8-core AMD processor, and 128 GB of physical memory.
+Source code.
+We provide a user-friendly and easily extensible PyTorch (paszke2019pytorch) implementation of ProteinTTT, available as the proteinttt Python package  [^3] https://github.com/anton-bushuiev/ProteinTTT .
+We provide Code snippet˜1 and Code snippet˜2 in Python to demonstrate the implementation of inference and customization with ProteinTTT, respectively.
+Code snippet˜1 demonstrates how inference with ESMFold can be enhanced with ProteinTTT by adding just a few lines of code to enable customization.
+Next, Code snippet˜2 shows how ProteinTTT can be easily implemented for a PLM of interest by inheriting from the abstract TTTModule class.
+To integrate ProteinTTT within a model (e.g., ESM2), the user needs to implement methods that define the model’s vocabulary, an interface for predicting logits, and a specification of which modules need to be fine-tuned or remain frozen.
+The rest, i.e., the test-time training logic itself, is implemented within the unified TTTModule class.
+Optimization.
+We minimize the loss defined in Equation˜2 using stochastic gradient descent (SGD) with zero momentum and zero weight decay (ruder2016overview).
+While a more straightforward option might be to use the optimizer state from the final pre-training step, this approach is often impractical because the optimizer parameters are usually not provided with the pre-trained model (hayes2024simulating; lin2023evolutionary).
+Moreover, many models are pre-trained using the Adam optimizer (kingma2015adam) or its variants (loshchilov2019decoupled).
+However, it was shown that Adam results in less predictable behavior of test-time training compared to the SGD optimizer, possibly due to its more exploratory behavior (gandelsman2022test).
+Customizing large models.
+We aim for customization to be applicable on the fly, i.e., without the need for any pre-computation and on a single GPU with a minimum computational overhead.
+Since state-of-the-art models for many protein-oriented tasks are typically large, with up to billions of parameters, our aim presents two key challenges.
+First, when using pre-trained Transformers on a single GPU, even for the forward pass, the batch size is typically limited to only several samples due to the quadratic complexity of the inference (vaswani2017attention).
+Second, for the backward pass, even a batch size of one is not always feasible for large models.
+To address the first challenge, we perform forward and backward passes through a small number of training examples and accumulate gradients to simulate updates with any batch size.
+We address the second challenge by employing low-rank adaptation (LoRA; hu2021lora), which in practice enables fine-tuning of any model for which a forward pass on a single sample is feasible, due to a low number of trainable parameters.
+Section˜H.3 details how ESMFold (lin2023evolutionary), with its 3B-parameter ESM2 backbone $f$, can be efficiently customized, retaining its speed advantage while enhancing performance.
+Appendix F Experimental details
+In this section, we describe the proposed benchmark suite for the three customization tasks considered in this work: protein structure prediction (Section˜F.1), protein fitness prediction (Section˜F.2), and protein function prediction (Section˜F.3).
 Each subsection describes the application of ProteinTTT to the respective models, along with details on the data, metrics, and models.
 Table˜A3 additionally summarizes the hyperparameters used for the application of ProteinTTT to individual models.
-E.1 Protein structure prediction
-E.1.1 Datasets
+F.1 Protein structure prediction
+F.1.1 Datasets
 CAMEO dataset.
-To evaluate the capabilities of ProteinTTT on protein structure prediction, we employ the CAMEO validation and test sets as described in Lin et al.
-(2023).
-Specifically, the validation set was obtained by querying the CAMEO (Continuous Automated Model Evaluation) web server [^4] https://www.cameo3d.org/modeling (Robin et al., 2021) for entries between August 2021 and January 2022, while the CAMEO test set consists of entries from April 1, 2022, to June 25, 2022.
-Most of the entries in the CAMEO sets are predicted with high accuracy and confidence (Lin et al., 2023).
+To evaluate the capabilities of ProteinTTT on protein structure prediction, we employ the CAMEO validation and test sets as described in lin2023evolutionary.
+Specifically, the validation set was obtained by querying the CAMEO (Continuous Automated Model Evaluation) web server [^4] https://www.cameo3d.org/modeling (robin2021continuous) for entries between August 2021 and January 2022, while the CAMEO test set consists of entries from April 1, 2022, to June 25, 2022.
+Most of the entries in the CAMEO sets are predicted with high accuracy and confidence (lin2023evolutionary).
 Therefore, we subselect the challenging validation and test sets where customization with ProteinTTT is suitable.
-Specifically, we apply two standard criteria: (1) preserving entries with ESMFold pLDDT scores below 70 to filter out high-confidence predictions (Jumper et al., 2021), and (2) selecting entries with ESM2 perplexity scores greater than or equal to 6, ensuring that the predictions are challenging due to poor sequence understanding rather than other factors.
-Additionally, most structures with perplexity scores below 6 are already associated with high-confidence predictions (Figure S5 in Lin et al.
-(2023)).
+Specifically, we apply two standard criteria: (1) preserving entries with ESMFold pLDDT scores below 70 to filter out high-confidence predictions (jumper2021highly), and (2) selecting entries with ESM2 perplexity scores greater than or equal to 6, ensuring that the predictions are challenging due to poor sequence understanding rather than other factors.
+Additionally, most structures with perplexity scores below 6 are already associated with high-confidence predictions (Figure S5 in lin2023evolutionary).
 After filtering, the resulting challenging validation and test sets consist of 27 (out of 378) and 18 (out of 194) targets, respectively.
-E.1.2 Metrics
-To assess the quality of the predicted protein structures with respect to the ground truth structures, we use two standard metrics averaged across the test dataset: TM-score (Zhang & Skolnick, 2004) and LDDT (Mariani et al., 2013).
+F.1.2 Metrics
+To assess the quality of the predicted protein structures with respect to the ground truth structures, we use two standard metrics averaged across the test dataset: TM-score (zhang2004scoring) and LDDT (mariani2013lddt).
 TM-score.
@@
 LDDT scores range from 0 to 100, with higher values indicating better local structural agreement.
-E.1.3 Models
+F.1.3 Models
 ESMFold.
@@
 After each ProteinTTT step, we run $h\circ f$ to compute the pLDDT scores, which allows us to estimate the optimal number of customization steps for each protein based on the highest pLDDT score.
-Since the backbone $f$ is given by the ESM2 model containing 3 billion parameters, we apply LoRA (Hu et al., 2021) to all matrices involved in self-attention.
+Since the backbone $f$ is given by the ESM2 model containing 3 billion parameters, we apply LoRA (hu2021lora) to all matrices involved in self-attention.
 This enables fine-tuning ESMFold + ProteinTTT on a single GPU.
@@
 Since ESMFold is a regression model, it only predicts one solution and does not have a straightforward mechanism for sampling multiple structure predictions.
-Nevertheless, the authors of ESMFold propose a way to sample multiple candidates (Section A.3.2 in Lin et al.
-(2023)).
+Nevertheless, the authors of ESMFold propose a way to sample multiple candidates (Section A.3.2 in lin2023evolutionary).
 To sample more predictions, the masking prediction (ME) method randomly masks 15% (same ratio as during masked language modeling pre-training) of the amino acids before passing them to the language model.
@@
 Since sampling multiple solutions with ESMFold + ME and selecting the best one via pLDDT is analogous to ESMFold + ProteinTTT, we employ the former as a baseline, running the method for the same number of steps.
+HelixFold-Single.
+HelixFold-Single is an MSA-free protein structure prediction model that combines representations from a pretrained protein language model with adapted AlphaFold2 geometric modules (EvoformerS and Structure) to directly predict atomic coordinates (fang2023method).
+We use the official implementation [^6] https://github.com/PaddlePaddle/PaddleHelix/tree/dev/apps/protein_folding/helixfold-single
+HelixFold-Single + ProteinTTT.
+HelixFold-Single shares the main concept with ESMFold, and we combine it with ProteinTTT in the same way as in ESMFold + ProteinTTT.
+DPLM2 Bit-based.
+The DPLM2 Bit-based discrete diffusion protein language model (hsieh2025elucidating) extends DPLM2 by using bit-wise discrete modeling to enhance structure generation capabilities (wang2024dplm).
+DPLM2 is a multi-modal model that jointly models protein sequences and discretized structural tokens within a single discrete diffusion framework.
+In this work, we evaluate DPLM2 Bit-based on the task of structure prediction.
+Structure prediction is performed by initializing the structural tokens with masks and gradually denoising them based on the sequential tokens.
+We use the official implementation [^7] https://github.com/bytedance/dplm with the standard 650M-parameter model, 100 denoising steps, and the denoising strategy set to annealing@1.1:0.1.
+DPLM2 Bit-based + ProteinTTT.
+To apply ProteinTTT to DPLM2 Bit-based, we use the standard masked language modeling objective (Equation˜2).
+See Appendix˜B for further discussion.
+Please also note that we do not use confidence function $c$ with DPLM2 Bit-based as it does not implement pLDDT or any other confidence function for protein structure prediction.
 ESM3.
-Unlike ESMFold, ESM3 is a fully multiple-track, BERT-like model (Devlin, 2018), pre-trained to unmask both protein sequence and structure tokens simultaneously (along with the function tokens).
-The structure tokens in ESM3 are generated via a separately pre-trained VQ-VAE (Razavi et al., 2019) operating on the protein geometry.
+Unlike ESMFold, ESM3 is a fully multiple-track, BERT-like model (devlin2018bert), pre-trained to unmask both protein sequence and structure tokens simultaneously (along with the function tokens).
+The structure tokens in ESM3 are generated via a separately pre-trained VQ-VAE (razavi2019generating) operating on the protein geometry.
 In our experiments, we use the smallest, publicly available version of the ESM3 model (ESM3_sm_open_v0) [^8] https://github.com/evolutionaryscale/esm.
@@
 In our experiments, we set $n=8$, which is the default value provided in the official GitHub repository.
-HelixFold-Single.
-HelixFold-Single is an MSA-free protein structure prediction model that combines representations from a pretrained protein language model with adapted AlphaFold2 geometric modules (EvoformerS and Structure) to directly predict atomic coordinates (Fang et al., 2023).
-We use the official implementation [^7] https://github.com/PaddlePaddle/PaddleHelix/tree/dev/apps/protein_folding/helixfold-single
-HelixFold-Single + ProteinTTT.
-HelixFold-Single shares the main concept with ESMFold, and we combine it with ProteinTTT in the same way as in ESMFold + ProteinTTT.
-E.2 Protein fitness prediction
-E.2.1 Datasets
+F.2 Protein fitness prediction
+F.2.1 Datasets
 ProteinGym.
-ProteinGym [^8] https://github.com/OATML-Markslab/ProteinGym is the standard benchmark for protein fitness prediction (Notin et al., 2024).
+ProteinGym [^9] https://github.com/OATML-Markslab/ProteinGym is the standard benchmark for protein fitness prediction (notin2024proteingym).
 The latest, second version of the dataset includes 217 deep mutation scanning experiments (DMSs) across different proteins.
@@
 Figure A3: Comparison of the standard ProteinGym dataset with the MaveDB dataset constructed in this work.
-A) MaveDB, mined from Esposito et al.
-(2019), includes novel assays even after filtering to ensure distinct proteins from the comprehensive ProteinGym dataset.
+A) MaveDB, mined from esposito2019mavedb, includes novel assays even after filtering to ensure distinct proteins from the comprehensive ProteinGym dataset.
 This is largely because most MaveDB assays post-filtering date to 2024, whereas the latest assays in ProteinGym date to 2023.
@@
 The box plots standardly depict quartiles, medians, and outliers.
-To establish a validation set disjoint from ProteinGym (Notin et al., 2024), we mined MaveDB [^9] https://www.mavedb.org (Esposito et al., 2019).
+To establish a validation set disjoint from ProteinGym (notin2024proteingym), we mined MaveDB [^10] https://www.mavedb.org (esposito2019mavedb).
 As of August 1, 2024, the database contains 1178 Multiplexed Assays of Variant Effects (MAVEs), where each assay corresponds to a single protein, measuring the experimental fitness of its variants.
 We applied quality control filters to remove potentially noisy data.
-Specifically, we ensured that the UniProt identifier (Consortium, 2023) is valid and has a predicted structure available in the AlphaFold DB (Varadi et al., 2022).
+Specifically, we ensured that the UniProt identifier (uniprot2023uniprot) is valid and has a predicted structure available in the AlphaFold DB (varadi2022alphafold).
 We also excluded assays with fewer than 100 variants, as well as those where at least one mutation had a wrongly annotated wild type or where most mutations failed during parsing.
@@
 Finally, for efficiently tuning hyperparameters for fitness prediction models, we sampled 50 proteins (Figure˜A3D), corresponding to 83 assays comprising 134 thousand variants.
-E.2.2 Metrics
+F.2.2 Metrics
 Protein fitness labels are not standardized and can vary across different proteins.
@@
 Spearman by phenotype.
-When computing Spearman correlations, we follow the evaluation protocol proposed in ProteinGym (Notin et al., 2024).
+When computing Spearman correlations, we follow the evaluation protocol proposed in ProteinGym (notin2024proteingym).
 First, for each protein, we compute Spearman correlation scores between the predicted ranks of mutations and their corresponding labels.
@@
 Spearman by MSA Depth.
-Following (Notin et al., 2024), we split the performance by the depth of available multiple sequence alignment (MSA), i.e., the number of homologous sequences available, as provided in ProteinGym: “Low depth”, “Medium depth”, and “High depth”, and report the Spearman correlation for each subset individually (Table˜A4).
-Specifically, the MSA depth categories in ProteinGym are determined using the following thresholds from Hopf et al.
-(2017): “Low” is defined as $N_{eff}/L<1$, “Medium” as $1<N_{eff}/L<100$, and “High” as $N_{eff}/L>100$, where $N_{eff}$ represents the normalized number of effective sequences in the MSA, and $L$ is the sequence length covered in the MSA.
-E.2.3 Models
+Following (notin2024proteingym), we split the performance by the depth of available multiple sequence alignment (MSA), i.e., the number of homologous sequences available, as provided in ProteinGym: “Low depth”, “Medium depth”, and “High depth”, and report the Spearman correlation for each subset individually (Table˜A4).
+Specifically, the MSA depth categories in ProteinGym are determined using the following thresholds from hopf2017mutation: “Low” is defined as $N_{eff}/L<1$, “Medium” as $1<N_{eff}/L<100$, and “High” as $N_{eff}/L>100$, where $N_{eff}$ represents the normalized number of effective sequences in the MSA, and $L$ is the sequence length covered in the MSA.
+F.2.3 Models
 ESM2.
-The ESM2 model is a bidirectional, BERT-like (Devlin, 2018) Transformer trained on millions of protein sequences using masked modeling (Lin et al., 2023).
-The goal of protein fitness prediction is to predict the effects of mutations, and PLMs are often adapted to this task using zero-shot transfer via log odds ratio (Notin et al., 2024; Meier et al., 2021).
+The ESM2 model is a bidirectional, BERT-like (devlin2018bert) Transformer trained on millions of protein sequences using masked modeling (lin2023evolutionary).
+The goal of protein fitness prediction is to predict the effects of mutations, and PLMs are often adapted to this task using zero-shot transfer via log odds ratio (notin2024proteingym; meier2021language).
 Specifically, for a given single- or multi-point mutation, where certain amino acids $T$ are substituted from $x_{i}$ to $x^{m}_{i}$ for each $i\in T$, the fitness prediction via the log odds ratio is defined as:
@@
 Mutations with negative values represent favorable mutations, while positive values correspond to disruptive mutations.
-We follow the ProteinGym benchmark and use this formula (Notin et al., 2024) to evaluate the fitness prediction capabilities of ESM2.
+We follow the ProteinGym benchmark and use this formula (notin2024proteingym) to evaluate the fitness prediction capabilities of ESM2.
 We use the implementation of ESM2 from ProteinGym.
@@
 SaProt.
-We also experiment with a structure-aware protein language model, SaProt (Su et al., 2023).
+We also experiment with a structure-aware protein language model, SaProt (su2023saprot).
 SaProt builds off the ESM2 model but incorporates structural information from predicted protein structures.
 Specifically, SaProt uses the same Transformer architecture but expands its vocabulary by combining the 20 standard amino acid tokens with 20 structural tokens from the 3Di vocabulary, increasing the total alphabet size to 400.
-The 3Di tokens capture the geometry of the protein backbone and are generated using VQ-VAE (Razavi et al., 2019), which projects continuous geometric information into discrete tokens and was trained as part of the Foldseek method (van Kempen et al., 2022).
-Since SaProt is also a protein language model, it also uses Equation˜5 to score variants.
-However, please note that SaProt, as implemented in ProteinGym (Notin et al., 2024), uses a slightly different version of the log odds ratio.
-In SaProt, the conditions in the log probabilities in Equation˜5 are replaced with $x_{\setminus T}$ instead of $x_{\setminus i}$, not assuming the independence of substitutions.
+The 3Di tokens capture the geometry of the protein backbone and are generated using VQ-VAE (razavi2019generating), which projects continuous geometric information into discrete tokens and was trained as part of the Foldseek method (van2022foldseek).
+Since SaProt is also a protein language model, it also uses Equation˜6 to score variants.
+However, please note that SaProt, as implemented in ProteinGym (notin2024proteingym), uses a slightly different version of the log odds ratio.
+In SaProt, the conditions in the log probabilities in Equation˜6 are replaced with $x_{\setminus T}$ instead of $x_{\setminus i}$, not assuming the independence of substitutions.
 During customization with ProteinTTT, we only mask sequential information and leave the structural part of the tokens unchanged, reflecting the original pre-training setup.
-We use the implementation of SaProt from ProteinGym8.
+We use the implementation of SaProt from ProteinGym9.
 SaProt + ProteinTTT.
@@
 ProSST.
-Finally, we experiment with the state-of-the-art fitness predictor, ProSST (Li et al., 2024).
-ProSST primarily improves upon SaProt (Su et al., 2023) by incorporating a larger vocabulary of structural tokens and employing disentangled attention mechanisms.
-Instead of relying on the 3Di alphabet optimized for protein structure search with Foldseek (van Kempen et al., 2022), Li et al.
-(2024) pre-train a new autoencoder to denoise corrupted protein backbones and cluster the resulting latent space using the $K$-means algorithm (Lloyd, 1982).
+We experiment with the state-of-the-art fitness predictor, ProSST (li2024prosst).
+ProSST primarily improves upon SaProt (su2023saprot) by incorporating a larger vocabulary of structural tokens and employing disentangled attention mechanisms.
+Instead of relying on the 3Di alphabet optimized for protein structure search with Foldseek (van2022foldseek), li2024prosst pre-train a new autoencoder to denoise corrupted protein backbones and cluster the resulting latent space using the $K$-means algorithm (lloyd1982least).
 Notably, optimal performance for fitness prediction is achieved with $K=2048$ tokens, compared to just 20 in the 3Di vocabulary used by SaProt.
@@
 ProSST, similarly to ESM2 and SaProt, is pre-trained using masked language modeling applied to protein sequence tokens.
-To score mutations on the ProteinGym benchmark (Notin et al., 2024), ProSST also uses the log-odds ratio, but in a slightly different way compared to ESM2 and SaProt.
+To score mutations on the ProteinGym benchmark (notin2024proteingym), ProSST also uses the log-odds ratio, but in a slightly different way compared to ESM2 and SaProt.
 Specifically, ProSST performs a single forward pass to predict log probabilities, which are then used to score all mutations.
-Formally, this approach modifies the log probability condition in Equation˜5, replacing $x_{\setminus i}$ with $x$.
+Formally, this approach modifies the log probability condition in Equation˜6, replacing $x_{\setminus i}$ with $x$.
 ProSST + ProteinTTT.
 Similarly to ESM2 and SaProt, we treat the Transformer encoder in ProSST as the backbone $f$, the masked language modeling head as the pre-training head $g$, and the log-odds ratio formula as the task-specific head $h$.
+ProGen2.
+For fitness prediction, we additionally experiment with one of the major autoregressive protein language models, ProGen2 (nijkamp2023progen2).
+Specifically, we experiment with ProGen of two sizes: ProGen2-small (151M parameters) and ProGen2-large (2.7B parameters).
+We obtain the pre-trained weights from the official GitHub repository [^11] https://github.com/salesforce/progen/tree/main/progen2.
+For ProGen2-large inference, we use floating-point 16 precision for computational efficiency.
+ProGen2 + ProteinTTT.
+To demonstrate the applicability of ProteinTTT in an autoregressive setting, we apply it to the ProGen2 (nijkamp2023progen2) language model.
+To perform the customization, we use the standard next-token prediction objective on a single target protein, following Equation˜4.
+Please see Appendix˜B for details.
 MSA Transformer.
-Finally, we experiment with MSA Transformer for fitness prediction (Rao et al., 2021).
-Similar to ESM2 (Lin et al., 2023), MSA Transformer is pre-trained on large protein sequence datasets; however, it is trained on multiple sequence alignments (MSAs) rather than individual sequences.
+Finally, we experiment with MSA Transformer for fitness prediction (rao2021msa).
+Similar to ESM2 (lin2023evolutionary), MSA Transformer is pre-trained on large protein sequence datasets; however, it is trained on multiple sequence alignments (MSAs) rather than individual sequences.
 Since MSA Transformer is also a protein language model, it can be used for fitness prediction in the same way as ESM2, as discussed above, by computing the log-odds ratio over the first sequence in the MSA in this case.
@@
 These changes significantly reduce computational time while also slightly improving performance compared to the results reported in ProteinGym.
-This improvement may be explained by the fact that the performance of MSA Transformer saturates with increasing MSA input size (Figure 4 in Rao et al.
-(2021)).
+This improvement may be explained by the fact that the performance of MSA Transformer saturates with increasing MSA input size (Figure 4 in rao2021msa).
 MSA Transformer + ProteinTTT.
 […]
-E.3 Protein function prediction
-E.3.1 Datasets
+F.3 Protein function prediction
+F.3.1 Datasets
 TPS dataset.
-For the evaluation of terpene substrate classification, we use the largest available dataset of characterized TPS enzymes from Samusevich et al.
-(2025) and repurpose the original 5-fold cross-validation schema.
+For the evaluation of terpene substrate classification, we use the largest available dataset of characterized TPS enzymes from samusevich2024discovery and repurpose the original 5-fold cross-validation schema.
 We focus on the most challenging TPS sequences, defined as those predicted by the TPS detector, proposed by the dataset authors, with confidence scores below 0.8.
@@
 setHard.
-For the test evaluation of subcellular location prediction, we use the setHard dataset constructed by Stärk et al.
-(2021).
-The dataset was redundancy-reduced, both within itself and relative to all proteins in DeepLoc (Almagro Armenteros et al.
-(2017); next paragraph), a standard dataset used for training and validating machine learning models.
+For the test evaluation of subcellular location prediction, we use the setHard dataset constructed by stark2021light.
+The dataset was redundancy-reduced, both within itself and relative to all proteins in DeepLoc (almagro2017deeploc; next paragraph), a standard dataset used for training and validating machine learning models.
 The setHard dataset contains 490 protein sequences, each annotated with one of ten subcellular location classes, such as “Cytoplasm” or “Nucleus”.
-Since we use ESM-1b (Rives et al., 2021) in our experiments with the dataset, we further filter the data to 432 sequences that do not exceed a length of 1022 amino acids.
-This step, consistent with Stärk et al.
-(2021), ensures that ESM-1b can generate embeddings for all proteins.
+Since we use ESM-1b (rives2021biological) in our experiments with the dataset, we further filter the data to 432 sequences that do not exceed a length of 1022 amino acids.
+This step, consistent with stark2021light, ensures that ESM-1b can generate embeddings for all proteins.
 DeepLoc.
-For hyperparameter tuning in the subcellular location prediction task, we use the test set from the DeepLoc dataset (Almagro Armenteros et al., 2017).
+For hyperparameter tuning in the subcellular location prediction task, we use the test set from the DeepLoc dataset (almagro2017deeploc).
 Similar to setHard, DeepLoc assigns labels from one of ten subcellular location classes.
@@
 Since setHard was constructed to be independent of DeepLoc, setHard provides a leakage-free source of data for validation.
-E.3.2 Metrics
+F.3.2 Metrics
 mAP, AUROC.
@@
 Therefore, we assess the quality of the predictions using standard multi-label classification metrics such as mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC) averaged across individual classes.
-These metrics were used in the original EnzymeExplorer paper (Samusevich et al., 2025).
+These metrics were used in the original EnzymeExplorer paper (samusevich2024discovery).
 We report the performance by averaging the metric values concatenated across all validation folds from the 5-fold cross-validation schema.
 Accuracy, MCC, F1-score.
-To evaluate the performance of subcellular location prediction methods, we use standard classification metrics as employed in Stärk et al.
-(2021).
-Accuracy standardly measures the ratio of correctly classified proteins, while Matthew’s correlation coefficient for multiple classes (MCC) serves as an alternative to the Pearson correlation coefficient for classification tasks (Gorodkin, 2004).
+To evaluate the performance of subcellular location prediction methods, we use standard classification metrics as employed in stark2021light.
+Accuracy standardly measures the ratio of correctly classified proteins, while Matthew’s correlation coefficient for multiple classes (MCC) serves as an alternative to the Pearson correlation coefficient for classification tasks (gorodkin2004comparing).
 The F1-score, the harmonic mean of precision and recall, evaluates performance from a retrieval perspective, balancing the trade-off between false positives and false negatives.
-E.3.3 Models
+F.3.3 Models
 EnzymeExplorer.
-EnzymeExplorer is a state-of-the-art method for the classification of terpene synthase (TPS) substrates (Samusevich et al., 2025).
+EnzymeExplorer is a state-of-the-art method for the classification of terpene synthase (TPS) substrates (samusevich2024discovery).
 The model consists of two parallel tracks.
-Given a protein sequence, EnzymeExplorer first computes its ESM-1v embedding (Meier et al., 2021) and a vector of similarities to the functional domains of proteins from the training dataset, based on unsupervised domain segmentation of AlphaFold2-predicted structures (Jumper et al., 2021).
+Given a protein sequence, EnzymeExplorer first computes its ESM-1v embedding (meier2021language) and a vector of similarities to the functional domains of proteins from the training dataset, based on unsupervised domain segmentation of AlphaFold2-predicted structures (jumper2021highly).
 The ESM-1v embedding and the similarity vector are then concatenated and processed by a separately trained random forest, which predicts TPS substrate class probabilities.
@@
 Light Attention.
-We use Light attention (Stärk et al., 2021) as a representative baseline for subcellular location prediction.
-Light attention leverages protein embeddings from a language model, which in our case is ESM-1b (Rives et al., 2021).
+We use Light attention (stark2021light) as a representative baseline for subcellular location prediction.
+Light attention leverages protein embeddings from a language model, which in our case is ESM-1b (rives2021biological).
 The model processes per-residue embeddings via a softmax-weighted aggregation mechanism, referred to as light attention, which operates with linear complexity relative to sequence length and enables richer aggregation of per-residue information, as opposed to standard mean pooling.
-We re-train the model using ESM-1b embeddings on the DeepLoc dataset (Almagro Armenteros et al., 2017) using the code from the official GitHub page [^11] https://github.com/HannesStark/protein-localization.
+We re-train the model using ESM-1b embeddings on the DeepLoc dataset (almagro2017deeploc) using the code from the official GitHub page [^13] https://github.com/HannesStark/protein-localization.
 Light attention + ProteinTTT.
@@
 Low-rank adaptation (LoRA) was only used with ESMFold, containing 3 billion parameters in the ESM2 backbone.
-Please note that we did not tune the number of customization steps, as adjusting the learning rate and batch size effectively controls the expected performance under the fixed number of steps, as shown in Figure˜A10.
+Please note that we did not tune the number of customization steps, as adjusting the learning rate and batch size effectively controls the expected performance under the fixed number of steps, as shown in Figure˜A9.
 Therefore, we used 30 steps in all our experiments.
 The only exception was ESM3 + ProteinTTT, where the number of steps was set to 50 during initial experiments with different models/tasks conducted in parallel before standardizing the number of steps to 30.
-Methods marked with an asterisk (“*”) used a slightly different calculation of the loss function.
+Bidirectional methods marked with an asterisk (“*”) used a slightly different calculation of the loss function.
 Specifically, the loss was propagated over all tokens, including special and non-masking tokens, while averaging the loss across all tokens simultaneously, rather than first averaging over sequences.
@@
 ESM2 (650M) + ProteinTTT * 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - -
+ProGen2-small (151M) + ProteinTTT 4e-5, 4e-4, 4e-3 4 4 1, 5, 10, 15, 20 - -
+ProGen2-large (2.7B) + ProteinTTT 1e-5 4 4 10, 15, 20 4, 8 8, 16
 SaProt (35M) + ProteinTTT * 4e-5, 4e-4, 4e-3 4 4, 8, 16, 32 30 - -
@@
 Structure prediction
+ESM3 + ProteinTTT 1e-4, 4e-4, 1e-3 2 1, 4, 16 50 (pLDDT) - -
+DPLM2 Bit-based + ProteinTTT 4e-6, 4e-5, 4e-4, 4e-3 2, 4, 8 2, 4, 8 10 - -
+HelixFold-Single + ProteinTTT 4e-4, 1e-3 4, 8, 16 1 30 (pLDDT) - -
 ESMFold + ProteinTTT 4e-4 4 4, 8, 32, 64 30 (pLDDT) 4, 8, 32 8, 16, 32
-HelixFold-Single + ProteinTTT 4e-4, 1e-3 4, 8, 16 1 30 (pLDDT) - -
-ESM3 + ProteinTTT 1e-4, 4e-4, 1e-3 2 1, 4, 16 50 (pLDDT) - -
 Function prediction
@@
 Light attention + ProteinTTT 4e-4, 1e-3, 3e-3 2 2, 4 30 - -
-Appendix F Case study details
-F.1 Modeling antibody-antigen loops
-We download the SAbDab dataset from the official website [^12] https://opig.stats.ox.ac.uk/webapps/sabdab-sabpred/sabdab(Dunbar et al., 2014).
+Appendix G Case study details
+G.1 Modeling antibody-antigen loops
+We download the SAbDab dataset from the official website [^14] https://opig.stats.ox.ac.uk/webapps/sabdab-sabpred/sabdab(dunbar2014sabdab).
 We apply ProteinTTT to targets with low-confidence ESMFold predictions (pLDDT < 70) and remove sequences longer than 400 residues due to GPU memory limitations.
@@
 We predict the full structures using ESMFold+ProteinTTT (with the same hyperparameters tuned on the CAMEO validation set specified in Table˜A3) and compute LDDT scores against the corresponding PDB structures to assess local errors, which are particularly relevant for loop regions.
-For antibodies, we evaluate the complete structures, while for complementarity-determining regions (CDRs), we extract the CDR substructures as annotated in SAbDab according to Chothia numbering (Chothia & Lesk, 1987) and calculate LDDT on these regions.
-F.2 Expanding known structures of viral proteins
-We use BFVD version archived/2023_02_v2131313https://bfvd.steineggerlab.workers.dev.
-This version contains maximum-pLDDT structures from predictions generated by two strategies: (i) ColabFold (Mirdita et al., 2022) with MSAs constructed using Logan (Chikhi et al., 2024), and (ii) ColabFold with 12 additional recycle steps and MSAs constructed using Logan.
+For antibodies, we evaluate the complete structures, while for complementarity-determining regions (CDRs), we extract the CDR substructures as annotated in SAbDab according to Chothia numbering (chothia1987canonical) and calculate LDDT on these regions.
+G.2 Expanding known structures of viral proteins
+We use BFVD version archived/2023_02_v2151515https://bfvd.steineggerlab.workers.dev.
+This version contains maximum-pLDDT structures from predictions generated by two strategies: (i) ColabFold (mirdita2022colabfold) with MSAs constructed using Logan (chikhi2024logan), and (ii) ColabFold with 12 additional recycle steps and MSAs constructed using Logan.
 In Figure˜5, we also report pLDDT values for BFVD version archived/2023_02_v1, where structures are simply obtained from ColabFold with MSAs from Logan, i.e., strategy (i).
@@
 We use the same hyperparameters tuned on the CAMEO validation set, as specified in Table˜A3, with the exception of 20 instead of 30 steps for computational efficiency.
-Appendix G Extended results
-In this section, we provide additional results on test sets (Section˜G.1), discuss validation performance (Section˜G.2), and analyze the runtime performance of customization (Section˜G.3).
-G.1 Detailed test performance
+Appendix H Extended results
+In this section, we provide additional results on test sets (Section˜H.1), discuss validation performance (Section˜H.2), and analyze the runtime performance of customization (Section˜H.3).
+H.1 Detailed test performance
 In this section, we provide details on the test performance.
@@
 Additionally, we provide a qualitative example illustrating how ProteinTTT substantially improves the correlation between ESM2-predicted fitness and ground-truth stability by better identifying disruptive mutations in the protein core (Figure A6).
-Next, Figure˜A6 shows the distribution of ProteinTTT effects: in many cases, customization has minimal impact on performance; often, it leads to substantial improvements; and in rare cases, customization results in a decrease in performance.
+Next, Figure˜A7 shows the distribution of ProteinTTT effects: in many cases, customization has minimal impact on performance; often, it leads to substantial improvements; and in rare cases, customization results in a decrease in performance.
 This positions ProteinTTT as a method for enhancing prediction accuracy, while a comprehensive analysis of its failure modes remains an important direction for future research.
@@
 We also observe that the overall trend of customization with ProteinTTT generally leads to improved performance, with robust consistency across random seeds.
-However, the progression of the performance curve can be rugged, particularly in classification tasks, where substantial changes in the underlying representations are required to shift the top-predicted class in the discrete probability distribution (Figure˜A8).
-G.2 Validation performance
+However, the progression of the performance curve can be rugged, particularly in classification tasks, where substantial changes in the underlying representations are required to shift the top-predicted class in the discrete probability distribution (Figure˜A12).
+H.2 Validation performance
 This section discusses the performance of ProteinTTT on validation data.
@@
 Table˜A3 provides the grid of hyperparameters explored for each model and its size, as well as specifies the optimal hyperparameters suitable for downstream applications.
-Figure˜A10 demonstrates the trend of hyperparameter tuning with optimal hyperparameter combination balancing underfitting and overfitting to a single target protein.
+Figure˜A9 demonstrates the trend of hyperparameter tuning with optimal hyperparameter combination balancing underfitting and overfitting to a single target protein.
 While most of reasonable hyperparameter configurations lead to overall improvements when using customization with ProteinTTT, poorly chosen hyperparameters can have detrimental effects due to rapid overfitting.
 However, with a reliable predicted confidence measure, such as pLDDT, the appropriate customization step for each protein can be selected to mitigate overfitting.
-Figure˜A11 demonstrates that when using ESM3 + ProteinTTT with pLDDT-based step selection for protein structure prediction, all hyperparameter configurations result in improved performance compared to the base ESM3 model.
-G.3 Runtime performance
+Figure˜A10 demonstrates that when using ESM3 + ProteinTTT with pLDDT-based step selection for protein structure prediction, all hyperparameter configurations result in improved performance compared to the base ESM3 model.
+H.3 Runtime performance
 In this section, we demonstrate that customization with ProteinTTT can be done efficiently, with an acceptable computational overhead.
-Specifically, we show that ESMFold, known for being a faster alternative to more performant methods such as AlphaFold2 (Jumper et al., 2021) or AlphaFold3 (Abramson et al., 2024), still remains in the category of lightweight methods even with ProteinTTT customization (Figure˜A4).
+Specifically, we show that ESMFold, known for being a faster alternative to more performant methods such as AlphaFold2 (jumper2021highly) or AlphaFold3 (abramson2024accurate), still remains in the category of lightweight methods even with ProteinTTT customization (Figure˜A4).
 This observation highlights the practical utility of ProteinTTT.
-For example, ESMFold enabled structural characterization of large metagenomics data (>617 million metagenomic sequences), which would be infeasible with AlphaFold2 (Lin et al., 2023).
+For example, ESMFold enabled structural characterization of large metagenomics data (>617 million metagenomic sequences), which would be infeasible with AlphaFold2 (lin2023evolutionary).
 Nevertheless, the original ESMFold has high confidence predictions only for 36% of sequences from the metagenomic database, while the other 392 million sequences remain with low or medium confidence predictions.
-At the same time, ESMFold + ProteinTTT enables more accurate predictions compared to the original ESMFold (Figure˜A6 suggests that ESMFold + ProteinTTT significantly improves predictions in almost 40% of challenging sequences).
+At the same time, ESMFold + ProteinTTT enables more accurate predictions compared to the original ESMFold (Figure˜A7 suggests that ESMFold + ProteinTTT significantly improves predictions in almost 40% of challenging sequences).
 It means that applying ESMFold + ProteinTTT to these remaining sequences could significantly expand the metagenomic atlas characterized by ESMFold.
 Here, we illustrate this on a similar case study by applying ESMFold + ProteinTTT to more than 300 thousand viral proteins in BFVD (Section˜5.2)
+Appendix I Limitations and future work
+We see two main limitations of the current version of ProteinTTT, which we discuss in detail below.
+Extension to other model types and tasks.
+The current form of the method is only applicable to protein language models (PLMs), i.e., Transformer-based (vaswani2017attention) models pre-trained using bidirectional masked language modeling (rives2021biological) or autoregressive next-token prediction (nijkamp2023progen2).
+Nevertheless, the concept of test-time training can also be extended to many other models in computational biology, which presents exciting opportunities for future research, as our work demonstrates the high potential of this paradigm for the field of computational biology.
+For instance, our central experiments in Section˜4.1 use ESMFold (lin2023evolutionary), which is known to often underperform (lin2023evolutionary) more specifialized multiple sequence alignment (MSA)-based structure predictors such as AlphaFold2 (jumper2021highly), AlphaFold-Multimer (evans2021protein), AlphaFold3 (abramson2024accurate), or Boltz-2 (passaro2025boltz).
+Nevertheless, all of these models can also be extended with test-time training akin to ProteinTTT.
+AlphaFold2, and subsequently AlphaFold-Multimer, use masked modeling of MSA as one of the training objectives to learn powerful pairwise representations in Evoformer.
+The Evoformer backbone could therefore be updated at test time to obtain more powerful representation of one input MSA at a time using the ProteinTTT objective (Section˜3.1).
+While AlphaFold3 and Boltz-2 do not use masked modeling, they can still be customized in a self-supervised way, for example using an optimization through distogram (cho2025boltzdesign1).
+Implementing the variants of ProteinTTT for the aforementioned models could enable customized structure prediction of protein multimers and protein-ligand complexes.
+Beyond structure prediction, test-time customization could also benefit de novo protein design.
+Our results with autoregressive ProGen2 on fitness prediction suggest that ProteinTTT can improve sequence design (Table˜4).
+Similarly, although our experiments with ESM3 are currently conducted in the context of structure prediction (Table˜2), ProteinTTT can be straightforwardly applied to ESM3 for protein design tasks such as inverse folding or structure inpainting by applying ProteinTTT to the corresponding ESM3 input tracks.
+Furthermore, BoltzGen (stark2025boltzgen) can be extended with test-time customization in a manner analogous to Boltz-2, discussed above, due to their shared architecture.
+Performing ProteinMPNN (dauparas2022robust) customization on part of a protein to guide generation of the remaining structure or its binder, as well as customizing RFdiffusion (watson2023novo) to a target structure for binder design, represent promising opportunities in protein design with the potential for higher success rates.
+Better control over failure cases.
+The failure modes of ProteinTTT are not yet fully understood.
+For instance, combining ESMFold with ProteinTTT decreases performance for several proteins in the CAMEO test set (Figure˜A7).
+A detailed analysis of these cases shows that the degradation can be attributed to ambiguity in the evaluation itself (Figure˜A13).
+These examples illustrate the challenge of identifying a general reason for the occasional degradation of performance.
+As discussed in Section˜H.2, confidence functions (such as pLDDT in structure prediction) allow effectively eliminating overfitting to a single protein and thereby mitigating such failure cases, making confidence prediction an essential component of customization.
+While confidence functions begin to emerge across tasks, such as fitness prediction (gurev2025variant; nijkamp2023progen2) and inverse folding (shuai2025sidechain), they are not yet universally available for use with ProteinTTT.
+In particular, for fitness (Section˜4.2) and function (Section˜4.3) prediction, controlling failure cases remains more challenging due to the absence of a reliable confidence metric.
+This motivates the development of general, task-agnostic, unsupervised confidence measures (for example, perplexity-based estimates (gurev2025variant)) or a dedicated confidence prediction module within ProteinTTT (abramson2024accurate; jumper2021highly).
+Another promising direction is deriving confidence estimates from mechanistic interpretability of protein language models (hubotter2025specialization; simon2025interplm; zhang2024protein).
 Table A4: ProteinTTT performance on ProteinGym depending on MSA depth.
-MSA depth reflects the number of available proteins similar to the target protein and, when using large protein language models, can be interpreted as a measure of the representation of similar proteins in the training data (Section˜E.2.2).
+MSA depth reflects the number of available proteins similar to the target protein and, when using large protein language models, can be interpreted as a measure of the representation of similar proteins in the training data (Section˜F.2.2).
 Customization with ProteinTTT primarily improves performance on difficult targets, with low MSA depth.
 Standard deviations are calculated over 5 random seeds but are omitted in the right panel for brevity, where the maximum standard deviation does not exceed 0.0004.
-Avg. Spearman $\uparrow$
-Spearman by MSA depth $\uparrow$
-Low depth Medium depth High depth
-ESM2 (35M) (Lin et al., 2023) 0.3211 0.2394 0.2707 0.451
-ESM2 (35M) + ProteinTTT (Ours) 0.3407 ± 0.00014 0.2445 0.3144 0.4598
-SaProt (35M) (Su et al., 2023) 0.4062 0.3234 0.3921 0.5057
-SaProt (35M) + ProteinTTT (Ours) 0.4106 ± 0.00004 0.3253 0.3972 0.5091
-ESM2 (650M) (Lin et al., 2023) 0.4139 0.3346 0.4063 0.5153
-ESM2 (650M) + ProteinTTT (Ours) 0.4153 ± 0.00003 0.3363 0.4126 0.5075
-SaProt (650M) (Su et al., 2023) 0.4569 0.3947 0.4502 0.5448
-SaProt (650M) + ProteinTTT (Ours) 0.4583 ± 0.00001 0.3954 0.4501 0.5439
-ProSST (K=2048) (Li et al., 2024) 0.5068 0.4731 0.5107 0.5749
-ProSST (K=2048) + ProteinTTT (Ours) 0.5087 ± 0.00004 0.4809 0.5104 0.5750
+Avg.
+Spearman $\uparrow$ Spearman by MSA depth $\uparrow$ Low depth Medium depth High depth ESM2 (35M) (lin2023evolutionary) 0.3211 0.2394 0.2707 0.451 ESM2 (35M) + ProteinTTT (Ours) 0.3407 ± 0.00014 0.2445 0.3144 0.4598 ProGen2-small (151M) (nijkamp2023progen2) 0.3255 0.2974 0.3136 0.3765 ProGen2-small (151M) + ProteinTTT (Ours) 0.3591 ± 0.0002 0.3319 0.3636 0.3917 ProGen2-large (2.7B) (nijkamp2023progen2) 0.3724 0.3504 0.3808 0.4090 ProGen2-large (2.7B) + ProteinTTT (Ours) 0.3817 ± 0.00158 0.3627 0.3791 0.4152 SaProt (35M) (su2023saprot) 0.4062 0.3234 0.3921 0.5057 SaProt (35M) + ProteinTTT (Ours) 0.4106 ± 0.00004 0.3253 0.3972 0.5091 ESM2 (650M) (lin2023evolutionary) 0.4139 0.3346 0.4063 0.5153 ESM2 (650M) + ProteinTTT (Ours) 0.4153 ± 0.00003 0.3363 0.4126 0.5075 SaProt (650M) (su2023saprot) 0.4569 0.3947 0.4502 0.5448 SaProt (650M) + ProteinTTT (Ours) 0.4583 ± 0.00001 0.3954 0.4501 0.5439 ProSST (K=2048) (li2024prosst) 0.5068 0.4731 0.5107 0.5749 ProSST (K=2048) + ProteinTTT (Ours) 0.5087 ± 0.00004 0.4809 0.5104 0.5750
 Table A5: Performance of ProteinTTT on the MaveDB dataset.
 In this work, we use our newly constructed MaveDB dataset as a validation fold for tuning the ProteinTTT hyper-parameters for fitness prediction.
-For computational efficiency, we only select a subset of 50 proteins (Section˜E.2.1) and do not run customization across multiple random seeds to estimate standard deviations.
+For computational efficiency, we only select a subset of 50 proteins (Section˜F.2.1) and do not run customization across multiple random seeds to estimate standard deviations.
 The performance shown was calculated by first aggregating correlations per assay, and then per protein (some assays correspond to the same protein).
 Avg. Spearman $\uparrow$
-ESM2 (35M) (Lin et al., 2023) 0.4458
+ESM2 (35M) (lin2023evolutionary) 0.4458
 ESM2 (35M) + ProteinTTT (Ours) 0.4593
-ESM2 (650M) (Lin et al., 2023) 0.4568
+ESM2 (650M) (lin2023evolutionary) 0.4568
 ESM2 (650M) + ProteinTTT (Ours) 0.4604
-SaProt (650M) (Su et al., 2023) 0.4926
+SaProt (650M) (su2023saprot) 0.4926
 SaProt (650M) + ProteinTTT (Ours) 0.4926
-SaProt (35M) (Su et al., 2023) 0.5251
+SaProt (35M) (su2023saprot) 0.5251
 SaProt (35M) + ProteinTTT (Ours) 0.5271
-ProSST (K=2048) (Li et al., 2024) 0.5444
+ProSST (K=2048) (li2024prosst) 0.5444
 ProSST (K=2048) + ProteinTTT (Ours) 0.5462
@@
 For ESMFold and its variants, the median and interquartile ranges of running times on the CAMEO test set are shown using a single NVIDIA A100 GPU.
-For AlphaFold2, we use estimates from Lin et al.
-(2023).
+For AlphaFold2, we use estimates from lin2023evolutionary.
 Specifically, a forward pass through AlphaFold2 is approximately 60 times more computationally expensive than ESMFold (e.g., AlphaFold2 (no MSA; estimate): $2\times 60=120$ seconds), with additional MSA construction taking at least 10 minutes using standard pipelines (AlphaFold2 (estimate): $2\times 60+10\times 60=720$ seconds).
 ESMFold + ProteinTTT (30 steps) involves LoRA parameter updates, along with forward passes at each customization step to estimate pLDDT and select the structure with the highest predicted confidence.
-Disabling pLDDT significantly reduces computational overhead (ESMFold + ProteinTTT (no pLDDT) compared to ESMFold + ProteinTTT), but may require careful parameter tuning (Section˜G.2).
+Disabling pLDDT significantly reduces computational overhead (ESMFold + ProteinTTT (no pLDDT) compared to ESMFold + ProteinTTT), but may require careful parameter tuning (Section˜H.2).
 Overall, ESMFold + ProteinTTT maintains the speed advantage of ESMFold, and is at least an order of magnitude faster than AlphaFold2.
+Figure A5: Test performance of EnzymeExplorer + ProteinTTT across customization steps.
+The performance is averaged across all 512 proteins in the dataset, with error bars representing the standard deviation across 5 random seeds.
 Figure A6: Example of protein fitness prediction upon single-sequence model customization with ProteinTTT.
-Fitness predictions from ESM2 (650M) show poor correlation with experimental fitness values in the ProteinGym test set measured by the stability assay “UBR5_HUMAN_Tsuboyama_2023_1I2T” (Tsuboyama et al., 2023) (left).
+Fitness predictions from ESM2 (650M) show poor correlation with experimental fitness values in the ProteinGym test set measured by the stability assay “UBR5_HUMAN_Tsuboyama_2023_1I2T” (tsuboyama2023mega) (left).
 ESM2 + ProteinTTT achieves significantly higher correlation, likely due to improved detection of disruptive mutations in the protein core that impact protein stability (middle).
@@
 While an increased number of ProteinTTT steps generally enhances performance, only a few steps (e.g., five) may suffice to achieve significant performance improvement.
-Figure A8: Test performance of EnzymeExplorer + ProteinTTT across customization steps.
-The performance is averaged across all 512 proteins in the dataset, with error bars representing the standard deviation across 5 random seeds.
-Figure A9: ESMFold + ProteinTTT pLDDT correlates with ESMFold + ProteinTTT LDDT.
-The evaluation was performed on 17,582 AlphaFold2 reference structures from the BFVD database with pLDDT > 90.
-Here, $r=0.875$ denotes the Pearson correlation coefficient.
 Figure A9: Dependence on ProteinTTT hyperparameters for customized fitness prediction.
@@
 We conducted a comprehensive grid search based on three key hyperparameters: learning rate (denoted as “lr”), number of gradient accumulation steps (denoted as “grad_steps”; with the batch size of two), and masking strategy (denoted as “mask”).
-We explored two learning rates, 4e-4 and 1e-3, three gradient accumulation step values of 1, 4, and 16, and five different masking strategies: uniform sampling of 0.05, 0.5, and 1.0 fractions of amino acids, as well as the “beta30” and “betalinear30” distributions proposed in the ESM3 paper (Hayes et al., 2024).
+We explored two learning rates, 4e-4 and 1e-3, three gradient accumulation step values of 1, 4, and 16, and five different masking strategies: uniform sampling of 0.05, 0.5, and 1.0 fractions of amino acids, as well as the “beta30” and “betalinear30” distributions proposed in the ESM3 paper (hayes2024simulating).
 Each row in the table presents the mean TM-score and LDDT metrics with standard deviations across five random seeds on the CAMEO validation fold.
@@
 For the hyperparameter search, we used 30 customization steps instead of 50 to reduce computation time.
+Figure A11: ESMFold + ProteinTTT pLDDT correlates with ESMFold + ProteinTTT LDDT.
+The evaluation was performed on 17,582 AlphaFold2 reference structures from the BFVD database with pLDDT > 90.
+Here, $r=0.875$ denotes the Pearson correlation coefficient.
+Figure A12: Magnitude of ESMFold pLDDT improvements after customization with ProteinTTT.
+The evaluation is performed on 317,882 proteins from the Big Fantastic Virus Database (BFVD).
+Percentage annotations indicate the fraction of proteins whose pLDDT increases by at least the corresponding value (e.g., 49.8% of proteins show an improvement of at least 5 pLDDT points).
+Figure A13: Detailed analysis of ProteinTTT failure cases on the CAMEO test set.
+The figure shows the two entries that consistently exhibit a decrease in TM-score after customization with ProteinTTT across most random seeds (see Figure˜A7).
+(a) For chain B of PDB entry 7QII (white), the ground-truth structure is part of a dimer in which the conformation of chain B depends on interactions with chain A (black).
+In the monomeric prediction setting, this context is absent, making the precise helix arrangement inherently ambiguous.
+Both ESMFold and ESMFold + ProteinTTT correctly capture the helical composition but differ in the global configuration, leading to different TM-scores.
+(b) For chain A of PDB entry 7MQ4 (white), the reference structure is an NMR ensemble with substantial conformational variability (black).
+Both ESMFold and ESMFold + ProteinTTT recover the stable substructure (right part of the structure in black consisting of a helix surrounded by beta strands), yet produce different conformations in the flexible regions, where multiple arrangements are plausible.